A Complete Bibliography of Publications in Limnology and Oceanography: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

$0.60 [Ano66a]. $0.75 [Ric69]. $1.80 [Cha66b]. $10.00 [Par66b]. $11.50 [Ise69]. $12.00 [Par66a]. $13.50 [Tyl68a]. $15.00 [Set67]. $15.95 [Pai68]. $16.50 [Ric67]. $17.50 [Koh68]. $18.00 [Edm69]. $19.95 [Gro69]. $2.75 [Spa66]. $2.95 [Pai69]. $23.50 [Bro68a, Bro68b]. $24.00 [Bar68]. $25.00 [PP67a]. $27.00 [Bar68]. $27.60 [Edm67]. $39.00 [Fre69]. $4.50 [Dee69]. $4.95 [Pai69, Ori68]. $45.00 [Wyr67]. $47.00 [Ban67b]. 4π [Mad66]. $5.00 [Bod68, Cha66a]. $5.28 [PP67b]. $5.75 [Par68]. $5.95 [Och67, Pai67]. $6.50 [Roy68]. $6.95 [Koh67, Owe69, Ray68]. $7.00 [Edm68a]. $7.50 [Dee67, Dic68]. $8.40 [Dug68]. $8.50 [Par67a, Stu69, Wyr66]. $8.95 [Aro69]. $9.00 [Ban67a, Bod68]. $9.50 [Fox66, Has68a, Lar68b]. ^{+}+ [WG65]. ^{137} [WH63b, Wil60]. ^{14} [AR67, BW68a, BH67, BB69c, Bug67, Cas62, JS61, LC69, McA61a, Mca61b, Nie65, OM67, Owe63, RM65]. ^{15} [NDDG62]. ^{2} [LP65]. ^{24} [LH62]. ^{32} [BB68, Cus67]. ^{65} [BO60, Cus67, Gut63, MO63, Ost62, Ver62]. ^{90} [LO62a]. _{1}2 [CS66, Dai69, Cur62, Gol64, MS62c]. _{2} [CD67, Mca61b, PH63, Par69, RV61, VBO60, Ver60, Ver61]. $A10.50 [Car68a]. k [Che67]. n [CB67].

-Factor [Che67]. -Paraffins [CB67].

0Microbiology [Bro68a].

100s [Koh68]. 150^{o} [Rei63]. 1952 [Rei63]. 1961 [Hob66, Hul62]. 1964 [Ano63f, BK68]. 1968 [CB69a].

21-cm [Bur68a]. 25c [KP68, KP69]. 27 [Val67]. 27s [Hay68, Hyn68]. 2nd [Pai68].

6-foot [ARNA64]. 60-liter [YBF69]. 67 [Par67a].

7 [CC62a]. 7-dibromo-8-hydroxyquinoline [Top69].

abatement [DH64]. aboard [TH69]. above [Vin67]. Absolute [vA65]. Absorption [Joy68, Yen62]. abstracts [Ano66e]. abundance [Mar68b, Taf60]. Acad [Val67]. Academia [Fre68]. Academic [Koh68]. Academy [Ano65e]. Acartia [Jef62b]. Acclimation [Hah63]. Accumulation [SO61, Jon60b, MT64, SS62a, WDH63]. accuracy [Car65a]. Acid [AM68, CP68b, HC67, TV69, BF62, BF63, KP67, Lew69b, McM67, McM69, Sut65, SS68c, WS66]. acid-soluble [McM69]. acids [GC68, Jef69, Lew69b, STH60, SS61, VBO60, WJ67]. Activation [FM62, RHMB62, MGPB65, WL68]. active [Jar67, JGST67, SS62b]. Activity [HM64, PT68, BW68a, BM65, JS61, LC69, Lor63, Nie65, SWBL63, VJ66]. actuated [Bro69b]. Adaptation [Bro69a]. adapted [BB69b]. addendum [Mul65c]. adder [Hir63]. Adenosine [HHH67, HHB66]. Adirondack [Web61]. adjacent [FAMR60, JM68]. Adsorption [Jar67]. Advancement [Bar68]. Advances [Koh68, Par67a]. Advisory [Ano66a]. Aegean [BK68]. Aequipecten [Mar60]. Aerial [Ell66]. Affecting [Mul63, Con66b, Gol68, PD65, YBW64]. affinis [Wel68b]. Afognak [DD61]. African [BPTW65, Mos69, Tal63]. After [Lam67, Her62]. Ag [THJ67]. against [HJK63]. agardhii [BBK69]. age [ZTTG65]. aggregates [Joh67a, Kan67, Ril63, RWV64, RVW65]. agilis [Par61]. agricultural [GLN64, NGL^{+}64]. aid [HO61b, Pfa62]. air [DB65, DS68, Ise69, Jar67, Rag62, vUB63]. alalunga [PO68]. Alaska [DW60, Fox66, Hob61]. Albacore [PO68]. Alewife [DW65]. Alfred [Sea65a, Sea65b]. alga [BL65, BB69c, EM63, WH63b, WS61]. Algae [Ber61, BO60, Bea69, BB69b, Bro65b, Cas60, FF67, Fit69, Gut63, Joh63b, Lew69a, Mos67b, Nal66, OC69, Sha69, SG69, Uke65, Wil63]. Algal [Mos67a, BB67a, Car68b, EM66, Hic69, Mac61, Mos69, PV69, Sha69, Sto67]. algebraic [Def65]. Algenvegetation [Fot68]. aligner [Fel64]. alkaline [Bea69, Hsi65]. Alkalinity [PWY69, Mai66, Par66c, Run65, TS67]. Allomyces [HM64]. along [Bum65, Des69, PPW62, PD65]. Also [Fre68]. altered [Mac61]. altitude [Tyl61a]. Altitudes [BPTW65]. Am [Par67a]. Amazon [GS67]. Amendment [Ano65a]. America [Hay64, HM69, PO68]. American [Ano60f, Ano61d, Ano64c, Ano65d, Ano66f, Ano67i, Ano68c, Ano69i, Bar68, Bod68, Bro68a, Bro68b, Edm69, Lar68b, Lau62, Lau63, Man64, Spa66, Tyl68a, Ano60a, Ano63a, Ano63b, AH64, Gei64, HM62]. americana [Her63b]. americanus [HM62]. Americas [Ber61]. Amino [TV69, AL64, Jef69, SS61, WJ67]. Ammonia [Mac67, Sol69, GDM64, MS62a, Bro67]. ammonia= [MP65]. ammonia=-nitrogen [MP65]. ammonium [ERM69]. Amphioxi [Cor67]. amphipod [Joh64b]. ampoules [Nie65]. Amsterdam [Dug68, Joy68]. anaerobic [EM63]. Anaerobiosis [Cha69]. analoga [BW68b]. analyses [FM62]. Analysis [TV69, Car65a, CS69b, CD67, FB68, FZ67, GK66, HHSW66, HD69, Klo63, LJ69, MGPB65, RHMB62, Val66, WL68]. Analytical [Par68]. analyzing [MK68]. anchovy [Far60]. Andrew [IJ61]. Andros [Wey63]. Anecdysis [Bar62]. Angeles [BIR64a]. Angino [Joy68]. angiosperm [GK66]. Anguilla [Jef60]. animals [FS67, VTK66, WW68a, vUB63, Pai68]. anion [Top69]. anions [Doy68]. Annapolis [Ise69]. Announcement [Ano62a, Ano64d, Ano65b, Ano67a, Ano69a, Ano64e]. Announcements [Ano64a, Ano67b]. annual [BK68, Gor64, Ham69a, HC68b, LJ69, McM69, Wyr65]. Anodonta [DW65]. anomalies [TS67]. anoxic [RCBA65]. Antarctic [Cor65, GMW63, GMH67, Man67a, RL64, Tib65, Wal69]. Antarctica [AAT64, AH62, BS65, BS61]. Antimicrobial [SWBL63]. antistapi [SWBL63]. antistapi-iylococcal [SWBL63]. Antiyeast [BM65]. Apparatus [HMA65, SV65, HAM68, Mac63, Pam65, Wat65b, WDS64]. apparent [BH67, KP67]. Application [MK68, TM64, Hep62]. Applications [Rit65]. approach [Def65]. aquaria [SCV69]. Aquatic [Edm68a, Ric67, Rit65, Bam62, BB68, Bri67a, Dug68, Fit69, GK66, GB66, Kor68, Sug67, TSB63, WW68a]. aquivalans [Dav62]. aquosus [Sco65]. Arabian [DGR64]. archives [Ege64]. Arctic [Hop69, HEHM60, Lew69b]. Area [AH62, BIR64a, SW61, BIR64b, BW68b, Bum65, Hob67, Hor62, Lor68, MR66, MSE66, SFM69, TvU64]. areal [Hol62]. areas [FAMR60, Pen62]. arenaria [Mat60a, Mat60b]. Arizona [Col61, CWB67]. arsenic [FM62]. artedii [CB69a]. Artemia [PT68]. arthropods [Tea60]. Articles [Ano63a]. articulated [AM67a]. artificial [Hil69a, KDCP67, KB65, KWV66, WDS64]. artificially [HL64]. Arts [Val67]. ascorbic [BF62]. ashfall [Kur66]. Aspects [HMA66]. assay [Gol64, Vis61]. assemblages [SB67]. assessing [HC68b]. assessment [Fag69, Roy68]. assimilated [BB69c]. Assimilation [Con66a, Con66b, CS65, DD65, Man67a, MG62, VJ66]. associated [GB66, Hoe60, SP60, Ver60]. Association [Ano61c, Bar68, KP68, Wil69b]. Associations [HS68, Hic69]. Astatic [Bro69a]. Asterionella [Rao69].

Atlantic [Bum65, GDM66, Man64, EWR65, Hop60, Jef60, Maz67, MCK64, MR64, RWV64]. Atlas [Bod68, Mai66, Coc63]. Atmosphere [Par66b, HB66, PH63]. Atomic [Ano61e, Fox66, Joy68]. Attached [Dri67, Cas60]. attachment [Nis64]. attenuation [HB62, SS62b]. attenuators [MS61]. Attractiveness [Bla60]. August [Rei63]. Australian [Car68a, Car68a]. autoclaved [Jon67b]. automated [Par65]. Automatic [Swa65, BA64, Hir63, Mül65a]. availability [But69, GK66]. available [Fre68, Geo64]. Averaged [Dux62]. Avoidance [FC65, Bri67b]. awatschensis [Heu69].

B [Edm66a, Och67, CS66, Cur62, Dai69, Gol64, MS62c]. Ba [THJ67]. Baas [Woo63]. background [PPF62]. bacteria [BB63, BC60, BC61, CM69, HHH67, Jan67, JSP68, Lig63, Sol64, TeI68, TG68, ZB65]. bacteria-free [Sol64]. Bacterial [Bro68c, AH64, HW65, HC69]. Bacteriastrum [SB66b]. bacteriological [Sie63]. Bacterium [AM68, Wat65a, MH64]. bag [ST61]. Bahama [TS67]. Bahamas [Wey63]. baja [LaF63]. balance [RL64]. Balanced [Zar69]. balances [TBLT68]. balanoides [Bar62, Moh61]. Balanus [Bar62, Moh61]. Balder [Owe69]. Baltimore [Stu69, Wyr66]. Band [Wyr67]. Bank [CT61, TS67]. Barbara [BW68b]. Bardach [Ray68]. barnacle [CP61]. Barnacles [Bar62]. Barnstable [Gor66, SGMH62]. based [MC65, VJ66]. bases [VBO60]. Basin [GE62, EHR62, GLN64, GG64, Hop69, Joh63a, NGL^{+}64, Rao68]. basins [Vin67]. Basis [Lag68]. Bass [MH61]. Batch [SG69, WH63b]. Bathyal [RM68]. Bathymetric [Wil68]. bathythermograph [Bro65c]. Bay [Anr64a, Anr64b, Bri67a, Car65b, FAMR60, GCV63, IJ61, Lid65, Phl65, Rao69, Hoe60, Kor68, Mar60, Anr64a, Anr64b, BW68c, Her63a, HWE67, Jef62a, JM68, Lyn66, Mar65, Mar68b, Mul64, Pat61a, Pra65, Pra66a, Ree69, San60, Squ69, Wie60a]. bays [HL64]. Beach [Lid66, Eme61, Joh67c, MTR64, SB68, BIR64b]. Beaches [Chi60]. beam [IJW61, LM64]. Beaufort [WM66]. beccarii [Bro67]. Becking [Woo63]. beer [YBF69]. Behavior [BB69a, Bur69a, MH61, PNH66, ARNA64, MB67]. Behrman [Aro69]. below [Vin67]. beneath [Hll67]. Bengal [Rao69]. Benthic [DH64, Mac64, PF68, Rot68, RM68, San60, Wie60a, BO60, BW68a, BB67a, Cas67, FAMR60, Gut63, Har69, Hil69a, Joh64b, MW69, PV69, Pam65, RR65, SG69, Wil63]. benthos [CU65, Fag69, HC68b, TO66, WL66]. Bentzel [Eve67]. Bering [CA66]. Berkeley [Edm68a, Koh68]. Berlin [Ban67a, Wyr67, Ban68]. Berlin-Nikolassee [Ban67a, Ban68]. Bermuda [Bee66, Cou69, HR63, IJW61, MS62a, MS62b, Neu66, Sha69, WD66]. Bernard [Edm67]. Between [Bur69b, Cas63a, DW65, Man69, PT68, Par60, Par61, Pra66a, AO63, Bur68b, CP61, DWW68, HR63, MF66, McL65, MSE66, PNH66, PSG65, SS63, SB61b, SB62b]. Biblical [Dee67]. Bibliography [Sea65a, Sea65b]. Bight [Bum65]. Billings [Joy68]. Binnengewässer [Fot68]. Bioassay [CS66, HW65, Mac61, ND66, CS69b]. bioassays [Fit69]. biodeposition [HMA66]. Biodialystat [Par67c]. biogeochemistry [Ano69g]. Biological [Ano60a, HEHM60, Lag68, LO62a, Ric69, Ano61f, BRZ68, CS69b, GC63a, Kur66, McC66, RMV61, Tau69, Vis61]. biologically [HHSW66]. Biologie [Ban67b]. Biology [Fre60b, Fre61b, HM62, Kes60, Pai68, Dic68, EMT65, Fre69, Koh68, Ori68, Pai67, Edm66a]. Bioluminescence [Nes67, CK65, YBW64]. Bioluminescent [Sol66, SFM69]. biomass [HH69]. Biotic [Hoe60]. Birge [Ege64]. Biscayne [FAMR60]. Bivalve [Bla60, Joh66]. bivalves [BR65]. Blackie [Hay68, Hyn68]. Blake [LM64]. Bloom [Rao69, BL65, KC65]. Blue [Sha69, BL65, BB69c, OC69, Rig68, WH63b]. Blue-green [Sha69, BB69c, OC69, WH63b]. boat [Pfa62]. boats [Dri64]. Body [Bur69b, Jaw69, Bur68b, McL65, MO63, Scr60]. bomb [Pai66]. Bonney [AAT64]. Book [Ano66a, Aro69, Ban67a, Ban67b, Ban68, Bar68, Bod68, Bro68a, Bro68b, Car68a, Cha66a, Cha66b, Dee67, Dee69, Dic68, Dug68, Edm66a, Edm66b, Edm67, Edm68a, Edm68b, Edm68c, Edm69, Fot68, Fox66, Fre66, Fre68, Fre69, Gro69, Has68a, Hay68, Hyn68, Ise69, Joy68, Koh67, Koh68, Lag68, Lar68b, Lew69a, Och67, Ori68, Owe69, Pai67, Pai68, Pai69, Par66b, Par66a, PP67a, Par67a, PP67b, Par68, Ray68, Ric67, Ric69, Roy68, Set67, Spa66, Stu69, Tyl68a, Val67, Wyr66, Wyr67]. Books [Ano67c, Ano67d, Ano68a, Ano69b, Ano69c, Ano69d]. Boric [CP68b, WS66]. Boring [EH61, Neu66]. Borntraeger [Wyr67]. Borntrager [Ban67a, Ban68]. Boston [Aro69]. Botanica [Ano60b, Lew69a]. Botanical [Ano63f]. both [Mca61b]. bottle [CDD62, DW60, Gas66]. Bottom [Lac61, Wel68b, AL68, Bar69, BA64, Bum65, CP68a, CH65, Clu65, Fag68, FA61, Fre61a, FP62, GLN64, LCEP61, MR68, OCO67, San60, Squ69, WW67, Whe67b, Whi62]. bottom-fauna [Fre61a]. bottom-type [AL68]. botttom [Joe64]. Bouma [Gro69]. bound [Bod68, MC65]. Bourcart [Cha66b]. Box [Mac64]. Brachyuran [CB60]. branchialis [SW61]. Branchinecta [Bro69a]. Bransfield [BS61]. Breeding [Bar62, SS68a, CP61]. Brief [Gro65a]. brightwellii [RR69]. Brillouin [SP69]. Brine [PT68]. British [TT66, PG60]. bromo [Top69]. bromo-oxine [Top69]. Brown [Aro69, Bea69, MCA65]. Bryozoan [App66, Den63]. Bubble [Ram62a, Ram62b]. Budget [TR69, Kue61a, LJ69, PWY69, WH67]. budgets [Gor64]. building [VT68]. built [YBF69]. Bull [App66]. Bulletin [Spa66]. Bumpus [PSS61]. Bunyoni [Bea66]. Bureau [Ric69]. Butcher [HT67]. Buttonwood [Lyn66]. Buzzards [Anr64a, Anr64b, San60, Wie60a]. By-laws [Ano63b, Ano65a].

C [Ano66a, Edm66a, Edm68a, Edm69, Koh68, Pai68, Sea65a, Sea65b, Set67, Spa66, Val67, AR67, BW68a, BH67, BB69c, Bug67, Cas62, JS61, LC69, McA61a, Mca61b, Nie65, OM67, Owe63, RM65]. C-ampoules [Nie65]. C-technique [RM65]. Ca [THJ67]. cable [Hug69]. caerulea [Taf60]. calanoid [Car65c, Col61]. Calanus [Hah63, Mul63, RR69]. Calcareous [Bar62, PH66]. calcicola [Sha69]. Calcification [MD67, WW66]. calcite [Bru69]. Calcium [Pyt65, Wey63, KP69, PK67, SS62a, TS67]. Calculation [TS67]. Calibration [Dux63, CD67]. Calif [Edm68a]. California [CC62a, AL66, BG65, BIR64a, BIR64b, BIR65, BW68b, BBH68, Bri67b, CC62b, Gas66, Gro65b, GE62, Hed60, HHSW66, LaF63, Lid66, Lig63, Lon68, vUB63]. californicus [LO62b]. Called [Bar62]. Callianassa [FCJ67]. calorimetry [Pai66]. Cambarus [SS62a]. Campbell [Has68a]. campestris [Bro69a]. Canal [Anr64a, Anr64b, Che68]. Canberra [Car68a]. Canyon [Wri60, Wri61, Wri65]. capability [Bri67b]. capacity [Wat61]. Cape [Fox66, ZTTG65, Anr64a, Anr64b, Wal65a, Wal65b, Wal66]. Carbamino [STH60]. carbohydrate [May60, RC61, Wal65a, Wal65b, Wal66, WD66]. carbohydrateatthe [BW68c]. Carbon [Fou68, Lor68, SB62a, CD67, DS64, Epp68, Jan67, MV64, Mul65b, Mul65c, MSE66, OCV63, RM65, SB61b, SB62b, Str67, SLC68, WG65, WM67a, WL66, Wil69c, Wil61]. Carbon-chlorophyll [SB62a]. Carbon/chlorophyll [Lor68]. Carbonate [Def65, Wey63, BYK69, Cha62, JCG64, Pyt65, WG65, Wan69]. carbonates [Sha69]. Carbonic [CP68b]. carboxylic [STH60]. carboy [McC62]. card [Mar67b]. Careers [Pie64]. Caribbean [Has68b, Hop60]. Carolina [GCV63, Kor68, THJ67, WG60, WM66]. Carotenoid [Man69]. carotenoids [Bro68c]. Castle [BG65]. catch [AC69]. catching [Mun64]. Caulobacter [Jon60a]. cause [Sha60, Wri60]. causes [PBK69]. causing [TG68]. cavity [HB62]. Cayuga [BB65, Ham69b, Lud67]. CCS [Woo65a]. cell [MSE66, OM67, Str67]. cells [Wil60]. cellulose [Kor68]. center [BR64, Ano62h]. central [EA65, Gra66, She63, She64, You61, Mos69]. centric [GC63b]. Centrifugation [Lam67]. cephalus [GW60, Odu68]. ceratium [Nor69]. Ceriodaphnia [Zar69]. certain [STK69]. cesium [Gut65, Wil60]. cf [SB65a]. Chaetoceros [Mul64, SB66b]. Chaetognatha [SS68a]. chain [Par65, TD64]. chain-forming [Par65]. Chairman [Par67a]. Challenge [Och67]. chamber [Lun62]. chambers [Daw60]. Champlain [FH69]. chamrer [FO65]. Change [Sco65, DT65, Liv65, Wey64]. Changes [Ano69e, BH67, CH65, Gil67, Hea67, Ver61, Bam62, Cas60, CK65, DM68, GC63a, Hoe60, Hop60, Mar68b, RV61, SH60, Ver60, Woo61, Yaa64]. Changing [BR61]. Channel [WM66]. Chaoboridae [Rot68]. chaoborus [Nor64, LaR68a, LaR69, Rot68, Sta66, TT66]. Characteristics [BBK69, Wat65a, AAB^{+}65, Jef62a, JSP68, LP65, TP65]. Characterization [TG68, JGST67, Smi68]. Chelan [Whe67b]. Chem [Par67a]. Chemical [AL66, HD69, PPW62, TBLT68, Yaa64, Bee66, BL69, DCL63, HM69, JGST67, RMV61, AH64, Dug68]. chemically [Mac61, TD64]. Chemistry [Par67a, BH69, Par68, RHP^{+}66]. chemostat [Car68b]. Cheng [Koh68]. Chesapeake [BW68c, Car65b, Mul64, Ree69]. Chicago [Dee69]. Chironomidae [IB69]. Chlamydomonas [EM63, TD68]. Chlorella [Lam67, TD64, TD68, Wil60]. chloride [Har67]. chlorinity [Lym69, RC64, TS67]. Chlorophyll [CMD64, GO61, Gor60, Man69, Mos67a, PP63, Sma63, SB65b, Tie68, AC63, And69, BA67, BB67a, BS61, Lor65, Lor67, Lor68, SC68b, SB61b, SB62a, SB62b, Str68, Wat61, WM66, YR63]. Chlorophyll-a [CMD64, GO61, Man69, Mos67a, PP63, Sma63, SB65b]. chlorophylls [Mos67b]. Christian [Car61]. chromatographic [JA67, SLC68]. chrysops [MH61]. Chukchi [BB63, Hun65a]. chute [DCL63]. chydorid [Fre65]. Cie [Cha66b]. ciiaracteristics [PD65]. Cincinnati [WM67a]. cinerea [Bla60]. circulation [Bye65, DR69, GCV63, Rei62]. circulations [Fal69, SMSW69]. Cirriformia [Geo64]. cirripede [BF62]. cis [WS66]. cis-diols [WS66]. Cisco [CB69a]. Cl [THJ67]. cladocera [Fre65, Bur68b, Cow67, MCC69]. cladocerans [RT66]. Clam [Mat60b]. Clams [Rop63]. Clarke [PSS61, San67]. classification [HR66]. clavata [MT64]. clay [But69, Lam62, RK62]. clear [CD64b, Sma63]. Cliffs [Koh67, Ori68, Pai69]. Cliona [Neu66]. closing [Clu65]. closterium [MJ64]. cm [Bur68a]. Co [Aro69, Dee67, Dug68, Lew69a, Par66b, Par66a, Tyl68a, CD67, PH63, Par69, RV61, VBO60, Ver60, Ver61, THJ67]. Coast [PO68, PF69, CS67, EWR65, HM69, HR63, Mur62, OPP64, PPW62, PD65, SC68b, WG60, Man64]. Coastal [Neu66, SS68a, Car65c, Joh64a, MP65, MPSS61]. Coasts [Ber61, And64, And65]. coated [BN68]. Cobalamin [VR61]. Cobalt [FZ67, VT68, PH69, Top69]. Cobalt-60 [FZ67]. Coccolith [WW66, Paa68]. Coccolithineae [WW66]. Coccolithophores [Mar68a, Mar66]. Coccolithophorid [Fou68]. Coccolithus [WW66]. Cod [Anr64a, Anr64b, SW61, ZTTG65, Wal65a, Wal65b, Wal66]. coefficient [Smi68, TH62]. coefficients [Gra66]. Coexistence [San67]. Coleman [Fag69]. coli [Jon67a]. collect [Wil63]. Collection [Art65, Gar65, BS65, CP68a, GB60, Har66b, Joe64]. collections [Joh63a]. Color [Sco65]. Colorado [HO61a, Pen69]. Colorimetric [BC69]. coloring [Hop61b]. Columbia [CS69c, Cus67, DM68, Hob66, OPP64, Par66c, PWY69, PNH66, PG60, Sil64, SR63, TT66, WDH63]. column [Lid66]. Columns [Gol62c]. Comment [PK63, Rod61]. Comments [LC69, Whi60]. Commercial [Ric69]. Commission [Fox66, Ano61e]. Committee [Ano65e, Ano66a, Coc63, SV65]. Common [Bar62, CG62]. Communication [Mar61]. Communities [Mos67a, And65, CS65, Har69, Jef69, JM68, KB69, Pam65, Pat61b, PM65, RV61, Sto67, Ver60]. Community [Con64, CD64a, Fag68, NS62, San60, Tau69, WL69, Wri64]. Company [Och67]. Comparative [Jef64, KB67, Tai60, IJW61, RA66]. comparator [Wol65]. compared [Bea66]. Comparison [BR67, HG69, PTL69, RM65, CS69b, Cow67, Raw60, Str68, TD67, Vin67, Wal69, Wil69c]. comparisons [WM64]. compatible [TD64]. Competition [Par60, Par61, Pra66a]. competitive [Man64]. compiler [Ric69]. complete [Owe69]. Complexes [WS66]. components [Cas62]. Composition [JS66, Bee66, Dic69, Gol62a, HM69, Jef69, Lew69b, MR64, PH66, SHHEL69]. compounds [Hel65, HC69, PMM63]. compressed [DS68]. compressibility [Enr63]. Computations [BA67]. concentrated [McN60]. concentrating [CS63, Den68, SSG69]. Concentratingplankton [DT64]. Concentration [BW68c, Cus67, Top69, FE63, Gil67, IP66, MJ64, WM66]. Concentrations [CG62, AL66, BA67, BRZ68, BY69, BC69, CM69, Dai69, Jan67, RPG69, SFM69, Sil64, Ste63, Str68, VTK62, VT68]. concept [Pla69]. Concepts [Ori68, Par67a, Pai69]. Concerning [Lun62]. Conditions [HT67, Hol68, AL68, Dun67, GS65, Hll67, OM67, Woo65a]. conductance [CW68, Wey64]. conductances [CW68]. conductivity [Smi62]. conductometric [Run65]. conference [Ano61a, Ano67h, Ano69h]. Congeneric [Col61, HS68]. Congress [Ano63f, Ano68d, Ano63e]. congresso [Ano61g]. Conn [Val67]. connaissance [Cha65]. Connecticut [Mar60, MH62]. Connell [Pai67]. consequences [RCBA65]. Conservation [Pie64]. conserve [HL64]. considerations [BM65]. constants [ERM69, KP67]. constituent [HB66]. Construction [AAR69]. Consumption [Bla60, Com68, LO62b, PT68, PB69, PTL69, LZS65]. Content [GO61, SB65b, BBH68, EM66, Epp68, HHH67, Lor65, MP65, May60, MGPB65, MSE66, Ost62, PBK69, Pro62, RC61, Rei62, RK62, Str67, Sut65]. continental [Bum65, Cor67, KC64, Mul64]. Continentales [Edm67]. continued [Far60]. continuous [MJ64, OK69, Pam65, Str68, WH63b]. continuous-flow [Pam65]. continuously [Woo65a]. contributing [SR63]. contribution [SC68b]. control [Cou69, DC62, Gor64, Ste62a]. controlled [Woo65a]. controlling [Cur62]. convenient [Wat65b]. convergence [Man67a]. Cooperation [Ric69]. coordinate [Nor65]. Copepod [CS67, Hah63, Jac61, Lew67, Whe67a, Car65c, FJ66, LO62b, McL65, MB67, She63, She64]. copepoda [Joh63a, Cow67, HS68]. Copepods [Col61, Mul63, AO63, Anr64a, Anr64b, Lan63]. copper [AC67, Jon67a, Wil69b]. copper-treated [Jon67a]. copropel [IB69]. Coprophagy [FS67]. coral [Eme68, Jef68, Joh67a]. corals [Col69, KS62, VT68]. Corbett [TT66]. core [Bur68a, Fel64, Mac69, RK62, Wal67]. Coregonus [CB69a]. corer [BK62, DD69, FK66, Mai69]. cores [Fel64, Liv67]. coring [Bur68a]. cornuta [Zar69]. coronatus [Jac61]. Correction [CC62a, Smi62]. corrections [HC69]. Correlation [Hol68]. cost [PST69]. Costa [Bro65a]. costatum [Cas63a, Cur62, Pra66a, SB66a]. cottid [Mor60]. Coulee [Cas60]. Council [Ano65e]. Counter [MK65, MK68]. counting [Daw60, JS61, Lun62]. counts [AH64]. Countty [Lig63]. County [BIR64a, IB69]. coverage [Gor68]. covered [Col64, LH62, Wri64]. Cowgill [Val67]. crab [BW68b]. Cram [Lew69a]. crane [TPS64]. Crassostrea [Spa66, US69]. Crater [HD69, Bea66, Byr65, KDB68]. crayfish [SS62a]. criticism [LS69]. critique [Fag69]. Cromie [Koh67]. Crop [Bro67, CU65, Hop69, Lyn66]. crops [AM67b, SHHEL69, Wal69]. Crosby [Dav62]. Cross [Che67]. Cross-Stream [Che67]. Crouch [She68]. Crustacea [MCC69, Tib65]. Crustacean [SH67, Bod68, JS66]. crustaceans [Enr63]. Crutchfield [Cha66a]. Cs [WH63b]. Cultivation [Jac61]. Culture [ARL63a, Pra66a, ARL^{+}63b, ACLL64, MJ64, MB67, Sol64, TD64, Uke65, WH63b]. cultures [Joh63b, SB62a]. Culturing [Lew67, SG69]. cumulative [Squ69]. current [Bri67b, KDB68, Lov68a, Ver64, WS61, Che67, Eve67, Hul62, Lon68, Woo61].

Currents [Rei63, Bra69, EHR62, Hug69, LBL60, SB61a]. curves [Wel68a]. Cushing [Dic68]. cyanocobalamin [GC63b]. cycle [BK68, Fre60a, Ham64, KC65, Sol66, WH63a]. Cyclic [GDM64]. cycling [Kuz68]. Cyclococcolithus [Fou68]. Cyclomorpihosis [Haz62]. Cyclops [Cha69, Par60].

D [Ano66a, Aro69, Ban68, Dee67, Dic68, Ise69, Lag68, Lar68b, Och67, Par68, Spa66]. D.C [Bar68, Par67a]. damage [Fra66, TP65]. Daphnia [BR67, Bur69a, Bur69b, Haz62, Haz66, Mar67a, MR65, Par61, Rig61, TD64, TD68, Wri65]. Dark [PTL69, DW60]. darkness [YR63]. Darwin [CP61]. Data [Ano62h, Che67, Gra67, Jac65, KS62, TS67, Wri61]. Davis [Roy68]. Daytime [Wel68b]. Deaths [Ano67e]. decalcification [RR65]. Decapoda [BW68b]. Decapods [RM68]. Decomposition [TV69, RCBA65, SSP68]. Decontamination [McA61a]. Dedicated [Art65]. deduced [EMT65]. Deep [Fou68, Ant66, BRS65, CS69a, Cou68, Dor61, Fra66, HEHM60, LM64, MCK64, RVW65, SHHEL69, TP65, Wey65, Whe67a, Whi60, Och67]. Deep-Living [Fou68]. deep-sea [Ant66, Fra66, HEHM60, LM64, TP65]. defant [FB68]. defined [TD64]. Definition [Val69]. Degradation [PP63, Mos67b]. deGruyter [Lew69a]. Dehydrogenase [HM64, PT68]. Dekker [Par68]. Delineation [Bro65c]. Delta [CDD62]. demersal [VTK66]. denitrification [GD66]. densities [Buz69, TvU64, Zei63, vUB63]. Density [Lam67, Lam62, Bra69, CW68, Kah63, Wri60]. Density-Gradient [Lam67]. density/conductance [CW68]. denticulatum [BR61]. deoxygenation [Bea66]. Deoxyribonucleic [AM68, SS68c]. dependence [Wri60]. dependent [GRB66]. Deposit [Gor66, EWR65]. Deposition [Wey63]. deposits [Mac69, OCO67]. depressor [Ber66, GV66]. Depth [AME67, Gra67, PB69, ARNA64, BCA66, Fee69, Klo63, PK63, Rao68, Ste64, Wan69]. depths [Str68]. derivatives [Gor60]. described [Lun62]. Description [Rei60, ARNA64, Jos66]. desert [GMH67]. design [GV66, TD64]. designed [Fel64]. desmidiaceae [Bro65b]. destruction [GRB66]. Detailed [LaF63, HHSW66]. detection [Orm66]. detectors [PLPP62]. Determination [BYK69, Her62, HH68, HH69, KP67, Lor67, Lov68b, RC64, SPCF65, Sol69, Ste63, AL64, BG64, BW68a, BC69, Cli69, FZ67, HC68a, JS61, KZ68, LC69, Nie65, OCV63, SS69a, WL68, WC62, Wil69c]. determinations [Mac67, Van60]. determined [MGPB65]. determining [CH68, FE63, Hic69, Par65, Run65, SLC68]. detritus [MG66, NS62, Roc63, Whe67a, Wil60]. Deutsche [Ban67a, Ban68]. developed [PM65]. Developing [CB60]. Development [Ori68, McL65, Rag62, ZTTG65]. Developmental [Lew67]. Developments [Edm69]. device [Bur68a, Cas67, CK65, Clu65, CH68, Hil69a, Nut61, VTK66, Wei65, ZMB68]. devices [Nis64]. Diagenesis [Ant66]. diagram [BY69]. diameter [Bur68a]. Diaptomus [Col61, Com65, Com68, HS68, Hea67, San67]. Diatom [RR69, Mul64, Pra65, SB67, Tai60]. diatoms [BW68a, Cas63b, GS67, GC63b, Hol69, Man67a, MW69, SB65a, SB66a, SB66b, WE69, WS62]. dibromo [Top69]. Dictionary [PP67b]. Diego [Lig63]. Diel [Hea67, PTL69]. Differences [HM66, WDH63]. different [MR65, Rig64]. Differential [VBO60, RC64]. Diffusion [OK69, CD64b]. Digging [LP65]. Digital [Gra67]. dilute [But69]. dilution [NA61]. dimensions [HR66]. dimictic [McM69]. Dinoflagellate [TLH62, US69]. dinoflagellates [PR68, SFM69, Sol66]. diols [WS66]. Diptera [Rot68]. Direct [RPA68, RM65]. disc [Gra66, Tyl68b]. discaudatus [Joh67b]. discharge [Cow67]. Discoloration [Rao69]. discrete [ARNA64, BCA66, Str68]. discussion [You61]. disk [SCV69]. dissociation [KP67]. Dissolved [BS63, Ram60, Ram62b, BC69, Car65a, Car65b, CD67, FS69, FE63, HC69, Joh64a, KB69, MB65, MV64, OH68, Par67c, Ram62a, Sha60, Wal65a, Wal65b, Wal66, WD66, WJ67, WM67a, Wil69b, Wil69c]. dissolved-oxygen [FE63]. distances [Lyn66]. distilled [Tyl61b]. Distribution [AC63, Bro67, CB67, Cor67, DS64, FH69, HC67, Hea67, Hil69b, HR63, Lon68, PP63, SS68a, Sma63, TT66, Wan69, Wel68b, vUB63, AC67, And64, Anr64a, Bri67a, Bro65a, Cas63b, Cur62, GS65, Hob67, HM69, JSP68, KB69, KS62, LB62, MH62, Mar66, McM69, MG66, ND66, Nor64, OM67, PH69, Pyt62, Rot68, RH60, SMH65, SS69a, She63, SR63, TeI68, TW66, Tyl61a, VR61, WD66, Wal69, Wil68, YBW64, Bod68, CS66, Man64]. District [Dai69, Her61]. Ditylum [RR69]. Diurnal [Bam62, Dun67, LaR68a, LaR69, Lor63, RMV61, CK65, GDM64, HT69, McM69, Mca63, NDT67, Wel68a, Wal65b]. divalent [Cur62]. diver [Wal67]. diver-operated [Wal67]. diversity [MF66]. dividing [Wat69]. Diving [LM64, Owe69]. Division [Fox66]. DM [Ban67a, Ban67b, Ban68, Fre68]. DM140 [Lew69a]. DM29.50 [Edm66b]. DM5 [Fot68]. Do [PP67b]. Dolichodorus [FH69]. Doliolum [BR61]. Dollar [ST65]. dome [Bro65a, CD64b]. domes [Maz67]. Dominant [Wil66]. Donau [Fre66]. Dorn [FM61]. dragonfly [Kor65]. Dredge [Mac64, Mer62]. Drift [MH62, Bum65, CDD62, Gas66, HG69, Mar67b, Mül65a, PK69, Squ69, Wat69]. drifters [Squ69]. drifting [Mun64]. driven [Bye65, PK69, SB61a]. drought [Hoe60]. dry [RR65, SS63, Yaa64]. dubius [Gei64]. dunaliella [YR63, HT67]. Dunbar [Ori68]. During [TLH62, Gil67, Gol65, Hob66, MJ64, Yaa64]. Dussart [Edm67]. dye [IP66]. dynamic [Dux62, vA65]. Dynamical [Rei60]. Dynamics [Wri64, Dug67, Ril63, Wri65, Dic68]. Dyne [LS69].

Early [Lew67]. east [HM69]. eastern [Bro65a, Gra66, Maz67, Ber61]. Eaux [Edm67]. echinometra [SU66]. echinoplutei [Hsi65]. echo [Nor64]. Ecological [Dav62, Odu68, WK61, BM65, Dee69, Eve67, FFF^{+}66, HHB66, SS68c, VTK66, VR61, LaR69, Ori68]. Ecology [Con64, Joh67a, Mat60b, RAF^{+}62, Sta66, Wei61, Car65c, Cor67, Den63, SB68, SHHEL69, TW66, Pai69]. ecosystem [Joh64a, Tau69]. ecosystems [CD64a]. ectology [Hun65b]. Ed [Bar68, Car68a, Dug68, Edm68a, Edm68b, Edm68c, Fre66, Fre68, Lag68, Pai68, PP67a, Par67a, PP67b, Roy68, Cha66a, Par66b]. eddy [Fal69]. Edinburgh [Ano63f]. Editor [Ano62b]. editorial [Ano69e]. Eds [Fox66, Has68a]. education [Cha66c]. educational [Ric69]. edulis [MD67]. Edward [Bea66]. Efects [Zei63]. Effect [AM68, Bar62, BB69b, But69, CP68b, Cur62, FF67, Gie69, Klo61, PTL69, PM65, Pyt62, Rag62, RR65, Sma63, WS61, BC61, Paa68, PH66, Tyl61a, You61]. Effective [Ano66a]. Effects [CG62, Gor66, Haz62, Haz66, HT67, HO61a, Joh63b, MD67, WW66, WL69, Dic69, HB62, McC66, McC68, MCA65, MH64, PV69, PH62, WW68b, YR63]. efficiency [LP65, Pla69]. Effluent [CMD64, Hob66, TD65]. effluents [CD64a]. Egg [Ans67, MCA65, McL65]. Eggs [CB60, Joh67b, Her63a, Ort62, Taf60, Wil68]. Eh [Whi69]. Eighth [Ano68d]. ejector [FBS68]. electrical [Wey64]. electrode [Dux63, MB65]. Electrokinetograph [Che67]. electronic [MK68]. elegans [SS68a]. element [BBH68, BR65]. Elemental [Gol62a, Sil64]. elements [DWW68, GS65, PH62, PG63, Par66a]. Eleven [Car68a]. Elminius [CP61]. Elsevier [Bro68a, Bro68b, Edm69, Joy68, Lar68b, Tyl68a]. elutriation [LCEP61]. Elvers [Jef60]. emergence [AL68]. emerging [Mun64]. Emerita [BW68b]. emitters [WDH63]. empirical [FB68, HB61]. Encyclopedia [PP67a]. end [Hoe60, RC64]. Endoprocta [Wei61]. Endothermy [Pai66]. Endrick [Hyn68, Hay68]. energetics [KB67, Pat61a, PV68]. Energy [Ano61e, Fox66, Sch65, KB65, Kue61b, MCC69, Pla69]. England [DS64, HR63, Jef60, She68, SW61]. Englewood [Koh67, Ori68, Pai69]. English [Dai69, Gor60, Her61, Nic60]. Enhanced [Man67a]. enigma [CT61]. enlarged [PSS61]. Enrichment [Lew67, Vac69]. entomostraca [SN68]. entrance [PH60]. Enumeration [BB63, McN60, BC60, BC61, SN68]. Environment [Bod68, Wey63, BRZ68, Dug68, Joh65b, Lid65, MCA65, Nri68, PH66, Fox66, Ise69]. environment1 [BM65]. Environmental [Bra68, HM66, HT67, Jac65, Jef62a, Ste62a, Buz69, Lac61]. Environments [Edm68a]. Enzymes [HT67, Doy68]. epifauna [Dri67]. epifaunal [Fag68]. epilimnion [SMSW69]. epipelic [EM66, Hic69]. epipsammic [Hic69]. epoch [Dee67]. equal [HMA65]. equations [Lor67]. equator [Des69]. Equatorial [Hul62, Rei63, Mac67]. equilibria [Def65]. Equilibrium [Par67a]. Equipment [FFF^{+}66, ZMB68]. equivalent [CW68, DW67]. Ergebnisse [Wyr67]. Erie [CH65, Dav64, Hil69b, Ver60]. Erosion [Neu66]. Errata [Ano62c, Gol62b]. Erratum [Ano60c, Ano67f, Ano67g, Wie60b]. erroneous [KS62]. error [AR67, Cas62]. erythraeum [BL65]. Escherichia [Jon67a]. esculentus [SU66]. Esquimalt [Scr60]. Essex [She68]. estimate [KWV66]. Estimates [Enr63, GDM66, GD66, Has68b, Cas62, Col69, MF66, Sut65]. Estimating [Str67, Epp68, Ham69a, Pat62, Wat61]. Estimation [BF63, Mos67a, EM66, Fee69, Mos67b, Ste64, Taf60, Wri61]. Estuaries [Jef62b, Jef60, Bar68, Bro68a, Bro68b]. Estuarine [Whi69, DWW68, Gun61, Jef64, Lan63, MP65, PSG65, TvU64, Tie68]. estuary [BLM63, Buz69, DH64, DR69, HR66, Jef62a, PV68, PMR68, She68, Heu69]. Étude [Edm67]. Euchaeta [Lew67]. euchlora [YR63]. Eucyclops [Par61]. Euglena [Wil60]. euphaushds [Ost62]. Euphausia [Jaw69, PTL69, SH67]. euphausiid [Bod68, JL66, LT65, TC67]. euphausiids [Bri67b, Mau65, OPP64]. euphotic [YBW64]. Europea [Edm68b, Edm68c]. Eutrophic [Hea67]. Eutrophication [BLM63, Bee65, Dav64]. evaluation [Jos66, Par69]. evaluations [Fit69]. Evidence [Dav64, Fra66]. Evolution [Ori68]. exchange [CA66, DT65, DWW68, PH63, PSG65, TK61, Top69]. Excretion [Hel65, Jaw69, MO63, PMM63, HG68, Mar68b, Nal66]. expenditure [MCC69]. Experimental [PTL69, SB65a, SB66a, SB66b, Cas63b]. Experiments [ARL63a, DW60, Hep62, PV68]. exploited [Mar67a]. extent [CS69c]. extinction [Gra66, SC68b, Smi68]. extracts [McC68, Mos67b, OM67]. extratropical [Val66].

F [Bro68a, Bro68b, Par67a, Par68, Roy68, Ano61g]. Factor [Che67, HB61, SWBL63]. Factorial [PV68]. Factors [Cha62, Con66b, Gol62c, Mul63, US69, YBW64, Gol68, SP60]. Failure [Far60]. Fairbridge [PP67a]. Fairy [Bro69a]. fall [Nob66]. Falmouth [Ric69]. farm [Den63]. Faro [TG68]. fashion [DT64]. fast [SCV69]. fast-flow [SCV69]. fate [BB69c]. fatty [Lew69b]. Fauna [Car68a, Hay68, Hyn68, CH65, Fre61a, GLN64, LCEP61, SGMH62]. faunal [SMH65]. Fayetteville [BL69, Bru69, BH69, CB69b, Lud69]. features [EA65, NDT67]. February [Hul62]. Fecal [FCJ67, Dil64, JS66, Sma69]. fecundity [McL65]. fed [Yaa64]. feeders [HMA66]. Feeding [Bur69a, Gor66, MR65, MH61, Mul63, RR69, US69, AO63, Anr64b, Bur68b, Col69, Con66b, MB67, Nor69, Edm66a]. Felsen [Fot68]. Fern [OCO67]. Ferrous [Gje64, McM67, McM69, Sha66]. Ferry [Wri60, Wri61, Wri65]. fertilization [Hep62]. fertilized [HL64]. fertilizer [HL64]. Festival [Par66b]. Field [BBG^{+}62, Fit69, KB63, DG68, Dux63, Gor68, PH63, PPF62, PLPP62, Tai60]. Fields [Par66b]. Figs [Stu69]. filament [Liv67]. Film [Gol62c]. filter [Bur68b, Den68, HMA66, McN60, Moo63, SCV69]. filter-feeding [Bur68b]. Filtering [BR67, Bur69b, Moo63]. filters [McA61a, Orm66, SS69b]. filtration [Mac63]. Fine [Odu68, Jon60b, SS69a]. fine-grained [SS69a]. finer [Mai69]. finger [HW67]. finmarchicus [Hah63]. First [SP60, HEHM60]. First-year [SP60]. Fischer [Edm68b, Edm68c]. Fish [Kat65, Lag68, Ort62, Pau65, GLN64, Her63a, Kah63, RPG69, Roy68, SO61, Sre64b, Sre64a, Wil68]. fish-food [SO61]. fishbite [Fra66, TP65]. Fisher [Bod68]. Fisheries [Cha66a, Ric69, Cur60, Dic68]. Fishery [Spa66]. fishes [Gol62a, Mor60]. fishing [Shi62]. fishponds [Hep62]. fission [BBG^{+}62]. Five [Man69, AH64, Man64]. fixation [DD62, DGR64, GN64, NDDG62, RM65]. fjord [RCBA65]. flagellate [YR63]. flat [Geo64, Joh65b]. flavicans [TT66]. float [AM67a]. Floating [KS62]. flood [Dee67]. floor [EMT65, HEHM60]. Florida [Lyn66, AC63, Bea69, Che67, FAMR60, IJ61, IB69, Wal69, You61]. Flotation [Dil64, SB65a, SB66a, SB66b, WL66]. Flounder [Sco65, Sai61]. flow [CS69c, Kue61b, OK69, Pam65, Pat61b, SCV69, WH63b]. flowering [Pra65]. flowing [Woo65a]. flows [HMA65]. fluctuation [Wal65b]. fluctuations [DG68, Dun67, Hol69]. fluorine [WL68]. fluorometer [NA61]. flushing [Scr60]. flux [DB62, Has68b]. Focke [MCC69]. folding [Ber66]. Folio [Bod68]. following [DH64]. Folsom [LS67, SN68]. Food [App66, Cla62, TD68, BS63, SO61, SU66, TD64]. food-chain [TD64]. foods [MR65]. foot [ARNA64]. Foraminifera [BIR64a, Bra68, Lid66, Phl65]. Foraminiferal [BIR64b, BIR65, Buz69, Lid65, Lyn66]. Forbes [Ric69]. Form [WW66, Sie63]. formation [JCG64, Paa68, Ram62b, Ril63, SEP67]. Formed [TLH62]. forming [Gar65, Par65]. forms [DG67]. Forschungsergebnisse [Ban67a, Ban68]. Forschungsgemeinschaft [Ban67a, Ban68]. forward [TH62]. foundation [Ano60e]. Four [Bur69b, Gil69]. Fourth [Ano61a]. fraction [BFC67, MF64]. Fractionation [And65, Mul65b, Mul65c]. fractions [MC65, Rig64]. fragilis [Fou68]. framework [Bar69]. Fredericella [App66]. Free [Chi60, AL64, Jef69, Rao68, Sol64, WJ67]. freeze [Ano63d, SSG69]. freeze-up [Ano63d]. freezing [MP65, Pro62]. frequency [CC62a, CC62b, Nor64]. Fresh [Fre60b, PH62, Roy68, WS61]. Fresh-Water [Fre60b, PH62, WS61]. Freshwater [DW65, Lag68, Man69, Mos67a, BRZ68, Bro68c, Cas60, Doy68, GMW63, HW65, McN60, Run65, Str64, Tau69]. front [Man67a]. Fuca [SR63]. function [KB63]. Fungal [TV69]. Further [AMP^{+}63, Fra66, RC61, Rig68, SB62b].

G [Bar68, Fre69, Joy68, Ric67, Tyl68a, Val67, Woo63]. gage [Red62]. Gaimard [BR61]. Galapagos [HD69]. Galtsoff [Spa66]. galvanic [Dux63]. Galveston [Phl65]. Gamma [PPF62, PLPP62, WDH63]. Gas [TK61, DT65, SLC68]. gaseous [CD64b]. gases [CD67, Klo61, Ree69, WW67]. Gastropod [Bla60]. gauge [TH69]. Gauthier [Edm67]. Gauthier-Villars [Edm67]. gear [AAB^{+}65]. Gebriider [Ban67a]. Gebrüder [Ban68, Wyr67]. Geiger [JS61]. General [Wal65a]. generation [Fal69]. gentle [DT64]. Genus [Mul63, FJ66, Mul64]. Geochemistry [Gro65a, Joy68, Ano69g, Par66b, Par66a]. geographic [And64]. Geographical [Bod68, Mok60]. geological [HEHM60]. Géologie [Cha66b]. Geology [EMT65, Gro65b, Joy68]. Geomagnetic [Che67]. Geophysics [Joy68]. Georges [CT61]. Georgia [DT65, Kue61b, TK61, TW66]. Gerking [Lag68]. Gerlache [BS61]. German [Fre60a]. Germany [Fot68]. geyser [Vin67]. gibberellins [Joh63b]. Gibbon [Vin67]. Gigantocypris [Tib65]. Glaciers [Ano62d]. Glasgow [Hay68, Hyn68]. glass [HJK63]. Glossary [Nic60]. Glucose [HC68a, VJ66, VJ67, VHJC68]. glucosein [HW65].

Gmelin [Spa66]. gnotobiotic [Tau69]. gold [FM62]. Goldman [Edm68a]. Golterman [Dug68]. Golubić [Fot68]. Gomont [BBK69]. Gonyaulax [GW60, TLH62]. Gould [Par67a]. Goulden [Val67]. gouldii [Gor66]. Government [Ano66a, Spa66]. grab [CP68a, LP65]. gracilis [Wei61]. graded [Pra66b]. Gradient [Lam67, Lam62, Lov68a, TSB63]. grained [SS69a]. Grand [Cas60]. grants [Ano60e]. graphic [Def65]. graphite [MB65]. gravitational [Rao68]. Gravity [Pyt63, Pyt62]. Grays [PH60]. grazing [PV69]. Great [Bee65, Ano61a, Dea63, DD66, FJ66, TS67]. Green [BL69, Bru69, BH69, CB69b, Lud69, BL65, BB69c, OC69, Sha69, WH63b]. Gross [Cas63a, KN69, McC65]. grossa [Rop63]. Group [Ano65e]. groups [Bee66, Str64]. Growing [RR69]. Grown [HM66]. Growth [Bar62, Cha66c, CG62, Haz66, Jan67, Jon67a, Sut65, WW66, ARL^{+}63b, CP61, Far60, GK66, Ham69b, McC65, MH64, Mos69, MB67, Par65, PR68, Ram62a, SP60, SHHEL69]. Gulf [BK68, Dux62, FV65, Gro65b, Haq67, Has68b, KC65, Pra66b, Rei60, SS68a, Zei64]. Gunner [Hah63]. Guppy [CG62]. Gynandromorphism [Fre65].

H [Bar68, Car68a, Dic68, Dug68, Gro69, Has68a, Lar68b, Lew69a, Och67, Pai67]. habit [AO63]. Habitat [Bro66, Bis68, Hed60, Dug68]. habitats [Bam62]. Half [ERM69]. Half-saturation [ERM69]. Hall [Koh67, Ori68, Pai69]. halocline [BW68c]. Hamburg [Ban67b, Lew69a]. Hamilton [Hyn69, JM68]. Handbook [Roy68]. Handbooks [Lew69a]. handling [Ano62f]. Hanford [PNH66]. Hansen [Jaw69, SH67]. Harbor [SGMH62, Gor66, PH60]. Hard [Rop63]. Hardbound [Pai69]. Harper [Ray68]. Harvest [Ray68]. Harveys [RT66]. hatch [Far60]. Hatching [Joh67b]. Hawaiian [HZ62]. Heat [DB62, WH67, Gor64, Has68b, HT69, Jon67a, LJ69, PBK69, RL64, Sho64]. heat- [Jon67a]. heating [BG65]. heavy [MF64]. helgolandicus [RR69]. helicopter [Ell66]. herbivore [McC65]. Herbivorous [TD65]. heterocephalus [FH69]. heterocyst [OC69]. heterotrophic [But69, VJ66]. Hidaka [Art65]. Higginson [Ise69]. High [DGR64, Joh63a, Nor64, VTK62, BBH68]. high-frequency [Nor64]. Hippidae [BW68b]. History [Val67, Liv65, Dee67]. Holding [CMD64, HMA65, TD65]. Hole [Ric69, WK61]. Holland [Dug68]. Homarus [HM62]. Home [Daw60]. Home-made [Daw60]. homogenates [BF63]. Hook [ZTTG65]. Hopkins [Stu69, Wyr66]. HOPPE [Lew69a]. horizontal [HS69, Joe64, PLH68]. Hosts [Koh68]. House [Set67]. Howell [GW60]. Hrbácek [Fre68]. humic [PR68]. Huron [Rod61, Bri67a, PA60]. hurricanes [You61]. Hutchinson [Val67, Fre69]. huxleyi [WW66]. hyalinum [SB66b]. Hydrobiological [Fre68]. Hydrobiology [Nic60]. hydrogen [Cli69, WC62]. hydrogeochemistry [Ano69g]. Hydrographic [Cou69]. hydrography [Anr64a, Anr64b, IJ61, MH62, RHP^{+}66, Zei64, Anr64a]. Hydrohiologie [Edm66b]. Hydrological [GC63a]. Hydrophylita [Dav62]. hydrophytes [HB66]. Hydrosphere [Par66b]. Hydrostatic [AM68, HM64, Klo61, ZB65].

Hydroxylamine [FSS67]. hydroxyquinoline [Top69]. Hynes [Fag69]. Hyperion [BIR65]. Hypolimnetic [BG65]. hypolimnion [Cou68, LBL60]. hypothesis [Rig68].

IBP [Roy68]. ice [Col64, Dee67, Hll67, LH62, Nut61, RR67, Whi60, Woo65b, Wri64]. ice-covered [Col64, LH62, Wri64]. ICITA [Jos66]. identification [Dug67, Par66c]. II [Fre69, Wyr67, Anr64b, ARL^{+}63b, Bra69, Bru69, DD62, GLN64, LaR69, Man64, PB69, SB66a, Wal65b, Wal69, Wie60a]. iii [Spa66, ARL63a, CS66, DC62, DD65, Lud69, Mar68b, San60, SB66b, Wal66, Wri60]. Ill [Dee69]. Illies [Edm68b, Edm68c]. Illumination [Haz66]. illus [Edm67]. im [Fre60a]. implicata [DW65]. Implications [MD67]. important [HHSW66]. impoundment [McC68]. Impregnation [HJK63]. Improved [CD67, Pat62, Ste62b, Ree67, Wei65]. Improvements [BCA66]. in-situ [ST61]. Inadequacy [TD68]. Inc. [Ano60f, Ano61d, Ano63a, Ano63b, Ano64c, Ano65d, Ano66f, Ano67i, Ano68c, Ano69i, Lau62, Lau63]. Incidence [Rop63, SW61]. include [BY69]. inconsistent [Rig68]. incorporation [Ano63a]. increased [ZB65]. incubation [BH67, Epp68]. incubator [Owe63]. incubators [DS68, TSB63]. Index [Ano60d, Ano61b, Ano62e, Ano63c, Ano64b, Ano65c, Ano66b, Ano68b, Ano69f, Man67b, Nor65]. India [Sre64a]. Indian [BRS65, JSP68, Mul65b, Mul65c, RHP^{+}66]. Indiana [Sta66]. Indicator [SW61, Hop61b, Van60]. indicators [Bro65b, CS67, PPW62]. indium [BB69a]. individual [HMA65, PV68]. induced [SH60]. inexpensive [BK62, Car68b, HAM68]. infaunal [Joh65b]. Influence [Anr64a, Anr64b, EHR62, Joh65a, PR68, RPG69, US69, AC69, GS65, Kur66, McM67, OC69]. influences [Hob66]. influencing [Cha62]. Information [Ano62f, Fox66]. ingested [Bur68b]. inhibition [GMW63]. Initial [Ver64, Har66b]. injection [CD67]. injury [GMW63]. Inland [Car68a]. Inlet [GG64]. inlets [PG60]. inner [Mul64]. Inorganic [Rig61, BFC67, DD65, Gil67, Ham64, Rig64]. INQUA [Ano63e, Ano68d]. insect [HG69, TD65, Vin67]. insecticides [GLN64, NGL^{+}64]. insects [BB68, Mun64]. inshore [Hop61b, KB69]. Institute [Ano60a, Ise69, Car65b]. Institution [Ric69]. Instrument [PF68, TH62]. instrumentation [PST69]. Integrated [TM64, Fee69]. intensity [BB69b, CL62, MJ64, Mca61b, Paa68]. Interaction [FS69, CP61, DB65, Man64]. interactions [MJ64]. Intercalibration [DJKMS65]. interface [Bro69b, FE63, Jar67]. Intermediate [Wyr66]. internal [AL66, BR60, CC62a, CC62b, HB66, Lee61]. International [Ano62g, Ano63f, Ano66c, Ano67h, Ano69g, Ano69h, Ano61g, Ano61c]. internationalis [Ano69j]. Interne [Wyr67]. Interpretation [Jac65]. Interspecies [WDH63]. interstitial [Har67, Joh67c, Ree67, BH69]. Intertidal [GO61, Gor66, Mat60a, Ell66, Geo64, PH66, PMR68, SGMH62]. intertropical [Stu69]. Introduction [Fre69, Edm69]. invertebrate [HMA66, Wat69]. Invertebrates [Chi60, Fag68, Hil69a, RR65, SS61, Sug67]. inverted [Daw60]. investigation [HS69, Par69]. Investigations [Mul64, STK69, Bie62, Eve67, RC61]. iodimetric [Van60]. iodometric [Van60]. ion [KP69]. ion-pairs [KP69]. ionic [BY69]. ions [Mil69]. Iowa [Sma63]. Irish [Hug69, MT64]. Iron [SU66, Gje64, McM67, McM69, MS62b, Sha66]. irradiance [IJW61]. irradians [Mar60]. irrorata [MO63]. Isaacs [ARNA64]. Isabela [HD69]. Island [KC64, VR61, Wey63, Wil68, DD61]. Islands [BL65]. Isolated [Gol62c, HM64, JSP68, RAF^{+}62]. isolating [Sol64]. Isolation [Doy68, FAMR60, Val63]. Isopod [EH61]. isotope [Ano69g]. isotopes [Ant66, DS64]. isotopic [TBLT68]. Issued [Ban67a, Ban68]. Itasca [CU65]. IV [ACLL64, BH69, DCL63, Wri61]. iylococcal [SWBL63].

J [Bod68, Bro68a, Bro68b, Cha66a, Cha66b, Edm66b, Edm68b, Edm68c, Fox66, Fre68, Ise69, Koh67, Lew69a, Ori68, Pai67, Pai68, Pai69, Ray68, Wyr66]. January [Hob66]. Japanese [PP67b, TeI68]. japonica [Lew67, Rao69]. Jerlov [Tyl68a]. Jersey [Koh67, Och67]. Joaquin [Heu69]. John [Fre69, Gro69]. Johns [Stu69, Wyr66]. Jolla [HWE67]. Jørgensen [Edm66a]. Journal [Gor68]. journals [Ano61f]. Jr [Edm69, Och67, Wyr66]. Juan [SR63]. Juday [Ege64]. July [CB69a, Rei63].

Kcs [Fre68]. keg [YBF69]. kelp [vUB63]. Ken [Par66b]. Kidd [ARNA64]. Kinds [TM64]. kindtii [MCC69]. Kinetic [Hsi65]. kinetics [HW65, VJ67]. Knudsen [Bar69]. Koji [Art65]. Kormondy [Pai69]. Kosmos [Edm66b]. Koyama [Par66b]. Krauss [Wyr67]. Krøyer [Car65c].

L [Bod68, Dug68, Ric69, Woo63, Wyr66]. L. [MT64, Moh61, SW61]. labeled [BW68a, MR65]. Laboratory [Jac61, MB67, Ric69, Fit69, LT65, MW69, MGPW64, MS66, PM65, SG69, Tai60]. Labrador [Car65c]. Lack [PTL69]. lactic [BF63]. lactuca [Gut61, Ver62]. lacustrine [IB69]. Lagoon [Scr60]. lagoons [EM63]. Laguna [Val67, BIR64b]. Lake [AH64, BG65, Bea66, Bea69, Bri67a, BL69, Bru69, BH69, BR67, Byr65, CB69a, CB69b, Dai69, DD62, Hea67, Her61, HD69, IB69, KDB68, Lud67, Lud69, OCO67, Rod61, San67, Sma63, Sta66, TT66, Wal66, Bro65b, BC63, Car65c, Col64, Dic69, Fre69, GD66, GS65, Har67, Hll67, Kur66, LH62, Mai66, McM67, McM69, Ram60, Rig68, Rot68, Ste63, Ste64, SH60, TBLT68, Tal63, Web61, Wri64, WH65, AAT64, BB65, CH65, CU65, Cou68, Dav64, DB62, EL68, FH69, GC63a, GPE69, HB61, Ham69b, Hil69b, Hob61, Hol68, Hol69, JM68, LBL60, Liv65, MH61, MGPB65, Nob66, Nri68, PA60, Red61, RT66, RCBA65, RA66, SB61a, SB67, TG68, Ver64, Ver60, Wel68b, Whe67b]. Lakes [BPTW65, BYK68, SP60, AH64, Bea66, BBH68, BR60, BR64, Bye65, Cas60, DD69, DS64, DD61, DD62, DD65, EA65, Fre60a, GMH67, GN64, Gor60, Gor64, HW67, Kuz68, LJ69, NDDG62, PH69, Pen69, Rag62, RL64, Raw60, Rig64, Roc63, Sre64a, TeI68, You61, Ano61a, Bea66, Bee65, Dea63, FJ66]. laminated [Bru69, Lud69]. lampa [Neu66]. lancelets [Cor67]. landlocked [Scr60]. Langmuir [Fal69, SMSW69]. large [AME67, AMP^{+}63, MPSS61, MR68, Raw60, ST61, Wat69]. large-volume [AME67, AMP^{+}63, MPSS61]. larva [Far60]. Larvae [LaR68a, LaR69, Lon68, TT66, US69, Her63a, Nor64, Ort62]. larval [IB69]. latent [Has68b]. lateral [Lyn66]. lauderi [SB66b]. Lauff [Bar68]. Lawrence [Bee65]. laws [Ano63b, Ano65a]. layer [Bro65c, TC67]. layers [PLH68, Sho64]. Lebistes [CG62]. Leidy [Wei61]. length [BH67]. Leptodora [MCC69]. leptopus [Hea67]. Lernaeocera [SW61]. Leucothrix [Bro66]. Levandowsky [Pat69]. level [Ano66h, DC62, Liv65]. levels [SHHEL69]. Levring [Lew69a]. liberation [MC65]. Liepolt [Fre66]. Light [Ano65e, DG68, DW60, GMW63, MS61, Nes67, PTL69, AAR69, BB69b, CL62, MJ64, Mad66, McM67, Mca61b, Paa68, SC68b, Spi68, Tyl61a, Yen62]. Light-Dark [PTL69]. lightweight [BK62]. Ligurian [Kan67]. Lilljeborg [Com65]. lime [SH60]. lime-induced [SH60]. Limitation [Mos69, Dug67, Ham69b]. Limited [BR61]. Limiting [Gol62c, Jan67]. limits [Mor60]. limnetic [Rot68]. Limnofauna [Edm68b, Edm68c]. Limnol [CC62a]. limnologiae [Ano69j]. Limnological [AH62, Dea63, McC68, Sre64a, AM67a, HO61b, MS66, Raw60, Wri61]. Limnologie [Edm67, Fre66]. limnologist [Val69]. Limnology [Ano60f, Ano61d, Ano63a, Ano63b, Ano64c, Ano65d, Ano66f, Ano67i, Ano68c, Ano69i, DC62, DCL63, Hay64, Lau62, Lau63, TM64, AAT64, Ano61g, BL69, Fre69, Gor68, Sre64b, Wri60, Wri61, Ano60d, Ano61b, Ano62c, Ano62e, Ano63c, Ano64b, Ano65c, Ano66b, Ano68b, Ano69f, Gol62b, Man67b]. limnoplankton [Fre69]. Limnoria [EH61]. line [Shi62]. lines [Fra66, TP65]. Linné [Bro67, HC67]. linsley [LB62, DS64]. liquid [LC69]. list [Ano60h, Ric69]. liter [Jit64, YBF69]. litter [McC68]. Little [Aro69, WM67a]. Littoral [FH69, Cas63b, Cha66b, Mok60, Pen62]. littorina [MO63]. live [Sug67]. Living [Chi60, Fou68, Koh67, Geo64, OM67]. Lobster [HM62]. Local [DM68]. Loch [Hay68, Hyn68]. Logical [Jac65]. Lomond [Hay68, Hyn68]. London [Fre69, Joy68, Koh68]. long [Liv67, Shi62, KC64, VR61, Wil68]. long-line [Shi62]. longa [Haq67]. longerostris [SS62a]. longulus [SS62a]. Looseleaf [Bod68]. loss [Kor65, Kor68]. lotic [BRZ68, WDS64]. Lovén [TV69]. Low [Joh67b, BC69, CM69, Har66a, PST69, Ste63, SHHEL69]. low-cost [PST69]. lower [Mul64, Cas60, CS69c]. lubs [MR68]. lucens [Haq67]. lucunter [SU66]. lunula [Nor69]. luteus [Pra66a]. lutheri [US69]. lydia [Kor65]. Lynch [Bro69a].

M [Bod68, Ori68, Stu69, Val67, Woo63, BPTW65, LP65, PB69]. M. [Haq67]. MacArthur [Pai67]. mackerel [MG62]. Mackinac [PA60]. Macmillan [Owe69, Lar68b]. macrobenthos [RA66, WM64]. macrogynus [HM64]. Macroinvertebrate [JM68]. macroinvertebrates [GB66]. macromolecular [SWBL63]. macromolecules [BB67a]. macrophytes [GB66, Kor68]. macrophytoplankton [Des69]. made [Daw60]. Madison [Dic68, Gil69, WH67, WM67b]. Madras [Sre64a]. Maggiore [MGPB65]. magna [MR65, Rig61]. Magnesium [KP68, KP69]. main [Str64]. Maine [Haq67, Mai66, SS68a]. Mainea [KC65]. mainland [Lid66]. Maintenance [LT65]. Maitland [Hyn68]. major [Bee66, BBH68, WS62, Wil66, FCJ67]. Malacobdella [Rop63]. maldanid [Man64]. Management [Cha66a]. maneuvers [Pfa62]. Manganese [PO68, MGPB65, Orm66, RHMB62, SU66]. Manganese-54 [PO68]. manual [Ano61f, Owe69]. mapping [SFM69]. maps [Fox66]. Maracaibo [Red61]. Marcel [Par68]. Margalef [Dee69]. Margaree [Hun65b]. Marina [Lew69a, Ano60b]. Marine [AM68, Ber61, Bod68, Bro66, CB67, GS67, Gro65b, Hah63, Hol62, Jac61, Kes60, Lew67, Mul63, Nes67, Pai68, PP63, Ric69, RR69, SH67, TLH62, Wat65a, Ano60g, AO63, BO60, BBG^{+}62, BB63, BC60, BC61, BM65, CM69, Cas63b, Cha66b, Che68, CS65, Dil64, DJKMS65, Enr63, ERM69, ECS69, FS69, Fel64, FS67, FM62, Gol65, GG64, GC63b, Gut63, HHH67, Hel65, Hob66, Hop60, Jan67, JS61, Joh64a, Joh64b, Joh65a, JS66, JSP68, Joh63b, Jon60b, KB69, Koh68, LO62a, Lew69b, McA61a, MW69, Mok60, Moo63, MH64, MB67, PR68, RC61, RAF^{+}62, SWBL63, SSP68, SFM69, SS69b, SB65a, SB66a, SB66b, SS61, Ste63, SCV69, ST61, Tai60, TvU64, Tea60, Uke65, Val66, VTK66, Vis60, WDH63, WJ67, WK61, YR63, ZB65, Koh68, Lew69a]. Marine [Par68, Cha66b]. marinus [AM68, HM66]. Marion [IB69]. marking [BB68]. marl [CB69b]. marsh [Bra68, DT65, Kue61b, Phl65, TK61, TW66, WK61, Yaa64]. Martha [WM64]. Martin [Par68]. Marukawa [Lew67]. Maruzen [Par66b, Par66a]. Maryland [Ise69, Wyr66]. Mass [ARL63a, Joh66, Ric69, ARL^{+}63b, ACLL64, Uke65]. Massachusetts [Anr64a, Anr64b, Lid65, Rop63, SGMH62, SMH65, Wal66, WK61, ZTTG65]. masses [Hop61b]. Masson [Cha66b]. material [Jar67, JGST67, PH66]. Materials [PP63, Gar65, Mun64]. Math [Dav62]. mathematical [Ril65]. mats [BB67a]. Matter [CMD64, SB65b, BS63, BFC67, Con66a, Con66b, Fre60a, Geo64, Hob67, HM69, Hop61b, KB69, MR64, OH68, Par67c, RCBA65, TeI68, Wil69b, Wil69a, Yen62]. Maucha [BY69]. Mauchline [Bod68]. maximum [And69, Bur68b, CS69a, Lor65]. maximun [Bro65c]. Mayfly [DC62]. McIntyre [CP68a]. McMurdo [AH62]. Md [Stu69]. mean [Ano66h]. means [Bro65c, Smi68]. measure [CD64b, GK66]. Measurement [BN68, LJ69, Mca63, SS68c, Wil61, Yen62, BB67a, GB60, HHB66, IP66, KEB62, KB63, McM67, MV64, MC65, NDDG62, PH63, RM65, Sha66, Wel68a]. Measurements [Ano65e, CK65, Jac65, MPSS61, Neu66, PB69, PKB66, AMP^{+}63, Car66, HB62, Hug69, KN69, Lov68b, NA61, PPF62, Sch65, Sma69, Smi62, VTK62]. measures [Rig68]. Measuring [PF68, AR67, Eme61, McA61a, MB65, Nie65, Pam65, SS62b, TO66, TH62]. mechanical [HO61b]. Mechanism [Gut61, Nis64, SMSW69, Ste62b].

Mechanisms [San67]. Mediterranean [BG69, Fou68, Fra66]. medium [TD64]. medusae [MH62]. Meereshöhlen [Ban67b]. meeting [Ano62g]. Meganyctiphanes [Bod68]. meiobenthos [Cou69, WM64]. meiofauna [Wie60a]. Mellita [ST65]. Melosira [Hol68]. Melt [Woo65b]. Mem [Val67]. members [Bar68]. membrane [McN60, Orm66]. membranefilters [CS63]. membranes [HL64]. Menai [BFC67]. Mendota [Nri68, DB62, HB61, LBL60, MH61, SB61a]. Mercenaria [Ans67]. Meridian [Dee67]. meromictic [Bru69, Car65c, CB69b, TBLT68]. mersbrane [Moo63]. Messina [TG68]. Metabolism [PF68, CD64a, Cor65, DD62, DD65, NDDG62, Nri68, Pam65, PM65, SH60, Wel68a, WL69]. metalimnetic [Sha60]. Meteor [Ban67a, Ban68]. meteorological [Dea63]. meter [AAR69, ARNA64, CL62, Jos66, Mad66]. meters [CK65]. Methane [HB66]. Method [CS63, CB60, Gol62c, OCV63, Sol69, AR67, BW68a, BC60, Car65a, Car65b, Dum69, Eme61, Epp68, Fag69, FB68, FZ67, FE63, FV65, GB60, Har66b, HC68b, JA67, KN69, LCEP61, McA61a, MW69, MR66, Moo63, Mos67b, Orm66, Pat62, Rig68, Run65, SSG69, Str64, SLC68, TO66, Uke65, Wat61, Woo62]. Methoden [Edm66b, Wyr67]. Methods [Hic69, BC61, CS69b, FF67, Mac67, Mca61b, Par69, Wil69c, Gro69, Joy68, Roy68, Par68]. Metridia [Haq67]. Mexico [Dux62, Has68b]. Miami [WM67a, Che67]. Michigan [CB69a, Hol68, Hol69, Nob66, RA66, Rot68, Ver64, Wel68b]. microbial [HH69, Vac69]. microbiological [Gol64]. Microbiology [Rit65, Roc63, Ano60g, Bro68b]. microconstituents [STK69]. microcosm [Nix69]. microcosms [McC62, McC65]. microcrustacean [PH62]. Microdistribution [Cas62]. microgram [BG64]. Microlayer [Har66b]. Microorganism [Bro66]. microorganisms [Fre60a, Kuz68, LO62a, LO62b]. microscope [Daw60]. microscopic [SS69a]. mid [ARNA64]. mid-water [ARNA64]. Middle [Bum65, DC62, DCL63, Jef60]. Midget [Eve67]. Midwater [Wis62, Ber66]. Migrating [SH67, Bis68]. Migration [EH61, Her63b, TT66, Bri67b, Nor64]. Mikroorganismen [Fre60a]. Miller [Set67]. Mineral [HC67, WS61]. mineralogy [Cha62]. Minimum [GC63b, Sha60, TC67]. Minnesota [CU65, SP60]. minor [Lew69b]. minutus [Car65c]. Mississippi [DC62, DCL63, CDD62]. Missouri [Cow67]. Mitchill [SP60]. Mitochondria [HM64]. mixed [Joh63b, PV68]. mixing [RH60, SMSW69]. Miyake [Par66b, Par66a]. Mn [WG65]. mobile [MS66]. Model [SS62a, Fee69, LS69, Ril65, SG69, Tau69]. modem [Hed60]. Moderate [BPTW65]. modern [Ori68]. modestus [CP61]. modification [BY69, CP68a, LM64, TH69]. Modified [FM61, MF64, SN68, Wel68a]. molal [Mil69]. mollusc [Joh66]. molluscan [JCG64, Val66]. Molluscs [Koh68, Gil69, MGPB65]. mollusk [PG63]. molted [SS62a]. Molting [JL66]. molybdenum [BG64, FM62, KZ68, Rig68]. monilata [GW60]. monitor [BRZ68]. Monochrysis [US69]. monomixis [CWB67]. monoxide [SLC68]. Montagu [Geo64]. Montana [Wri65]. Monterey [Squ69]. Monthly [Dai69]. moored [DB65]. mooring [Fra66, PST69, TP65]. morphological [Dav62]. Morphometric [Gor64]. Morphometrics [Bro67].

Morphometry [Byr65]. Mortalities [CB69a]. mortality [Joh66]. motile [Wil63]. motor [DS68]. Mount [Sto67]. mountain [Pen69]. Mountains [Web61]. mouth [AO63, DM68, WDH63]. mouth-parts [AO63]. Movements [LH62, BRS65, CS67, Sai61]. mucor [Bro66]. Mud [BR64, FE63, Geo64, LB62, Bea69, IB69]. Mud-center [BR64]. muds [DT65, Gor60, Tea60]. Mugil [GW60, Odu68]. Mullet [Odu68]. Mullin [WC62]. multiple [FK66]. mundana [EM63]. municipal [NGL^{+}64]. Munida [RM68]. Murdoch [Jaw69]. Mussel [DW65, BRZ68, Kue61a, Kue61b]. Mya [Mat60a, Mat60b]. Myers [Sta66]. Mysis [RPA68, TV69]. Mytilus [MD67].

N [Fox66, Ori68, Pai69, Tyl68a, NDDG62, Wat65a]. Nägele [Fre66]. naiads [SC68a]. nana [SB65a]. nannoplankton [Lun62]. Nantucket [Lid65, Rop63]. Narragansett [Her63a, Mar65, Mar68b, Pra65, Pra66a]. National [Ano60e, Ano65e, Ano66d, Car68a, Gil69, Sto67, Vin67, WH67, WM67b, Ano66g, Ano62h]. Natural [BG69, Hop61b, Pie64, Sol69, BG64, BH67, Bis68, BB69b, Car66, Cli69, CD64b, CS65, DS64, Epp68, Fee69, GC68, HC68a, HC69, KEB62, Lor63, MK68, NA61, OM67, Par67a, Sha66, Smi68, Tyl61b, VJ67, Vac69, Ver61]. Naturally [BBK69]. Nature [LaR68a, BB69c, SWBL63]. naval [Ise69]. navigation [DC62]. near [Clu65, FE63, IJW61, Joe64, Ram62a, Ram62b, WDH63, WK61]. near-shore [Ram62b]. Negentropy [Pat61b]. Nelson [Car61]. Nematoda [FH69]. Nematodes [Chi60, Mur62, Tea60, TW66, WK61]. Neomysis [Her63b, Heu69, Jaw69]. Nephtys [Cla62]. neritic [NDT67]. net [ARNA64, AC69, Jos66, MF66, PMR68, Str64, WW68b]. nets [FC65, HG69]. netting [Wil63]. Neuse [THJ67]. Newport [Lid66]. Niantic [Mar60, MH62]. niches [Lac61]. nickel [SU66]. Nicol [Pai68]. Nikolassee [Ban67a, Ban68]. Nitinat [RCBA65]. Nitrate [CS69a, HZ62, VTK62, ZB65, ERM69, ECS69, GDM64, Her61, WC62]. Nitrification [CM69]. Nitrifying [Wat65a]. nitrite [CS69a, Pro62, VTK62]. Nitrogen [DD62, DD65, GDM66, GN64, Jaw69, Man69, NDDG62, DD61, DGR64, DG67, Fit69, Ham64, HH68, MP65, OC69, OH68]. nitrogen-rich [OH68]. Nitrogenous [Jaw69, Tho66]. nitroso [FZ67]. nitroso-r [FZ67]. Nitrosocystis [Wat65a]. nitrosyl [Jon60b]. nitzschia [MJ64, SB65a]. Nkugute [Bea66]. No [Ban67a, Ban68, Bar68, Edm69, Fre66, Roy68, Stu69, Wyr66]. nondestructive [CD67]. nonmembers [Bar68]. nonsubscribers [Ban68]. NORPAC [Coc63]. North [Dug68, GCV63, THJ67, WM66, Gra66, PK67, RWV64, She63, Wyr65, AH64, Gei64, Hay64, Hul62, Man64, MCK64, MR64, PO68, SB65b, WG60]. northeast [And69, CS66, Hob67, Mac67, WG60]. northeastern [Tho66, Val66]. Northern [SB65b, Car65c, Gas66, Raw60]. northwest [Mar68a]. northwestern [Mar66]. norvegica [Bod68]. Nostrand [Och67].

Note [Mos67a, Ano62b, Hll67, Lor65, Rao68, Van60, You61]. Notes [BS65, Col61, DW65, LZS65, Wat61, ARNA64, Dav62]. Notice [SV65]. Nova [DT65, RK62]. November [BK68]. Ntroduction [Sch62]. Nuclear [Par66b, TLH62]. numbers [EM66]. Numerical [Val66, Fee69]. Nutrient [Dug67, Gol62c, Ham69b, Par67b, ACLL64, AL66, GK66, Joh65a, MJ64, SR63, Str68, SHHEL69]. nutrients [Bro65a, But69, DM68, Par67b, Tho66]. Nutritional [Haq67, TD68]. nutritive [JS66].

O [Has68a, Lew69a, Mca61b, RV61, Ver60]. Oak [Fox66]. Obermiller [Fre66]. obligate [MH64]. observation [Dea63, FO65, Gol65]. Observations [Cla62, Coc63, Den63, Fou68, HM62, Joh67b, Lee61, Mar66, Mat60b, Mca61b, Neu66, Nob66, Ree69, RPA68, ST65, SSP68, SS68a, Spi68, Squ69, Sto67, Tib65, Eme68, Gra66, HT69, HEHM60, Hun65a, JL66, KDB68, LBL60, PG60, Rig68, Sho64, SB65a, SB66a, SB66b, Wri60]. Obtained [HM64]. Occurrence [MS62a, MS62b, MS62c, TvU64, SEP67, She64, VHJC68, NGL^{+}64]. Occurrences [Col61, Jef60]. Occurring [BBK69, Gil67]. Ocean [BYK68, HHB66, HH69, HT69, IP66, Klo63, Mca63, MG66, Pyt65, Sho64, SS68c, WG65, Wan69, Yen62, And69, Bro65a, BB67b, CS66, Des69, GDM66, Gra66, Hob67, HEHM60, JSP68, Mac67, Maz67, Mul65b, Mul65c, PK67, Rei62, RWV64, RHP^{+}66, Stu69, Tho66, Tib65, VT68, Wal69, Wey65, Wyr65, Wyr66]. Oceanic [Ano66e, Nor65, Par69, PBK69, PF69]. oceaniques [Cha65]. Oceanogr [CC62a]. Oceanographic [Ano62h, Coc63, Ric69, Stu69, Wyr66, Cha66c, Nis64, PST69, TD67]. Oceanography [Ano60f, Ano61d, Ano63a, Ano63b, Ano64c, Ano65d, Ano65e, Ano66a, Ano66f, Ano67i, Ano68c, Ano69i, Art65, Cur60, Lau62, Lau63, FBS68, Hor62, PP67a, Ric69, Tyl68a, Ano60d, Ano61b, Ano62c, Ano62e, Ano63c, Ano64b, Ano65c, Ano66b, Ano68b, Ano69f, Gol62b, Man67b]. Oceans [Bro68a, Bro68b, Bro65c, BRS65, PP67b, Aro69]. oceanus [Wat65a]. October [BK68]. Off [Hed60, And64, And65, AL66, Bee66, CS69a, CC62a, CC62b, CDD62, Che67, CS67, DM68, EWR65, HM69, HHSW66, LaF63, Lid66, MS62b, Mul64, OPP64, Par67b, PBK69, PF69, Rao69, RK62, SC68b, SR63, WD66, WG60, Wel68b, vUB63]. Office [Ano66a, Spa66]. offshore [WM64]. Ohio [WM67a]. Oil [CMD64, CD64a, McC66, TD65]. old [AM67b]. oligochaetes [Bri67a, Hil69b]. Olisthodiscus [Pra66a]. one [ARNA64, Jos66]. one-meter [ARNA64, Jos66]. Ontario [JM68, San67]. Onthe [Ste64]. open [WG65]. operated [Wal67]. Operation [Ano63d]. Operational [Whi69]. Ophiophteus [Gei64]. Opossum [Her63b]. Opportunities [Pie64]. optical [AH64, Smi68, Wol65, Tyl68a]. Optimal [HM66]. orange [Mer62]. orange-peel [Mer62]. orcanisms [WDS64]. order [Cor67]. Oregon [And64, And65, Byr65, CS67, KDB68, Mur62, OPP64, PPW62, Par67b, PD65, PBK69, PF69, SC68b]. Organic [CMD64, Fou68, Geo64, Kan67, Man69, MCK64, Ril63, RWV64, RVW65, SB65b, BS63, But69, Con66a, Con66b, CD67, Doy68, GC68, Hel65, HC69, HM69, HH68, Jan67, Joh64a, Joh67a, KB69, McC68, MR64, MV64, MG66, Mul65b, Mul65c, OCV63, OH68, Par67c, PMM63, RCBA65, RHP^{+}66, Str67, TeI68, Vac69, WG65, WM67a, WS66, Wil69b, Wil69c, Wil69a, Wil61, WH65]. organically [MC65]. organism [Den68]. Organisms [CB67, Mac64, BBG^{+}62, Dil64, FM62, GDM64, Gun61, Har68, Jon60b, SO61, Sol64, Sut65, TSB63, WDH63, Ric67]. Orientation [RM68]. oriented [Fel64]. Origin [Tal63]. Ort [SS62a]. orthophosphate [Rig68]. oscillations [Rao68]. Oscillatoria [BBK69]. Oscillaxanthin [BC63, GPE69]. Osmotic [Bro69a]. Ostracoda [Tib65]. Other [Chi60, US69, ARL^{+}63b, Cas62, DS64, Doy68, HMA66]. outcrop [NS62]. outdoor [ARL^{+}63b, ACLL64]. Outer [ZTTG65]. Outfall [BIR64a, BIR64b, BIR65]. overturn [Nob66]. oxidation [MC65, Wil69a]. oxidized [Orm66]. oxidizing [GS65]. oxine [Top69]. Oxygen [Bla60, Com65, Com68, PT68, PB69, PTL69, BC69, CM69, Car65a, Car65b, Car66, Dux63, FE63, Har69, LZS65, MB65, Ram60, Ram62a, Ram62b, Sha60, TC67, Wey65]. Oyster [Spa66, Wal65b, Wal66]. Oysters [Hop61a, HMA65, HMA66]. Ozark [AM67b]. Ozeanographie [Wyr67].

P [Hyn68, Owe69, Spa66, Ano66a, Aro69, Ban67a, Ban67b, Ban68, Bar68, Bod68, Car68a, Cha66a, Cha66b, Dee67, Dee69, Dic68, Dug68, Edm66b, Edm67, Edm68a, Edm69, Fot68, Fox66, Fre68, Fre69, Gro69, Has68a, Hay68, Ise69, Koh67, Koh68, Lar68b, Lew69a, Mca63, Och67, Ori68, Pai67, Pai68, Pai69, Par66b, Par66a, PP67a, Par67a, PP67b, Par68, Ray68, Ric67, Ric69, Roy68, Set67, Stu69, Tyl68a, Wyr66, Wyr67, BB68, Cus67]. p. [Edm66a, Edm68b, Edm68c, Fre66, Lag68, Val67]. Pacific [Dee67, Stu69, Wyr66, And69, BRS65, Bro65a, BB67b, CS66, Coc63, Des69, Gra66, Hob67, Mac67, MG62, Mor60, PK67, Rei62, She63, She64, Tho66, Val66, VT68, Wey65, Wyr65, vUB63]. pacifica [Jaw69, PTL69, SH67]. pail [Fre61a]. pairs [KP69]. Paleoecological [MD67]. paleolimnology [Ano66c, Edm69]. Panama [Che68]. Panel [Ano66a]. paper [JA67]. paperback [Pai69]. Papers [Art65]. paraffin [BN68]. paraffin-coated [BN68]. Paraffins [CB67]. Parameter [Whi69]. parameters [Bra68, Lid65]. Parapagurus [RM68]. Parasite [SW61]. Parey [Ban67b]. Paris [Cha66b, Edm67, Ano68d]. Park [Gil69, Vin67, WH67, WM67b]. Parks [Sto67]. part [Rei62, Mul64]. partial [CW68, DW67, Mil69]. Particle [Bur69a, MK65, Odu68, Bur68b, MK68]. particles [Jon60b, Orm66, SS69a, SEP67, SS69b]. Particulate [Fou68, SB65b, WG65, BS63, BW68c, Hob67, HM69, HH68, MR64, MV64, Mul65b, Mul65c, SC68b, SB61b, SB62b, Yen62]. parts [AO63]. patchiness [WW68b]. Patrick [Val67]. Patten [Dee67, LS69]. Patterns [Phl65, VR61, Woo65b]. Paul [Ban67b]. peat [EWR65]. Pectinaria [Gor66]. Pedro [GE62]. peel [Mer62]. Pelagic [Cor65, Gol62a]. Pellets [FCJ67, Dil64, JS66, Sma69]. penetration [Nut61]. Pennsylvania [Roy68, DD62]. percentage [Mos67b]. performance [AAR69]. Pergamon [Edm66a]. period [BH67, BK68, CC62a, CC62b]. periphyton [BN68, Dum69, GB60, KWV66, PM65, SS63, Wat61]. permanent [CD67]. persistence [KC64]. Persistent [LaR68a, LaR69]. personnel [Ano66d]. Perspectives [Kes60, Dee69]. persulfate [MC65]. Peru [CS69a, Woo61]. pesticide [BRZ68]. pesticides [SC68a]. Petenxil [Val67]. Petroleum [Hed60]. pH [BYK68, CP68b, Gie69, PKB66, Ver61]. phaeodactylum [ARL^{+}63b, ACLL64]. Phaeodactyylum [ARL63a]. phaeophytin [Lor65]. phagotrophic [Nor69]. Pharmacology [Sco65]. Phenolhypochlorite [Sol69]. pheo [Lor67, Mos67b, Tie68]. pheo-pigments [Lor67, Mos67b, Tie68]. Philadelphia [Roy68]. phosphatase [Hsi65]. Phosphate [BB65, CS69b, Gil67, Her61, Her62, PSG65, PMM63, PK67, Rei62, Ste63]. phosphate-phosphorus [Rei62]. phosphoric [KP67]. Phosphorus [HG68, Kue61a, Rig61, Cor65, DD61, FF67, Fit69, Joh64a, Joh64b, KC65, LB62, MCK64, MR65, MC65, PMM63, Rei62, Rig64, SG69, WH63a]. phospiiate [HJK63]. photic [DG68]. photocell [NA61]. Photochemical [Ham64]. Photodensitometric [IP66]. photographs [EMT65, HEHM60]. photography [Ell66]. photometer [RW69, SFM69]. photometric [CK65]. photon [WL68]. Photoperiod [Haz62, Paa68]. photophosphorylation [BB67a]. Photosynthesis [HT67, Lam67, Nal66, BB69b, McA61a, McC65, MS61, Mca61b, STH60, Ste62a, ST61, Tai60, Wri60, YR63]. Photosynthetic [BBK69, HT67, Jef68, CS65, Fee69, GDM64, Lor63, Man67a, RM65, Sol66, TeI68, TG68, Wat61]. photosynthetically [BB69c, SS62b]. phycomycetes [Vis60]. Phylum [Wei61]. Physical [Hor62, Rei60, Wri61, BL69, DCL63]. Physicochemical [AAT64, Dun67]. Physiological [HM66, WK61]. Physiology [Haq67, Sco65]. Phytoplankton [BS61, Hul62, Man69, Mar65, Mar68b, RV61, Ver60, WM67a, WM66, And65, BH67, CS63, CS65, Dav64, Den68, Dic69, Epp68, ERM69, ECS69, GMW63, Ham69b, Hel65, Hob66, Hol62, HR63, JA67, Jef68, Lew69b, Lor63, McN60, MS61, Moo63, MK68, MSE66, OM67, Owe63, Par65, PR68, RH60, SSG69, Sol64, Str67, SHHEL69, Tho66, Wal69, Woo62, Wri64]. Phytoplankton-zooplankton [Mar65, Mar68b]. Piedmont [NS62]. pigment [EM66]. Pigments [Cas63a, JA67, Jef68, Lor67, Mos67b, Tie68]. pilosimanus [RM68]. piston [BK62, DD69]. Plains [DD66]. planipes [Lon68]. Plankter [BR61]. Plankton [HO61a, Pat61a, WW68b, ARNA64, AAB^{+}65, BCA66, Che68, Clu65, Cus67, Dor61, DJKMS65, Eme68, FJ66, GE62, Gus61, HB62, Har68, Har66a, Jef69, Jos66, KC65, Lam62, LS67, NDT67, Ort62, PSS61, Pat61b, Pat62, PV68, Ril65, SN68, Str64, SHHEL69, Swa65, TH69, Wol65]. Planktonic [Bro65b, Her63a, Lew67, Lid66, BL65, Lan63, Mar61, Mca61b, MB67, Nal66, Tai60, Wil66, Wil68]. Plant [Cas63a]. plants [GK66]. plasma [Str67]. plastic [AMP^{+}63, CD64b, Jit64, Mar67b, MPSS61, Ste62b, ST61]. plate [BC60, BC61, LM64]. plates [Bod68]. platform [DB65]. plathemis [Kor65]. Pleistocene [Gol65]. Pleuroncodes [Lon68]. plume [Par66c]. plumosus [HC67]. pneumatic [FBS68, Wal67]. Pocasset [SMH65]. point [Whi60]. pointy [RC64]. Polar [Ori68, Joh63a, Man67a]. Polarographic [KEB62]. policy [Ano69e, Has68a]. Polikarpov [Ric67]. Polluted [Fre61b, Jef62a, KB67]. Pollution [Kat65, Ano67h, Ano69h, DH64, GLN64, McC66, NGL^{+}64]. Polychaete [Gor66, Geo64]. polychaetes [Man64]. polyedra [TLH62]. Polyethylene [Gol62c, Her62]. Polymorphism [Zar69]. Polyphosphate [SS68b]. Pomolbus [DW65]. Pond [Mat60b, Dun67, LB62, Sre64b, TD65, DS64, Wal65b, Wal66]. Ponds [CMD64, HS68, CWB67, Den63]. Pontellid [She63, She64]. Pontoporeia [Wel68b]. pools [Wei65, WL69]. Population [BBK69, Con64, Gil69, Dic68, Kah63, Kue61a, Kue61b, NDT67, TvU64, Wri65, Zei63, vUB63]. Populations [Mat60a, RR69, SW61, BH67, But69, DC62, EM66, Lor63, Mar61, Mar67a, MK68, PV69, PH62, PV68, TD65, VJ67, Vac69, Vin67, Pai67]. portable [Bug67, NA61, Red62]. Possible [Nor69, SW61, AR67, KS62, PH62, Ram62b]. Potential [FCJ67]. potentiometric [RC64]. pour [BC61]. practical [Red62]. Prague [Fre68]. Precipitates [Jon67b]. Precipitation [Bru69]. précontinent [Cha66b]. predaceous [MCC69]. Predation [Zar69, RT66]. Predation-Balanced [Zar69]. preformed [Par67b]. Preliminary [AO63, Eme68, Str64, BB69a]. Prentice [Koh67, Ori68, Pai69]. Prentice-Hall [Ori68, Pai69]. Preparation [KDCP67]. prepared [Coc63]. Presence [BF62, VTK62, WC62]. preservation [FF67, RPG69]. preserving [Moo63]. President [Ano66a]. Press [Car68a, Cha66a, Dee69, Dic68, Edm68a, Has68a, Koh68, Stu69, Wyr66]. Pressure [AM68, CP68b, HM64, Mac63, Fal69, Klo61, Lov68b]. Pressure-filtration [Mac63]. pressures [ZB65]. previously [Lun62]. Prey [Bla60]. Primary [BB67b, Cas63a, DD66, Hep62, KB65, MGPW64, RHP^{+}66, And64, AMP^{+}63, AR67, BO60, BK68, BB67a, Cas62, CB69b, DJKMS65, DG67, Hic69, KN69, MPSS61, Nie65, Pla69, Sre64b, SB61b, SB62b, TO66, Wat65b, Wri61, Edm68a]. Princeton [Och67]. Principal [WS62]. Principles [Rit65]. print [Ric69]. Printing [Ano66a, Spa66]. Probe [BYK68, PPF62]. Problèmes [Cha66b]. Problems [Cha66a]. Procedure [Whi60]. procedures [RR65]. Proceedings [Ano60f, Ano61d, Ano64c, Ano65d, Ano66f, Ano67i, Ano68c, Ano69i, Lau62, Lau63]. Processes [PD65, SR63, Ham64, Kur66]. Production [Ans67, Cas63a, Fag69, SB68, AMP^{+}63, BK68, Cas62, CB69b, Fee69, GC63a, Ham69a, Hep62, HC68b, IB69, KB69, LC69, Mac61, MPSS61, MCA65, MGPW64, Nie65, OC69, OH68, Pla69, Roy68, RHP^{+}66, Sre64b, SB61b, SB62b, TeI68, TO66, Wat65b, WM66, Wri61, Wri65, ACLL64, Lag68]. Productivity [Edm68a, McC62, WM67b, And64, AR67, BO60, BB67a, Bug67, BB67b, Dic69, DJKMS65, DD66, DG67, Hic69, JS61, KN69, KB65, KWV66, LS69, Mca63, NS62, PV68, VBO60]. Products [PP63, BBG^{+}62].

Professor [Woo63]. profile [HHSW66]. Profiles [BYK68, Eme61, HH69, Str68]. profondeurs [Cha65]. profundal [CU65]. program [Ano68d, Ano60a]. Prolonged [Bea66, YR63]. propagation [KB63]. Properties [GC68, AH64, PPW62, Pyt63, RMV61, Tyl61b]. proposed [Fag69]. Protected [TD67]. protection [Cha66b]. protection-La [Cha66b]. Protein [AM68]. protozoa [Joh65a]. Provincelands [ZTTG65]. pseudocalanus [McL65, Car65c]. Pseudodiaptomus [Jac61]. pseudoharengus [DW65]. Psychrophilic [AM68, MH64]. Publ [Dee67, Tyl68a]. Public [Ric69, Has68a]. Publication [Ano64d, Ano64e, Bar68]. Publications [Ano61e]. Publishing [Dug68]. Puget [Chi60, PPF62]. pulex [Par61, TD68]. Pulsed [Nes67]. pump [Bug67, FBS68]. pure [Car66]. pyrite [Val63].

quality [PH60]. Quantitative [Col69, Mac64, Mer62, Pen62, Dum69, GB66, GB60, SSG69, Str64, WM64]. quantities [BG64]. Quaternary [Ano61c]. question [Con66b]. quinquiesperforata [ST65]. Quoy [BR61].

R [Ban67b, Bod68, Dee69, Edm68a, Fre66, Has68a, Pai67, PP67a, Par67a, Val67, FZ67]. Racek [Val67]. Radiation [Mar67a, PPF62, SS62b]. Radioactive [Gut61, MR65]. Radioactivity [TLH62]. Radiocarbon [BRS65, BG69, Gol64]. radioecology [Ano66g, MTR64, Ric67]. radioisotopes [LO62a]. radiometer [SS62b]. Radiometric [PH63]. radionuclides [MT64, PNH66]. radiosodium [LH62]. radiostrontium [MG62]. radiozinc [RPG69]. Rafinesque [MH61]. Raia [MT64]. rain [MS62a, Yaa64]. rain-fed [Yaa64]. Rainier [Sto67]. raising [Liv67]. rajids [BS65]. Raleigh [GCV63]. Random [Set67]. ranges [Val66]. rapid [SFM69]. Raritan [DH64, Jef62a, Pat61a]. Rate [Bur69b, HT67, MD67, Neu66, BH67, CP61, KEB62, McC65, McL65, MR65, MO63, MB67, Par65, Sma69]. Rates [BR67, GDM66, Bis68, CD64b, CH68, DGR64, GD66, Gol65, Mar68b, Zei63]. Ratio [Man69]. ratios [CS65, Man67a]. ray [MT64]. rayii [Jaw69]. Reactions [Har68]. Readings [Con64, BA67]. received [Ano67c, Ano67d, Ano68a, Ano69b, Ano69c, Ano69d]. receiving [CD64a]. recently [SS62a]. Reconnaissance [AH62]. Record [Moh61, Nor64, RR67]. Recorder [DB69, TR69]. Recording [Gra67, AME67, Str68, WW68a]. records [Dav64]. recover [Fel64]. recovery [ZMB68]. rectangular [Rao68]. Red [HWE67, HB62, TG68, SP60]. Redefinition [Lym69, WLD69]. Redfield [Sea65a, Sea65b]. reduced [BB69b]. reducing [GS65, Lig63]. reductase [ECS69]. reduction [CS69a, ZB65]. reef [Eme68, Jef68, Joh67a, VT68]. reef-building [VT68]. Reeves [Edm69]. reference [Bro65b, Joh63a, Lew69b, Mar60]. Refinement [Moo63]. Refinery [CMD64, CD64a, TD65]. regenerated [DG67]. regeneration [BB65, Joh65a, Par67b]. Region [Fox66, BS65, PA60]. regional [Ril65]. Regions [Ori68, PK67]. register [Ano66d]. Reid [Wyr66]. Reihe [Ban67a, Ban68]. Reinhold [PP67a, Ric67]. related [Lov68a, OPP64]. Relation [Bur69b, AH64, BRS65, DC62, GC63a, MH62, Mar68b, MO63, Ram62b, RHP^{+}66, She63]. Relations [SB61b, Dri67, Gun61, McC62, SB62a, SB62b]. Relationship [Cas63a, DW65, Man69, McC65, MSE66, PT68, SS63, AO63, AL66, Bur68b, CW68, MF66]. relationships [KB65, Lor68, Mar65, Mar68b, McL65]. Relative [Haz66, PK67, SC68a, SC68b]. Release [KB69, Rig61, FF67, Joh64a, Joh64b, WJ67]. Relevance [Fou68]. reliable [PST69]. relicta [RPA68, TV69]. remote [ZMB68]. removal [LO62a]. renewal [Dic69]. repeated [WW68b]. Reply [Hyn69, Pat69]. repopulation [DH64]. Report [Ano65e]. representatives [Joh64a]. Reproducibility [PKB66]. Reproduction [EH61]. requirements [GC63b, Vis60]. Research [Ano65e, Ano61a, Ano67h, Ano69h, Lew69a, Wol65, Ano61c]. Researches [Par66b]. Reservoir [Lig63, Wri65, AM67b, Cow67, App66, Wri60, Wri61]. Reservoirs [HO61a, CH68]. Residual [Bum65]. resin [Top69]. resistance [Hop60]. Resonant [HB62]. Resource [Cha66a]. Resources [Pie64, Has68a, Ric69]. respect [HR63]. Respiration [HC69, SH67, TC67, Col69, KEB62, WS61, YR63, Zei63, Anr64b]. Respiratory [Bis68, MCC69]. respirometer [WW68a]. Response [BR61, PLPP62, Vac69]. Results [Shi62, CDD62, Har66b, Str64]. retainer [Bur68a]. Retention [SS69b, Gut65, WH63b]. reticulatus [CG62]. reversal [CS69c]. reversing [Nis64, TD67, Whi60]. reversing-thermometer [Nis64]. Review [Ano66a, Aro69, Ban67a, Ban67b, Ban68, Bar68, Bod68, Bro68a, Bro68b, Car68a, Cha66a, Cha66b, Dee67, Dee69, Dic68, Dug68, Edm66a, Edm66b, Edm67, Edm68a, Edm68b, Edm68c, Edm69, Fot68, Fox66, Fre66, Fre68, Fre69, Gro69, Has68a, Hay68, Hyn68, Ise69, Joy68, Koh67, Koh68, Lag68, Lar68b, Lew69a, Och67, Ori68, Owe69, Pai67, Pai68, Pai69, Par66b, Par66a, PP67a, Par67a, PP67b, Par68, Ray68, Ric67, Ric69, Roy68, Set67, Spa66, Stu69, Tyl68a, Val67, Wyr66, Wyr67]. rhenium [FM62]. Rhizosolenia [SB66a]. Rhythm [LaR68a, LaR69, PTL69]. Rhythmicity [LaR68a]. Ribonucleic [AM68, Sut65]. Rica [Bro65a]. rich [OH68]. Ricker [Roy68]. Ridge [Fox66]. Riedl [Ban67b]. riffle [Vin67]. Rift [Tal63]. Riley [WC62]. ring [ARNA64]. River [Gil69, Hay68, Heu69, Hyn68, Kat65, Mar60, MH62, SMH65, THJ67, WH67, WM67b, Dor61, DCL63, GLN64, Har67, Lam62, McC66, NGL^{+}64, Pra66b, SC68b, CS69c, Cow67, Cus67, DH64, DC62, DCL63, DM68, Hob66, OPP64, Par66c, PWY69, PNH66, She68, Sil64, SR63, Vin67, WDH63, WM67a]. Robert [Set67]. Roccus [MH61]. rock [NS62]. rock-outcrop [NS62]. Rocky [GO61]. Rohrbough [Ise69]. Role [EM63, NS62, Fre60a, IB69, Kuz68, SMSW69, VHJC68, DD62]. Rolle [Fre60a]. rosea [BR67]. rostrata [Jef60]. Rotenone [HO61a]. rotifers [Wil66]. rotifiers [Ano60h]. rotula [SB65a]. rough [AL68]. routine [LC69]. Row [Ray68]. rubidium [SPCF65]. Russian [Ric67, Nic60]. Russian-English [Nic60]. ruthenium [Jon60b].

S [Ano66a, Fot68, Fox66, Hyn68, Lag68, Spa66]. sa [Cha66b]. Saanich [GG64]. Sacramento [Heu69]. Saginaw [Bri67a]. Sagitta [SS68a]. salina [PT68]. Saline [Bro69a, Cas60, CWB67, EA65]. Salinity [DR69, Gra67, Joh67c, MD67, SMH65, Bro65c, DM68, Gun61, Hop60, HR63, Lan63, Lov68b, Lym69, Mor60, TR69, WLD69]. Salmo [MCA65]. salmon [Hun65b]. salps [Ost62]. Salt [Mat60b, Vis60, DT65, HD69, Kue61b, TK61, TW66, WK61]. salt-marsh [DT65]. salts [CW68, DW67, Pyt62]. sample [CD67, FF67, Gil67, LS67]. sample-injection [CD67]. Sampler [PLH68, ARNA64, Bar69, BA64, BCA66, Bro69b, Dor61, FM61, FA61, GB66, Gus61, HS69, Jit64, Joe64, Mac69, MR68, Mül65a, Mun64, PSS61, Par67c, PK69, Ree67, Sie63, Wal67, WW67, Whi62, YBF69]. samplers [Ste62b]. samples [CS63, EMT65, Epp68, Fre61a, Joe64, LCEP61, May60, Ort62, Wat69, WL66]. Sampling [Col64, Mac64, TM64, AME67, AAB^{+}65, Cas67, Clu65, Cow67, HAM68, Hil69a, HO61b, Kah63, Lac61, LP65, LM64, Mai69, Mer62, Pen62, PMR68, Swa65, VTK66, Wei65]. San [Heu69, Lig63, GE62, Lig63]. Sanctuary [DD62]. Sand [ST65, Sco65, BW68a, BW68b, Dil64, Fag68, Joh65b, Mai69]. sand-bottom [Fag68]. sandy [Joh67c, SB68]. Santa [BW68b]. sardine [Taf60]. Sardinops [Taf60]. Sargasso [Bee66, DGR64, GDM64, Hul62, Mar66, Mar68a, MS62a, MS62b, MS62c, RVW65, RMV61, WD66]. Saronicos [BK68]. Sars [Bod68, Zar69]. Saskatchewan [HS68, Raw60]. saturation [BYK69, ERM69, Pyt65, PK67]. saturometer [BYK69]. Say [DW65]. scallop [Mar60]. Scandinavian [Gei64]. Scattering [Tyl61b, IJW61, Spi68, SP69, TH62]. schimper [Nor69]. Schizothrix [Sha69]. SCHMID [Lew69a]. Schooling [MH61]. Schools [Ric69]. Schrader [Hob61]. Schweizerbart’sche [Fot68, Fre66, Fre68]. Schwoerbel [Edm66b]. Sci [Val67]. Science [Ano66a, Bar68, Ano60e, Ano62f]. Sciences [Ano60a, Ano65e]. scientific [Ano66d, Dee67]. scintillation [LC69, PPF62, PLPP62]. scleractinian [Col69]. Scopthalamus [Sco65]. Scotia [DT65, RK62]. Screened [Fre61a]. screens [SS69b]. scuba [FFF^{+}66]. Sea [BK68, Hed60, Jon60a, Ano66h, Ant66, BF63, CC62a, CC62b, CK65, Clu65, CFC68, DB65, Dug67, EMT65, Fra66, Gar65, Ham64, Har66b, HM69, HEHM60, IJW61, Jar67, JA67, LaF63, LO62a, LM64, May60, MPSS61, ND66, PV69, PH63, Pro62, Pyt62, Pyt63, Ram62a, RHMB62, Ste62a, SS62b, SU66, TP65, TH62, Vis61, WH63a, Whe67a, Whi60, Wil61, Wyr65, Ano66a, Bee66, BB63, BG69, DGR64, Fou68, GDM64, Has68b, Hug69, Hul62, Hun65a, Ise69, Kan67, Koh67, Mar66, Mar68a, MT64, MS62a, MS62b, MS62c, Ray68, RVW65, RMV61, Set67, SB65b, WD66]. Seabed [PB69, Squ69]. sealed [SCV69]. Seashore [GO61]. season [Yaa64]. Seasonal [BW68b, Cas60, HS68, Hea67, Hol69, Jef69, And64, BFC67, Her61, Hob67, Mar68b, Wal66]. Seasonality [Mul64]. Seattle [Cha66a, Dee67, Has68a]. seawate [SS68b]. Seawater [CP68b, Gie69, PKB66, AC67, AME67, AL64, BB69a, BS63, BYK69, CS66, CS69a, CW68, DW67, FSS67, FZ67, Gil67, Gol64, GRB66, Jan67, JGST67, Jon67a, Jon67b, KP67, KDCP67, KP68, KP69, KZ68, LZS65, MV64, MC65, Mil69, OCV63, PD65, Ril63, SEP67, SPCF65, Spi68, SP69, SLC68, Top69, VJ66, VJ67, VHJC68, Vac69, Wey64, WL68, WS66, Wil69b, Wil69c, Wil69a, Woo65a]. seaweeds [Gut65, KB69]. Secchi [Gra66, Tyl68b]. Second [Ano66g]. Sediment [CB67, AH64, Bur68a, CP68a, DD69, Doy68, FA61, GPE69, GE62, Liv67, SSP68]. Sedimentary [Lid65, Gro69, Nri68]. Sedimentation [Bur69a, Liv65, Lud67, She68, Bru69, CH68, Gol65]. sedimentology [Edm69]. Sediments [Gor66, Gro65a, Ant66, AH64, BC63, Bro68c, Bru69, BH69, Cha62, Col64, DWW68, FS69, GS65, GG64, Har67, HM69, JM68, KEB62, LJ69, Lud69, Mai69, PSG65, Ree69, SB67, Tie68, Val63, WW67, Zei64]. Seen [Fre60a]. seiche [FB68, Ste64]. Seiches [EL68]. Selected [San67, Ric69]. Selection [Odu68]. Selective [RT66]. selectively [BB68]. Self [Clu65]. Self-closing [Clu65]. semen [BF62]. semiannual [Wyr65]. semidrainage [Pen69]. seminar [Ano69g]. semipermeable [HL64]. sensible [Has68b]. sensitive [AAR69, RW69, SLC68]. separate [Fox66, Stu69]. separating [Dil64, Tea60]. separation [JA67, Lam62, WL66]. Ser [Par67a]. Sergestes [PF69]. Serial [Bod68]. seriata [SB65a]. series [TD65]. seston [WM67a]. set [EL68]. set-up [EL68]. setigera [SB66a]. settle [SSG69]. settle-freeze [SSG69]. Several [CG62, SC68a]. shaker [DS68]. shallow [Bam62, Cas67, Dun67, Fag68, FO65, HAM68, Kah63, Lee61, Mai69, PV68, Ram62b, Sie63, Wei65, Whi62, Woo65b]. shallow-water [Kah63]. shear [Lov68a, OK69]. shelf [Bum65, Cor67, KC64, Lid66, Mul64, Val66]. Shell [Bar62, JCG64]. shells [PG63]. ship [TH69]. shipboard [JL66]. Shoals [App66]. shore [BA64, Ram62b]. short [CC62a, CC62b, Lyn66, Mac69, Sho64]. short-period [CC62a, CC62b]. short-term [Sho64]. Shrimp [Bro69a, Her63b, PT68, PF69]. shrimps [JL66, LT65]. Si [THJ67]. Sibert [Pat69]. siciloides [Com65]. Sicily [TG68]. Sierra [BBH68]. sifting [Fre61a]. Significance [FCJ67, LaR69, Odu68, CS69b, Dug67, HHB66, SS68c, VR61]. Silica [Har67, FS69, WE69]. silicate [Her61]. silty [RK62]. silver [VT68]. Similarity [Har69]. similis [PF69]. Simocephalus [Par60]. Simple [HO61b, BA64, Car68b, Dri64, Eme61, HC68b, Mai69, MR66, Orm66, RW69, Sie63, Uke65]. simultaneous [CP68a, EMT65, Lov68b]. sinking [Sma69]. siphon [Har66a]. situ [ST61, Ano65e, BYK68, DB69, GDM66, Gol62c, Gra67, MK65, Nes67, PF68, PB69, Wat65b]. six [Jit64]. six-liter [Jit64]. Size [Bur69a, Bur69b, Mul65b, Mul65c, Bur68b, MF66, McL65, MO63, SS69a]. skeletal [PH66]. Skeletonema [Cas63a, Cur62, Pra66a, SB66a]. skeletons [SU66]. Skill [LS67]. skimmer [FP62]. skin [Owe69]. slick [Gar65]. slick-forming [Gar65]. Small [Chi60, Hea67, BS65, Dil64, Dri64, DS68, Kah63, McM69, MR66, Pfa62, Scr60, SEP67, Sut65, Tea60, TSB63, Wei65, WK61, ARL63a]. Small-Sscale [ARL63a]. smelt [RT66]. Smith [CP68a, Her63b, RM68]. Smith-McIntyre [CP68a]. So-Called [Bar62]. Soc [Par67a]. Societas [Ano69j]. Society [Ano60f, Ano61d, Ano63a, Ano63b, Ano64c, Ano65d, Ano66f, Ano67i, Ano68c, Ano69i, Bod68, Lau62, Lau63]. Sodium [KP69]. Soft [Mat60b, Liv67, San60]. soft-bottom [San60]. solubility [Car66, Klo61, Klo63, PK63]. soluble [McM69, PMM63]. solutes [WH65]. Solution [Lew67, Hir63]. solutions [BF63, JS61]. Some [BBK69, Bar62, CDD62, Col61, Dic69, EA65, Gol68, Gun61, HM66, HS68, HT67, Hob66, HT69, JCG64, Joh67b, LBL60, MJ64, MCA65, McL65, Mul63, NDT67, PH62, PG60, RCBA65, Sho64, Sma69, SH60, TP65, WM64, ARNA64, BR65, Dai69, Enr63, Gol62a, GS65, GC63b, Hel65, Jef68, Lan63, Mos69, RMV61, Sut65, Wri60]. Son [Hay68, Hyn68]. sonic [ZMB68]. Sons [Fre69, Gro69, Lag68, Pai68]. sorption [HJK63]. sorter [MF64]. sorting [LCEP61, Str64]. Sound [AH62, Lyn66, Rop63, HB62, Lov68b, Chi60, PPF62, VR61, Wil68]. sounder [Nor64]. source [AR67, BS63, KS62, OK69]. Sources [DD61, DD65, JCG64]. sous [Cha66b, Cha66b]. sous-marine [Cha66b, Cha66b]. south [BB67b, KC64, She64, WM64, HM69, Hop60]. Southampton [EH61]. southeastern [Cor67]. Southern [CB69a, AL66, CC62a, CC62b, DD66, GC63a, HHSW66, HR63, Mor60, Rot68, vUB63, Hed60, Wal69]. southwestern [Hul62, McC68]. sp [GDM66, Jon60a, MJ64, Wat65a]. spawning [CT61]. special [Bro65b, Joh63a, Lew69b, Mar60]. speciation [Man64]. Species [Bur69a, Bur69b, Col61, Hol68, Jef62b, Man69, ARL^{+}63b, Buz69, Dic69, Gil69, HR63, Man64, Mor60, PV68, Rot68, SC68a, TvU64, vUB63]. specific [Par66c]. spectra [CC62a, CC62b]. Spectrometry [Joy68]. Spectrophotometric [Cli69, BA67, Lor67, Mos67b]. speed [AC69, Eve67, KB63]. speedometer [Dri64]. sphere [AMP^{+}63, MPSS61]. spherules [Val63]. sphincter [Bur68a]. splitter [LS67, SN68]. Sponge [Neu66]. spp [HS68, PV69, San67]. spread [BC60, BC61]. spring [Bro69b, Hll67, Pra65]. spring-actuated [Bro69b]. Square [Koh68]. Sr [LO62a, THJ67]. Sscale [ARL63a]. St. [Bee65, IJ61]. stability [Gol68, Pat62, Rag62, You61]. Stabilization [May60, Pro62, EM63]. stabilized [Van60]. Stabilizing [MP65, Bar69]. stable [DB65, MGPB65]. Stages [Lew67]. stagnant [GG64]. stained [Den68]. stained-organism [Den68]. standard [CS69b]. Standing [Bro67, AM67b, CU65, Hop69, Lyn66, RR67, Wal69, Wel68a]. starch [Van60]. State [Sre64a, But69]. States [Spa66, Cha66c, Cor67, EWR65, WS62, Wil66]. station [Mca63]. statistical [Taf60]. steady [But69]. steady-state [But69]. Stewart [Och67]. stiff [Col64]. Stimulation [Nes67]. Stizostedion [SP60]. Stoffkreislauf [Fre60a]. stoneflies [Ano62d, SC68a]. storage [Gil67, Her62, RPG69]. storm [Hun65a]. Strait [CA66, SR63, Wal69]. Straits [BS61, AC63, BFC67, PA60]. Stratification [BPTW65, Bea66, Cou68, Har68, Tal63]. stratified [HS69]. stratigraphie [Cha66b]. Straus [MR65, Rig61]. Stream [Che67, TM64, BN68, DD66, Fag69, GB60, HG69, Hil69a, HC68b, KB67, Mül65a, NS62, SG69, TO66, WE69, Wat61, Wat69, WL66, WDS64, FV65, Pra66b, Rei60]. streams [KB65, MGPW64, Mun64, PM65, Whi62]. stress [Mar67a]. Striped [Odu68]. stromatolites [Sha69]. strong [VBO60]. strongylocentrotus [PV69]. strontium [SS62a]. Structure [Kue61b, AO63, Mar61, San60]. structures [LaF63, Gro69]. students [Ano60h]. Studies [ARL^{+}63b, ARL63a, ACLL64, Art65, Bee66, BFC67, ECS69, Fre68, Man64, Mar60, PH69, Stu69, VJ66, Wal65a, Wal65b, Wal66, WJ67, Whi69, Wyr66, BBG^{+}62, Bug67, CDD62, DB65, DD65, FFF^{+}66, Gil69, Hsi65, Jef64, JS61, Kuz68, LC69, MS61, Owe63, PLH68, PA60, San60, SSG69, Sre64a, Tai60, Ver64, VBO60, Wat65b, Wie60a, WK61, WH67, WM67b, CB69b, Hyn68, SHHEL69, Car68a, Hay68]. Study [Ori68, San67, AC69, BB69a, Cas63b, Dee67, Dic68, Dum69, FB68, Gas66, Gro69, HB61, IJW61, MW69, RA66, Sai61, SGMH62, ST61, Taf60, TD64, TD67, VTK66, WH63a, WDS64, Woo62]. Studying [Gol62c, DWW68]. Stumm [Par67a]. Stuttgart [Edm66b, Edm68b, Edm68c, Fot68, Fre66, Fre68]. Style [Ano61f]. sub [Gus61, Whe67b, BS65]. sub-Antarctica [BS65]. sub-bottom [Whe67b]. sub-sampler [Gus61]. subaqueous [Mac69]. subarctic [GD66]. subject [Scr60]. Subjected [Joh67b]. Sublethal [CG62]. sublittoral [CU65]. Submarine [Hug69, RPA68]. submerged [EWR65, ZMB68]. Subsampler [Wat69, Hop62]. subscribers [Ban67a, Ban68]. substances [HHSW66, Kuz68, PR68]. substrata [KWV66]. substrate [CM69, Dri67, Hil69a]. substrates [BN68, RAF^{+}62]. Subsurface [And69]. Subtidal [PF68]. Subtropical [Ber61, Bro65c, RWV64]. Succession [HS68, Jef62b, SB67]. successive [Nut61]. Succinate [PT68]. sugars [AL64]. Sugawara [Par66b]. Sulfate [Lig63, KP68, KP69]. Sulfate-reducing [Lig63]. sulfide [Cli69, WC62]. Sulfur [Nri68, Cur62, TeI68, TG68]. sultana [App66]. summary [Sch62]. Summer [Hob61, WG60, CU65, Ham69b]. Sun [Tyl61a]. Sun-altitude [Tyl61a]. Sunfish [App66]. sunlight [BB69b]. superfluous [Con66b]. Supplement [Ano64f, Ano65f, Gol62b, Ano62c]. Supplies [SV65]. supply [Wey65]. support [Ano68d]. supported [Top69]. surf [Sch65]. Surface [Des69, JGST67, MH61, Tho66, BB67b, CK65, CS67, Fal69, Gar65, Gor60, GCV63, Har66b, HJK63, Jar67, Kan67, Mul65b, Mul65c, NDT67, Ram62a, RWV64, RVW65, She63, Sho64, Wyr65]. surface-active [Jar67, JGST67]. Surge [DB69, Hun65a]. Survey [Gro65a, AO63, Dai69, Lew69a, PLPP62, Wal65a]. surveying [MR66]. surveys [Ell66, Hol62, Wis62]. Suspended [CMD64, GE62, BFC67, Hob67, SS69a, Wan69]. Suspension [Edm66a]. Suspensions [BR67]. swash [Sch65]. sweetwater [Ram60]. Sydney [Fre69]. Sylvester [Has68a]. Symbioses [Koh68]. sympatric [Man64]. Symposium [Ano60g, Ano66h, Jac65, Ano66c, Ano66g, Ano69g]. Synchronously [RR69]. Synthesis [AM68]. synthetic [Jon67a, KP68, Nix69]. syringes [Bro69b]. System [IJ61, BCA66, MB65, Par69, PST69, Red61, Sol64, SCV69, Swa65, Woo65a]. systems [CD67, WH63b, Par67a].

T [Has68a, Koh68, Lew69a, Par66b]. Tagus [HD69]. taken [Str68]. Tanganyika [Cou68, GC63a, Liv65]. tank [SHHEL69]. tanks [ARL^{+}63b]. tape [Liv67]. taxonomy [RAF^{+}62]. Technical [Fox66, Ano66d]. Technique [Kah63, BB68, Car65b, Den68, Dil64, DWW68, Ell66, Par65, RM65, Sug67, Tea60]. Techniques [PTL69, TM64, DJKMS65]. temperate [Gor64, Jef69]. Temperature [BYK68, BR61, Bur69b, DB69, DT65, Gie69, GRB66, Gra67, Hah63, Haz62, Joh65b, KDB68, MD67, MH64, Mor60, Scr60, Smi62, WW66, LaF63, Lov68b, Lov68a, MJ64, McL65, MO63, Paa68, PM65, Wey64, Wyr65]. Temperature-Limited [BR61]. Temperatures [HM66, Joh67b, Hll67, Hob61]. ten [Par69]. Tendipes [HC67]. Tennessee [Fox66]. tentaculata [Geo64]. term [Sho64]. tertiolecta [HT67]. Tessiarsuk [Car65c]. Test [TLH62, Hor62]. tests [Whi60]. tetrachloride [WL66]. tetraselmis [MJ64]. Texas [Phl65]. Thalassiosira [SB65a]. Their [Bla60, Car68a, Fou68, AL66, KB69, MH62, PBK69, Ril63, SMSW69]. them [Ver60]. Theoretischen [Wyr67]. theory [Dee69]. theperiod [Ste64]. Thermal [Cou68, Maz67, TSB63, BB69b]. thermodynamic [Klo63]. thermometer [Nis64, TD67]. thermometers [Whi60]. thermophilic [SG69, Sto67]. thermopile [SS62b]. thiamine [GRB66, Gol68, ND66, Vis61, VR61]. thickness [BW68a, JS61]. thin [PLH68]. Thompson [Fox66]. thorium [Ant66]. thornback [MT64]. Three [Mur62, TM64, Col69, Mor60, Rot68, SC68a, Sre64a]. Thunnus [PO68]. Thurlow [Car61]. tidal [BLM63, Hug69, Red61, Scr60]. Tide [Hun65a, Red62]. Tides [Lar68b]. tigriopus [LO62b]. time [Fee69, Rig64]. Timing [CS69c]. Tissue [GK66]. titration [VBO60]. Tokyo [Par66b, Par66a, PP67b]. Tokyo-Do [PP67b]. tolerance [Lan63]. Tonga [BL65]. tool [Wol65]. topographies [Dux62]. topography [Whe67b, vA65]. Tortanus [Joh67b]. total [CD67, KB69, Mai66, McM69, MC65, RPG69, Ver61]. Towable [SFM69]. towed [FC65]. towers [Dea63]. Towing [AAB^{+}65, ARNA64]. Town [Ric69]. tows [WW68b]. Toxaphene [HO61a, SO61]. Toxicants [CG62, WL69]. toxicities [SC68a]. toxicity [GW60]. Trace [BBH68, BR65, PG63, DWW68, PH62]. Tracer [WH63a, FZ67, DD65]. traces [Bro65c, HJK63]. tracking [FV65]. Trailing [Owe63]. transfer [Sho64]. Transl [Ric67]. Translation [Ano60a]. transmittance [IJW61]. transparency [CFC68]. Transport [Che68, Cus67, PNH66, PA60, Rod61]. trap [AL68]. Travel [Ano68d]. trawl [ARNA64, Ber66, LM64, Wis62]. treated [Jon67a, NGL^{+}64]. treatise [Fre69]. trends [BIR64b, BIR65]. Trichocysts [US69]. Trichodesmium [BL65, GDM66]. triphosphate [HHH67, HHB66]. tripneustes [SU66]. tripping [Ste62b]. tripton [SC68b]. Trophic [FCJ67, Hol68]. Tropical [BPTW65, Ber61, AC67, Bea66, Bro65a, Bro65c, Dun67, Maz67, RWV64, Sre64b, Tho66, GDM66]. trout [MCA65]. trutta [MCA65]. Tsuchiya [Stu69]. Tsukada [Val67]. tube [Eve67]. tuna [Ano62g]. tundra [BR64]. tungsten [FM62]. tunicates [STK69]. turbidity [AL66, EHR62, PG60]. Turbulence [Haz66, Lov68a]. turnover [Rig64]. tween [BC61]. twelve [Raw60]. twenty [AH64]. twenty-five [AH64]. twin [Jit64]. Two [Bur69a, CC62b, GMH67, HO61a, Jef62b, RM68, CWB67, GN64, HG69, LJ69, Mac67, PK67, RL64, STK69, Wil69c, CC62a]. type [AL68, Dux63]. types [Bro65b, HG69, Mok60, Rig64, She63].

U [Ano66a, Fox66, Val67]. U.S [Ise69, Ric69]. U.S. [Ano61e]. Uganda [Bea66]. Ulva [Gut61, Ver62]. Underwater [TLH62, AAR69, CL62, HB62, MR66, RW69, TPS64, Tyl61a]. undescribed [Mur62]. undisturbed [MR68]. unialgal [Sol64]. uniform [OK69]. United [Spa66, Cha66c, Cor67, EWR65, WS62, Wil66]. Univ [Dee69, Edm68a, Hay68, Hyn68]. University [Car68a, Cha66a, Dic68, Has68a]. unstratified [Bye65]. Unusual [CWB67, Web61]. upland [Sre64a]. Upon [Bar62, HO61a, EHR62, GS65, Klo61]. upper [DD69, Rei62, Stu69]. Uptake [BO60, Com65, DG67, GDM66, Gut61, Gut65, Joh64a, Joh64b, Kor65, MG62, Rig61, SS61, TLH62, WH63b, Wil60, BH67, BBG^{+}62, BR65, ERM69, GDM64, Gut63, Har69, HW65, HC69, Owe63, SG69, VJ67, Ver62, WS61, WH65]. upwelling [Lor68, PPW62, PH60, WG60]. uranium [VT68]. urchins [PV69, SU66]. Urnatella [Wei61]. Urosalpinx [Bla60]. Use [Ano66a, App66, CD64b, KWV66, Nor64, SW61, SS62b, Wel68a, Wil63, AAR69, BA64, DS68, Dux63, FFF^{+}66, FBS68, HL64, LC69, Liv67, LS67, MS61, PPF62, PLPP62, SHHEL69, TH69, Van60, Whi62]. used [JS61, MF66]. Using [BYK68, AMP^{+}63, Bro69b, CS63, CK65, FFF^{+}66, Gol64, MPSS61, Mca61b, RC64, Squ69, ST61, Top69]. utilization [ACLL64, Ril63, Lew69a].

v [Fot68, Koh68, Lew69a, Spa66, Val67, CB69b, Mul64]. vacuum [Bug67]. Valid [Jac65]. valida [RM68]. value [JS66]. vanadium [FM62]. Vancouver [PNH66]. variability [RK62]. variable [Rao68]. variables [Buz69]. Variation [Che67, Klo63, Lyn66, WE69, BW68b, BFC67, Her61, Joh65b, Lor63, McM69, Mca61b, Mca63, PK63, SB61a, Wal66, Wyr65]. Variations [CS65, OM67, Sil64, VJ67, GDM64, Lov68a, PBK69, Ril65, RMV61, Scr60, THJ67]. varicans [Cha69]. Various [PP63, Nal66, RAF^{+}62]. vascular [HB66]. Veen [LP65]. vegetation [Pen62]. vellar [DR69]. velocity [Har66a, Lov68b]. Venezuela [Red61, Zei64]. Verlagsbuchhandlung [Fot68, Fre66, Fre68]. Vertical [Bra69, Bri67b, Bro67, Her63b, LB62, Lov68a, TeI68, TT66, WD66, Wal69, Cas63b, Hob67, Mar66, McM69, RH60, SB61a, YBW64]. Vertically [SH67]. vetulus [Par60]. Vibrio [AM68, HM66]. Vicente [Lig63]. vicinity [Joh67a]. VII [Ano62c, Ano63e, AH64, Gol62b]. Villars [Edm67]. Vineyard [WM64]. virginica [Spa66, US69]. viridis [Par60]. visible [Yen62]. vitamin [CS66, Cur62, Dai69, Gol64, MS62c]. vitreum [SP60, SP60]. vitro [CB60]. vol [Ano64f, Ano65f]. volcanic [Kur66]. Volume [Fre69, Par66b, Par68, AME67, Ano61b, Ano62c, Ano62e, Ano63c, Ano64b, Ano65c, Ano66b, Ano68b, Ano69f, AMP^{+}63, Gol62b, MPSS61, MSE66, SS69a, Str67, TH69]. volumes [DW67, Mil69, Rei62]. volumetric [WW68a]. vs [Cur60, VBO60].

W [Dee67, Fot68, Koh67, PP67a, Par67a, Roy68, Wyr67, Rei63]. Wadachi [PP67b]. walleye [SP60]. Waltair [Rao69]. Wandering [Mau65]. Washington [Ano66a, Bar68, Cha66a, Has68a, Par67a, Spa66, And64, And65, Cas60, EA65, GPE69, OCO67, PNH66, SB67, Whe67b]. waste [EM63]. Water [BR67, Fre60b, Gol62c, GLN64, HC67, Jon60a, NGL^{+}64, OCO67, PH60, PA60, RK62, Rod61, AL68, Ano67h, Ano69h, AH64, ARNA64, BG64, BF63, BA64, Bea69, BRS65, BH69, CP68a, Car66, Cas67, CA66, CS67, CD67, DC62, DWW68, DR69, Fag68, FM61, FF67, FE63, FO65, Gje64, HAM68, HB62, Har68, HMA65, Her62, Hll67, HWE67, Hop61b, HS69, Jit64, Joe64, Joh67c, Kah63, KC64, Lam62, LO62a, Lee61, Lid66, Liv65, Mai69, May60, MPSS61, MCK64, McM67, MS62a, OPP64, Par67a, PH62, PSG65, Pro62, Pyt62, Pyt63, Ram62b, Ree67, Rig68, RHMB62, Scr60, She63, Ste62b, TG68, TH62, VBO60, Vis61, WH63a, Wei65, Wel68a, WS61, Wil61, WH65, YBF69, vUB63, EH61, Has68a]. water-level [DC62]. Waters [Bro69a, Car68a, Fre61b, Gei64, Rei63, SS68a, Sol69, Wyr66, BB67b, Cli69, CD64b, Cor65, Dai69, Dor61, EHR62, Fee69, GC68, GCV63, Har67, Her61, HC68a, HZ62, HC69, HS69, Jef68, Kan67, Mac61, Mai66, MP65, MS62a, Mos69, Mul64, Mul65b, Mul65c, NA61, RWV64, RVW65, Sha66, Smi68, Ste63, Stu69, Tyl61b, Ver61, Wal65a, Wal65b, Wal66, Woo65b, Yaa64, Roy68]. waters.1 [SWBL63]. waterways [WS62, Wil66]. waterwheel [PK69]. watery [Sie63]. wave [CC62a, CC62b, HT69]. wavelength [KB63]. waves [AL66, BR60, Fal69, KB63, Lee61, Rag62, RR67]. Weatherley [Car68a]. weeds [Fit69]. weighing [Sug67]. Weight [Kor68, SS63]. weights [RR65]. Wellen [Wyr67]. Wequaquet [Wal66]. West [PO68, Wey63]. western [CH65, Hil69b, MCK64, MR64, Mul65b, Mul65c, RHP^{+}66, Ver60]. wet [SS63, Wil69a]. White [MH61]. Widespread [Bro66]. Wildlife [Pau65]. Wiley [Fre69, Gro69, Lag68, Pai67, Pai68]. Wilimovsky [Fox66]. Williams [Ise69, Jac61]. Wilson [DW65]. Wind [Bye65, Sma63, HB61, KB63, Rag62, SB61a]. Wind-driven [Bye65, SB61a]. Winkler [Car65a, Car65b]. Winnipeg [EL68]. Winter [Jef60, Ano62d, KC64, Pra65, RH60, Sai61]. Wisconsin [Dic68, GN64, LJ69, Nri68]. within [LH62]. Wolfe [Fox66]. Wood [EH61, Bro68a, Bro68b]. Wood-Boring [EH61]. Woods [Ric69, WK61]. work [Dux63]. Working [Ano65e]. World [Koh67, Ano60h, Aro69].

X [Ano63f]. xi [Fre69]. xiv [Lag68]. Xv [Ano61g, Ano66a, Car68a, Edm68b, Edm68c]. xvi [Fox66, Gro69]. Xxiv [Edm67]. xxviii [Ric67].

year [SP60]. Yearly [Woo61]. Yeast [BR67, TvU64, vUB63]. yeasts [FAMR60, RAF^{+}62]. yellow [GC68]. Yellowstone [Gil69, Sto67, WH67, WM67b, Vin67]. yield [Sre64a]. York [Bod68, Bro68a, Bro68b, Edm66a, Edm69, Fre69, Gro69, Lag68, Lar68b, Owe69, Pai67, Pai68, PP67a, Par68, Ray68, Ric67, Set67, Tyl68a, BL69, Bru69, BH69, CB69b, Joy68, KC64, Lud67, Lud69, Web61].

Zealand [BR65]. zero [BW68a, JS61]. zero-thickness [JS61]. Zinc [Gut61, OPP64, PO68, Gut65, Kor65, RPG69, RHMB62, Top69]. Zinc-65 [PO68, Kor65]. Zn [BO60, Cus67, Gut63, MO63, Ost62, Ver62]. ZoBell [Sie63]. Zonation [Mat60a, Mok60]. Zone [FH69, DG68, PMR68, Sch65, YBW64]. Zones [GO61]. zooplankter [MCC69]. Zooplankton [AM67b, Bie62, HW67, Hop69, Joh63a, BB65, Bee66, Bis68, Con66a, Con66b, CB69b, FC65, GLN64, HG68, Hop62, Hop60, Jef64, Lew69b, Mar65, Mar68b, MF66, MF64, OCO67, Pen62, PMM63, RC61, Rei62, WJ67, Zei63].

References

Aron:1965:TCP

[AAB^{+}65] W. Aron, E. H. Ahlstrom, B. McK. Bary, A. W. H. Bé, and W. D. Clarke. Towing characteristics of plankton sampling gear. Limnology and Oceanography, 10(3):333–340, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Ackefors:1969:CPS

[AAR69] Hans Ackefors, Gunnar Ahnström, and Carl-Gustaf Rosén. Construction and performance of a sensitive light meter for underwater use. Limnology and Oceanography, 14(4):613–617, July 1969. CO-DEN LIOCAH. ISSN 0024-3590.

Angino:1964:PLL

[AAT64] Ernest E. Angino, Kenneth B. Armitage, and Jerry C. Tash. Physicochemical limnology of Lake Bonney, Antarctica. Limnology and Oceanography, 9(2):207–217, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Alexander:1963:DCS

[AC63] James E. Alexander and E. F. Corcoran. Distribution of chlorophyll in the Straits of Florida. Limnology and Oceanography, 8(2):294–297, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Alexander:1967:DCT

[AC67] J. E. Alexander and E. F. Corcoran. The distribution of copper in tropical seawater. Limnology and Oceanography, 12(2):236–242, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Aron:1969:SIN

[AC69] William Aron and Sneed Collard. A study of the influence of net speed on catch. Limnology and Oceanography, 14(2):242–249, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Ansell:1964:SMC

[ACLL64] Alan D. Ansell, J. Coughlan, K. F. Lander, and F. A. Loosmore. Studies on the mass culture of phaeodactylum. IV. Production and nutrient utilization in outdoor mass culture. Limnology and Oceanography, 9(3):334–342, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Armitage:1962:LRA

[AH62] Kenneth B. Armitage and Hugh B. House. A limnological reconnaissance in the area of McMurdo Sound, Antarctica. Limnology and Oceanography, 7(1):36–41, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anthony:1964:LWS

[AH64] E. H. Anthony and F. R. Hayes. Lake water and sediment. VII. Chemical and optical properties of water in relation to the bacterial counts in the sediments of twenty-five North American lakes. Limnology and Oceanography, 9(1):35–41, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Antia:1964:DFA

[AL64] N. J. Antia and C. Y. Lee. The determination of “free” amino sugars in seawater. Limnology and Oceanography, 9(2):261–262, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Armstrong:1966:CNC

[AL66] F. A. J. Armstrong and E. C. LaFond. Chemical nutrient concentrations and their relationship to internal waves and turbidity off southern California. Limnology and Oceanography, 11(4):538–547, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1968:BTE

[AL68] David W. Anderson and Douglas W. Larson. A bottom-type emergence trap for rough water conditions. Limnology and Oceanography, 13(1):181–182, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1967:ALF

[AM67a] D. V. Anderson and D. H. Matheson. An articulated limnological float. Limnology and Oceanography, 12(3):162–163, April 1967. CO-DEN LIOCAH. ISSN 0024-3590.

Applegate:1967:ZSC

[AM67b] Richard L. Applegate and James W. Mullan. Zooplankton standing crops in a new and an old Ozark reservoir. Limnology and Oceanography, 12(4):592–599, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Albright:1968:EHP

[AM68] Lawrence J. Albright and Richard Y. Morita. Effect of hydrostatic pressure on synthesis of protein, ribonucleic acid, and deoxyribonucleic acid by the psychrophilic marine bacterium, Vibrio marinus. Limnology and Oceanography, 13(4):637–643, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Andersen:1967:DRL

[AME67] Neil R. Andersen, Robert E. Morris, and Gerald H. Efinger. Depth recording of large-volume sampling of seawater. Limnology and Oceanography, 12(4):712–714, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Antia:1963:FMP

[AMP^{+}63] N. J. Antia, C. D. McAllister, T. R. Parsons, K. Stephens, and J. D. H. Strickland. Further measurements of primary production using a large-volume plastic sphere. Limnology and Oceanography, 8 (2):166–183, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1964:SGD

[And64] George C. Anderson. The seasonal and geographic distribution of primary productivity off the Washington and Oregon coasts. Limnology and Oceanography, 9(3):284–302, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1965:FPC

[And65] George C. Anderson. Fractionation of phytoplankton communities off the Washington and Oregon coasts. Limnology and Oceanography, 10 (3):477–480, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1969:SCM

[And69] G. C. Anderson. Subsurface chlorophyll maximum in the northeast Pacific Ocean. Limnology and Oceanography, 14(3):386–391, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:AIB

[Ano60a] Anonymous. American Institute of Biological Sciences Translation Program. Limnology and Oceanography, 5(4):433, April 1960. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:BM

[Ano60b] Anonymous. Botanica marina. Limnology and Oceanography, 5(4): 434, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:E

[Ano60c] Anonymous. Erratum. Limnology and Oceanography, 5(3):348, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:ILO

[Ano60d] Anonymous. Index to Limnology and Oceanography. Limnology and Oceanography, 5(4):437–440, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:NSF

[Ano60e] Anonymous. National science foundation grants. Limnology and Oceanography, 5(1):109, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:PAS

[Ano60f] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 5(1):115–119, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:SMM

[Ano60g] Anonymous. Symposium on marine microbiology. Limnology and Oceanography, 5(4):434, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1960:WLS

[Ano60h] Anonymous. A world list of students of rotifiers. Limnology and Oceanography, 5(3):343–345, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1961:FCG

[Ano61a] Anonymous. Fourth conference on Great Lakes research. Limnology and Oceanography, 6(1):97, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1961:ILO

[Ano61b] Anonymous. Index to Limnology and Oceanography, volume 6, 1961. Limnology and Oceanography, 6(4):492–496, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1961:IAQ

[Ano61c] Anonymous. International Association for Quaternary Research. Limnology and Oceanography, 6(1):97, April 1961. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1961:PAS

[Ano61d] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 6(1):101–104, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1961:PUA

[Ano61e] Anonymous. Publications by the U.S. Atomic Energy Commission. Limnology and Oceanography, 6(1):97, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1961:SMB

[Ano61f] Anonymous. Style manual for biological journals. Limnology and Oceanography, 6(4):491, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1961:XIC

[Ano61g] Anonymous. Xv international congresso f limnology. Limnology and Oceanography, 6(1):97, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:A

[Ano62a] Anonymous. Announcement. Limnology and Oceanography, 7(4):490, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:EN

[Ano62b] Anonymous. Editor’s note. Limnology and Oceanography, 7(S1):ii, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:ESV

[Ano62c] Anonymous. Errata — Supplement to volume VII of Limnology and Oceanography. Limnology and Oceanography, 7(S1):xcii, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:GWS

[Ano62d] Anonymous. Glaciers and winter stoneflies. Limnology and Oceanography, 7(1):109, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:ILO

[Ano62e] Anonymous. Index to Limnology and Oceanography, volume 7, 1962. Limnology and Oceanography, 7(4):491–496, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:IHS

[Ano62f] Anonymous. Information handling and science information. Limnology and Oceanography, 7(3):437, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:IMT

[Ano62g] Anonymous. International meeting on tuna. Limnology and Oceanography, 7(2):271, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1962:NOD

[Ano62h] Anonymous. The National Oceanographic Data Center. Limnology and Oceanography, 7(1):108, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1963:AIA

[Ano63a] Anonymous. Articles of incorporation of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 8 (1):133–134, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1963:LAS

[Ano63b] Anonymous. By-laws of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 8(1):134, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1963:ILO

[Ano63c] Anonymous. Index to Limnology and Oceanography, volume 8, 1963. Limnology and Oceanography, 8(4):495–501, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1963:OF

[Ano63d] Anonymous. Operation freeze-up. Limnology and Oceanography, 8

(3):365, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1963:VIC

[Ano63e] Anonymous. VII INQUA Congress. Limnology and Oceanography, 8

(3):365, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1963:XIB

[Ano63f] Anonymous. X International Botanical Congress—Edinburgh, 1964. Limnology and Oceanography, 8(4):493, April 1963. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1964:A

[Ano64a] Anonymous. Announcements. Limnology and Oceanography, 9(3):

462–465, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1964:ILO

[Ano64b] Anonymous. Index to Limnology and Oceanography, volume 9, 1964. Limnology and Oceanography, 9(4):608–614, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1964:PAS

[Ano64c] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 9(1):159–162, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1964:PAa

[Ano64d] Anonymous. Publication announcement. Limnology and Oceanogra-

phy, 9(2):248, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1964:PAb

[Ano64e] Anonymous. Publication announcement. Limnology and Oceanogra-

phy, 9(2):274, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1964:SV

[Ano64f] Anonymous. Supplement to vol 9. Limnology and Oceanography, 9

(2):i–xxxii, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1965:AL

[Ano65a] Anonymous. Amendment to the by-laws. Limnology and Oceanography, 10(4):617–618, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1965:A

[Ano65b] Anonymous. Announcement. Limnology and Oceanography, 10(3): 496, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1965:ILO

[Ano65c] Anonymous. Index to Limnology and Oceanography, volume 10, 1965. Limnology and Oceanography, 10(4):619–620, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1965:PAS

[Ano65d] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 10(1):162–165, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1965:RLS

[Ano65e] Anonymous. Report of the In Situ Light Measurements Working Group of the Committee on Oceanography of the National Academy of Sciences—National Research Council. Limnology and Oceanography, 10(1):161–162, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1965:SV

[Ano65f] Anonymous. Supplement to vol. 10. Limnology and Oceanography, 10(1):i–xxix, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1966:BRP

[Ano66a] Anonymous. Book review: Panel on Oceanography, President’s Science Advisory Committee. 1966. Effective Use of the Sea. U. S. Government Printing Office, Washington, D. C. xv + 144 p. $0.60. Limnology and Oceanography, 11(4):650–651, April 1966. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1966:ILO

[Ano66b] Anonymous. Index to Limnology and Oceanography volume 11, 1966. Limnology and Oceanography, 11(4):652–659, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1966:ISP

[Ano66c] Anonymous. International symposium on paleolimnology. Limnology and Oceanography, 11(4):651, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1966:NRS

[Ano66d] Anonymous. National register of scientific and technical personnel. Limnology and Oceanography, 11(2):318, April 1966. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1966:OA

[Ano66e] Anonymous. Oceanic abstracts. Limnology and Oceanography, 11(3): 446, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1966:PAS

[Ano66f] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 11(1):138–141, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1966:SNS

[Ano66g] Anonymous. Second national symposium on radioecology. Limnology and Oceanography, 11(3):446, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1966:SMS

[Ano66h] Anonymous. Symposium on mean sea level. Limnology and Oceanography, 11(4):650–651, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:Ab

[Ano67a] Anonymous. Announcement. Limnology and Oceanography, 12(4): 728, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:Aa

[Ano67b] Anonymous. Announcements. Limnology and Oceanography, 12(4): 557–558, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:BRa

[Ano67c] Anonymous. Books received. Limnology and Oceanography, 12(2): 365, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:BRb

[Ano67d] Anonymous. Books received. Limnology and Oceanography, 12(2): 365, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:D

[Ano67e] Anonymous. Deaths. Limnology and Oceanography, 12(4):728, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:Ea

[Ano67f] Anonymous. Erratum. Limnology and Oceanography, 12(4):558, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:Eb

[Ano67g] Anonymous. Erratum. Limnology and Oceanography, 12(4):728, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1967:ICW

[Ano67h] Anonymous. International conference on water pollution research. Limnology and Oceanography, 12(3):187, April 1967. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1967:PAS

[Ano67i] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 12(3):184–186, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1968:BR

[Ano68a] Anonymous. Books received. Limnology and Oceanography, 13(2): 400–401, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1968:ILO

[Ano68b] Anonymous. Index to Limnology and Oceanography volume 13, 1968. Limnology and Oceanography, 13(4):733–739, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1968:PAS

[Ano68c] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 13(1):209–210, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1968:TSP

[Ano68d] Anonymous. Travel support program for Eighth INQUA Congress in Paris. Limnology and Oceanography, 13(4):732, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:A

[Ano69a] Anonymous. Announcement. Limnology and Oceanography, 14(1): 174, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:BRa

[Ano69b] Anonymous. Books received. Limnology and Oceanography, 14(1): 173–174, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:BRb

[Ano69c] Anonymous. Books received. Limnology and Oceanography, 14(4): 651, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:BRc

[Ano69d] Anonymous. Books received. Limnology and Oceanography, 14(6): 968, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:CEP

[Ano69e] Anonymous. Changes in editorial policy. Limnology and Oceanography, 14(6):970, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:ILO

[Ano69f] Anonymous. Index to Limnology and Oceanography volume 14, 1969. Limnology and Oceanography, 14(6):971–978, November 1969. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:ISH

[Ano69g] Anonymous. International symposium on hydrogeochemistry and biogeochemistry and seminar on isotope geochemistry. Limnology and Oceanography, 14(5):816, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:IWP

[Ano69h] Anonymous. International water pollution research conference. Limnology and Oceanography, 14(6):970, November 1969. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1969:PAS

[Ano69i] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 14(1):175–178, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1969:SIL

[Ano69j] Anonymous. Societas internationalis limnologiae. Limnology and Oceanography, 14(2):316, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Anraku:1964:ICCa

[Anr64a] Masateru Anraku. Influence of the Cape Cod Canal on the hydrography and on the copepods in Buzzards Bay and Cape Cod Bay, Massachusetts. I. Hydrography and distribution of copepods. Limnology and Oceanography, 9(1):46–60, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Anraku:1964:ICCb

[Anr64b] Masateru Anraku. Influence of the Cape Cod Canal on the hydrography and on the copepods in Buzzards Bay and Cape Cod Bay, Massachusetts. II. Respiration and feeding. Limnology and Oceanography, 9(2):195–206, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Ansell:1967:EPM

[Ans67] Alan D. Ansell. Egg production of Mercenaria mercenaria. Limnology and Oceanography, 12(3):172–176, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Antal:1966:DTI

[Ant66] Paul S. Antal. Diagenesis of thorium isotopes in deep-sea sediments. Limnology and Oceanography, 11(2):278–292, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Anraku:1963:PSR

[AO63] Masateru Anraku and Makoto Omori. Preliminary survey of the relationship between the feeding habit and the structure of the mouthparts of marine copepods. Limnology and Oceanography, 8(1):116– 126, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Applegate:1966:UBF

[App66] Richard L. Applegate. The use of a bryozoan, Fredericella sultana, as food by sunfish in Bull Shoals Reservoir. Limnology and Oceanography, 11(1):129–130, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Arthur:1967:PSE

[AR67] C. R. Arthur and F. H. Rigler. A possible source of error in the ^{14}C method of measuring primary productivity. Limnology and Oceanography, 12(1):121–124, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Ansell:1963:SMCb

[ARL63a] A. D. Ansell, J. E. G. Raymont, and K. F. Lander. Studies on the mass culture of Phaeodactyylum. III. Small-sscale experiments.

Limnology and Oceanography, 8(2):207–213, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Ansell:1963:SMCa

[ARL^{+}63b] A. D. Ansell, J. E. G. Raymont, K. F. Lander, E. Crowley, and P. Shackley. Studies on the mass culture of phaeodactylum. II. The growth of phaeodactylum and other species in outdoor tanks. Limnology and Oceanography, 8(2):184–206, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Aron:1964:DDD

[ARNA64] William Aron, Newell Raxter, Roy Noel, and William Andrews. A description of a discrete depth plankton sampler with some notes on the towing behavior of a 6-foot Isaacs–Kidd mid-water trawl and a one-meter ring net. Limnology and Oceanography, 9(3):324–333, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Arons:1969:BRB

[Aro69] Arnold Arons. Book review: Behrman, D. 1969. The new world of the oceans. Little, Brown and Co., Boston. 436 p. $8.95. Limnology and Oceanography, 14(6):966–967, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Arthur:1965:SOC

[Art65] Robert S. Arthur. Studies on oceanography — a collection of papers dedicated to Koji Hidaka. Limnology and Oceanography, 10(3):497, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Bayes:1964:SAB

[BA64] J. C. Bayes and A. D. Ansell. A simple automatic bottom water sampler for shore use. Limnology and Oceanography, 9(4):600–601, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Banse:1967:CCC

[BA67] K. Banse and George C. Anderson. Computations of chlorophyll concentrations from spectrophotometric readings. Limnology and Oceanography, 12(4):696–697, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Bamforth:1962:DCS

[Bam62] Stuart S. Bamforth. Diurnal changes in shallow aquatic habitats. Limnology and Oceanography, 7(3):348–353, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Banse:1967:BRM

[Ban67a] K. Banse. Book review: “Meteor” Forschungsergebnisse. (Issued by Deutsche Forschungsgemeinschaft). Reihe A, No. 1, 79 p. Gebriider Borntrager, Berlin-Nikolassee. For subscribers, DM 36 ($9.00). Limnology and Oceanography, 12(3):181, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Banse:1967:BRR

[Ban67b] Karl Banse. Book review: Riedl, R. 1966. Biologie der Meereshöhlen. Paul Parey, Hamburg. 636 p. DM 186. (About $47.00.). Limnology and Oceanography, 12(4):725–726, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Banse:1968:BRM

[Ban68] Karl Banse. Book review: “Meteor” Forschungsergebnisse. 1967. (Issued by Deutsche Forschungsgemeinschaft.) Reihe D, No. 1, 88 p. To subscribers, DM 68.40; to nonsubscribers, DM 95. No. 2, 98 p. To subscribers, DM 66.60; to nonsubscribers, DM 92.50. Gebrüder Bornträger, Berlin-Nikolassee. Limnology and Oceanography, 13(2): 397–398, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1962:CAB

[Bar62] H. Barnes. So-called anecdysis in Balanus balanoides and the effect of breeding upon the growth of the calcareous shell of some common barnacles. Limnology and Oceanography, 7(4):462–473, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1968:BRL

[Bar68] Clifford A. Barnes. Book review: Lauff, G. H. [Ed.]. 1967. Estuaries. Publication No. 83, American Association for the Advancement of Science, Washington, D.C. 757 p. To members $24.00, to nonmembers $27.00. Limnology and Oceanography, 13(4):725, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Barnett:1969:SFK

[Bar69] P. R. O. Barnett. A stabilizing framework for the Knudsen bottom sampler. Limnology and Oceanography, 14(4):648–649, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Boyd:1963:EMB

[BB63] William L. Boyd and Josephine W. Boyd. Enumeration of marine bacteria of the Chukchi Sea. Limnology and Oceanography, 8(3):343– 348, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Barlow:1965:PRZ

[BB65] John P. Barlow and John W. Bishop. Phosphate regeneration by zooplankton in Cayuga Lake. Limnology and Oceanography, 10(S1): R15–R24, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Brock:1967:MCP

[BB67a] Thomas D. Brock and M. Louise Brock. The measurement of chlorophyll, primary productivity, photophosphorylation, and macromolecules in benthic algal mats. Limnology and Oceanography, 12(4): 600–605, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Burkholder:1967:PPS

[BB67b] Paul R. Burkholder and Lillian M. Burkholder. Primary productivity in surface waters of the south Pacific Ocean. Limnology and Oceanography, 12(4):606–617, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Bishop:1968:TSM

[BB68] J. E. Bishop and J. B. Bishop. A technique for selectively marking aquatic insects with ^{32}P. Limnology and Oceanography, 13(4):722– 724, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Baric:1969:BIS

[BB69a] A. Barić and M. Branica. Behavior of indium in seawater (a preliminary study). Limnology and Oceanography, 14(5):796–798, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Brock:1969:ELI

[BB69b] Thomas D. Brock and M. Louise Brock. Effect of light intensity on photosynthesis by thermal algae adapted to natural and reduced sunlight. Limnology and Oceanography, 14(3):334–341, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Brock:1969:FNP

[BB69c] Thomas D. Brock and M. Louise Brock. The fate in nature of photosynthetically assimilated ^{14}C in a blue-green alga. Limnology and Oceanography, 14(4):604–607, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Berner:1962:FSU

[BBG^{+}62] Leo Berner, Jr., Robert Bieri, Edward D. Goldberg, DeCourcey Martin, and Robert L. Wisner. Field studies of uptake of fission products by marine organisms. Limnology and Oceanography, 7(S1):lxxxii–xci, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Bradford:1968:TME

[BBH68] G. R. Bradford, F. L. Bair, and V. Hunsaker. Trace and major element content of 170 High Sierra lakes in California. Limnology and Oceanography, 13(3):526–530, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Baker:1969:SPC

[BBK69] A. L. Baker, A. J. Brook, and A. R. Klemer. Some photosynthetic characteristics of a naturally occurring population of Oscillatoria agardhii Gomont. Limnology and Oceanography, 14(3):327–333, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Buck:1960:SPM

[BC60] John D. Buck and Robert C. Cleverdon. The spread plate as a method for the enumeration of marine bacteria. Limnology and Oceanography, 5(1):78–80, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Buck:1961:ETE

[BC61] John D. Buck and Robert C. Cleverdon. The effect of tween 80 on the enumeration of marine bacteria by the spread and pour plate methods. Limnology and Oceanography, 6(1):42–44, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Brown:1963:OLS

[BC63] Seward Brown and Brian Colman. Oscillaxanthin in lake sediments. Limnology and Oceanography, 8(3):352–353, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Broenkow:1969:CDD

[BC69] William W. Broenkow and Joel D. Cline. Colorimetric determination of dissolved oxygen at low concentrations. Limnology and Oceanography, 14(3):450–454, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Bourbeau:1966:IDD

[BCA66] F. Bourbeau, W. D. Clarke, and W. Aron. Improvements in the discrete depth plankton sampler system. Limnology and Oceanography, 11(3):422–426, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Beadle:1966:PSD

[Bea66] L. C. Beadle. Prolonged stratification and deoxygenation in tropical lakes. I. Crater Lake Nkugute, Uganda, compared with Lakes Bunyoni and Edward. Limnology and Oceanography, 11(2):152–163, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Beard:1969:MLF

[Bea69] M. E. Beard. Mud Lake, Florida; its algae and alkaline brown water. Limnology and Oceanography, 14(6):889–897, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Beeton:1965:ESL

[Bee65] Alfred M. Beeton. Eutrophication of the St. Lawrence Great Lakes. Limnology and Oceanography, 10(2):240–254, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Beers:1966:SCC

[Bee66] John R. Beers. Studies on the chemical composition of the major zooplankton groups in the Sargasso Sea off Bermuda. Limnology and Oceanography, 11(4):520–528, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Bernatowicz:1961:MAE

[Ber61] A. J. Bernatowicz. Marine algae of the Eastern tropical and subtropical coasts of the Americas. Limnology and Oceanography, 6(1): 99–100, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Bercaw:1966:FMT

[Ber66] J. S. Bercaw. A folding midwater trawl depressor. Limnology and Oceanography, 11(4):633–635, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1962:PAA

[BF62] H. Barnes and D. M. Finlayson. Presence of ascorbic acid in cirripede semen. Limnology and Oceanography, 7(1):98–99, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1963:ELA

[BF63] H. Barnes and D. M. Finlayson. Estimation of lactic acid in sea water solutions and homogenates. Limnology and Oceanography, 8 (2):292–294, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Buchan:1967:SSV

[BFC67] S. Buchan, G. D. Floodgate, and D. J. Crisp. Studies on the seasonal variation of the suspended matter in the Menai Straits. I. The inorganic fraction. Limnology and Oceanography, 12(3):419–432, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Bachmann:1964:DMQ

[BG64] Roger W. Bachmann and Charles R. Goldman. The determination of microgram quantities of molybdenum in natural water. Limnology and Oceanography, 9(1):143–146, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Bachmann:1965:HHC

[BG65] Roger W. Bachmann and Charles R. Goldman. Hypolimnetic heating in Castle Lake, California. Limnology and Oceanography, 10(2):233– 239, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Broecker:1969:NRM

[BG69] Wallace S. Broecker and Robert Gerard. Natural radiocarbon in the Mediterranean Sea. Limnology and Oceanography, 14(6):883–888, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Barnett:1967:CAR

[BH67] Arthur M. Barnett and Jed Hirota. Changes in the apparent rate of 14C uptake with length of incubation period in natural phytoplankton populations. Limnology and Oceanography, 12(2):349–353, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Brunskill:1969:FGLc

[BH69] Gregg J. Brunskill and Robert C. Harriss. Fayetteville Green Lake, New York. IV. Interstitial water chemistry of the sediments. Limnology and Oceanography, 14(6):858–861, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Bieri:1962:ZI

[Bie62] Robert Bieri. Zooplankton investigations. Limnology and Oceanography, 7(S1):xxix–xxxi, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Bandy:1964:FAC

[BIR64a] Orville L. Bandy, James C. Ingle, Jr., and Johanna M. Resig. Foraminifera, Los Angeles County Outfall Area, California. Limnology and Oceanography, 9(1):124–137, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Bandy:1964:FTL

[BIR64b] Orville L. Bandy, James C. Ingle, Jr., and Johanna M. Resig. Foraminiferal trends, Laguna Beach outfall area, California. Limnology and Oceanography, 9(1):112–123, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Bandy:1965:FTH

[BIR65] Orville L. Bandy, James C. Ingle, Jr., and Johanna M. Resig. Foraminiferal trends, Hyperion outfall, California. Limnology and Oceanography, 10(3):314–332, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Bishop:1968:RRM

[Bis68] John W. Bishop. Respiratory rates of migrating zooplankton in the natural habitat. Limnology and Oceanography, 13(1):58–62, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Byrne:1962:ILP

[BK62] John V. Byrne and L. D. Kulm. An inexpensive lightweight piston corer. Limnology and Oceanography, 7(1):106–108, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Becacos-Kontos:1968:ACP

[BK68] Theano Becacos-Kontos. The annual cycle of primary production in the Saronicos Gulf (Aegean Sea) for the period November 1963– October 1964. Limnology and Oceanography, 13(3):485–489, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Bowman:1965:BPB

[BL65] Thomas E. Bowman and L. J. Lancaster. A bloom of the planktonic blue–green alga, Trichodesmium erythraeum, in the Tonga Islands. Limnology and Oceanography, 10(2):291–293, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Brunskill:1969:FGLa

[BL69] G. J. Brunskill and S. D. Ludlam. Fayetteville Green Lake, New York. I. Physical and chemical limnology. Limnology and Oceanography, 14 (6):817–829, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Blake:1960:OCB

[Bla60] John W. Blake. Oxygen consumption of bivalve prey and their attractiveness to the gastropod, Urosalpinx cinerea. Limnology and Oceanography, 5(3):273–280, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Barlow:1963:ETE

[BLM63] John P. Barlow, Carl J. Lorenzen, and Richard T. Myren. Eutrophication of a tidal estuary. Limnology and Oceanography, 8(2):251–262, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Buck:1965:AAM

[BM65] John D. Buck and Samuel P. Meyers. Antiyeast activity in the marine environment1 1. ecological considerations. Limnology and Oceanography, 10(3):385–391, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Beers:1968:MSP

[BN68] Gary D. Beers and John M. Neuhold. Measurement of stream periphyton on paraffin-coated substrates. Limnology and Oceanography, 13(3):559–562, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Bachmann:1960:UZP

[BO60] Roger W. Bachmann and Eugene P. Odum. Uptake of Zn^{65} and primary productivity in marine benthic algae. Limnology and Oceanography, 5(4):349–355, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Boden:1968:BRM

[Bod68] Brian P. Boden. Book review: Mauchline, J., and L. R. Fisher 1967. Distribution of the euphausiid crustacean Meganyctiphanes norvegica (M Sars). Serial Atlas of the Marine Environment, Folio 13. American Geographical Society, New York 2 p. + 3 plates. Looseleaf $5.00, bound $9.00. Limnology and Oceanography, 13(4):727–728, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Baxter:1965:STA

[BPTW65] R. M. Baxter, M. V. Prosser, J. F. Talling, and R. B. Wood. Stratification in tropical African lakes at moderate altitudes (1,500 to 2,000 m). Limnology and Oceanography, 10(4):510–520, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Bryson:1960:IWL

[BR60] Reid A. Bryson and Robert A. Ragotzkie. On internal waves in lakes. Limnology and Oceanography, 5(4):397–408, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Berner:1961:RCT

[BR61] Leo D. Berner and Joseph L. Reid, Jr. On the response to changing temperature of the temperature-limited plankter Doliolum denticulatum Quoy and Gaimard 1835. Limnology and Oceanography, 6(2): 205–215, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Bryson:1964:MCT

[BR64] Reid A. Bryson and Robert A. Ragotzkie. “mud-center” tundra lakes. Limnology and Oceanography, 9(1):146–147, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Brooks:1965:TEU

[BR65] Robert R. Brooks and Martin G. Rumsby. Trace element uptake by some New Zealand bivalves. Limnology and Oceanography, 10(4): 521–527, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Burns:1967:CFR

[BR67] Carolyn W. Burns and F. H. Rigler. Comparison of filtering rates of Daphnia rosea in lake water and in suspensions of yeast. Limnology and Oceanography, 12(4):492–502, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Bradshaw:1968:EPM

[Bra68] John S. Bradshaw. Environmental parameters and marsh Foraminifera. Limnology and Oceanography, 13(1):26–38, April 1968. CODEN LI-OCAH. ISSN 0024-3590.

Bradley:1969:VDC

[Bra69] W. H. Bradley. Vertical density currents — II. Limnology and Oceanography, 14(1):1–3, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Brinkhurst:1967:DAO

[Bri67a] Ralph O. Brinkhurst. The distribution of aquatic oligochaetes in Saginaw Bay, Lake Huron. Limnology and Oceanography, 12(1):137– 143, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Brinton:1967:VMA

[Bri67b] Edward Brinton. Vertical migration and avoidance capability of euphausiids in the California current. Limnology and Oceanography, 12 (3):451–483, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Broenkow:1965:DNC

[Bro65a] William W. Broenkow. The distribution of nutrients in the Costa Rica dome in the eastern tropical Pacific Ocean. Limnology and Oceanography, 10(1):40–52, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Brook:1965:PAI

[Bro65b] A. J. Brook. Planktonic algae as indicators of lake types, with special reference to the desmidiaceae. Limnology and Oceanography, 10(3): 403–411, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Brown:1965:DLM

[Bro65c] Robert P. Brown. Delineation of the layer of maximun salinity in tropical and subtropical oceans by means of bathythermograph traces. Limnology and Oceanography, 10(1):157–160, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

Brock:1966:HLM

[Bro66] Thomas D. Brock. The habitat of Leucothrix mucor, a widespread marine microorganism. Limnology and Oceanography, 11(2):303–307, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Brooks:1967:SCV

[Bro67] Albert L. Brooks. Standing crop, vertical distribution, and morphometrics of Ammonia beccarii (Linné). Limnology and Oceanography, 12(4):667–684, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Brock:1968:BRWa

[Bro68a] Thomas D. Brock. Book review: Wood, E. J. F. 1967. 0Microbiology of Oceans and Estuaries. American Elsevier, New York. $23.50. Limnology and Oceanography, 13(1):205, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Brock:1968:BRWb

[Bro68b] Thomas D. Brock. Book review: Wood, E. J. F. 1967. Microbiology of Oceans and Estuaries. American Elsevier, New York. $23.50. Limnology and Oceanography, 13(1):205, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Brown:1968:BCF

[Bro68c] Seward R. Brown. Bacterial carotenoids from freshwater sediments. Limnology and Oceanography, 13(2):233–241, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Broch:1969:OAF

[Bro69a] Edmund S. Broch. The osmotic adaptation of the fairy shrimp Branchinecta campestris Lynch to saline astatic waters. Limnology and Oceanography, 14(4):485–492, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Broenkow:1969:ISU

[Bro69b] William W. Broenkow. An interface sampler using spring-actuated syringes. Limnology and Oceanography, 14(2):288–291, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Bien:1965:RPI

[BRS65] G. S. Bien, N. W. Rakestraw, and H. E. Suess. Radiocarbon in the Pacific and Indian Oceans and its relation to deep water movements. Limnology and Oceanography, 10(S1):R25–R37, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Brunskill:1969:FGLb

[Bru69] G. J. Brunskill. Fayetteville Green Lake, New York. II. Precipitation and sedimentation of calcite in a meromictic lake with laminated sediments. Limnology and Oceanography, 14(6):830–847, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Bedford:1968:FMB

[BRZ68] J. W. Bedford, E. W. Roelofs, and M. J. Zabik. The freshwater mussel as a biological monitor of pesticide concentrations in a lotic environment. Limnology and Oceanography, 13(1):118–126, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Burkholder:1961:PCG

[BS61] Paul R. Burkholder and John M. Sieburth. Phytoplankton and chlorophyll in the Gerlache and Bransfield Straits of Antarctica. Limnology and Oceanography, 6(1):45–52, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Baylor:1963:DOM

[BS63] E. R. Baylor and W. H. Sutcliffe, Jr. Dissolved organic matter in seawater as a source of particulate food. Limnology and Oceanography, 8(4):369–371, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Bigelow:1965:NSC

[BS65] Henry B. Bigelow and William C. Schroeder. Notes on a small collection of rajids from the sub-Antarctica region. Limnology and Oceanography, 10(S1):R38–R49, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Bugbee:1967:PVP

[Bug67] Stephen L. Bugbee. A portable vacuum pump for ^{14}C productivity studies. Limnology and Oceanography, 12(4):707–709, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Bumpus:1965:RDA

[Bum65] Dean F. Bumpus. Residual drift along the bottom on the continental shelf in the Middle Atlantic Bight area. Limnology and Oceanography, 10(S1):R50–R53, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Burke:1968:SCD

[Bur68a] John C. Burke. A sediment coring device of 21-cm diameter with sphincter core retainer. Limnology and Oceanography, 13(4):714–718, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Burns:1968:RBB

[Bur68b] Carolyn W. Burns. The relationship between body size of filterfeeding Cladocera and the maximum size of particle ingested. Limnology and Oceanography, 13(4):675–678, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Burns:1969:PSS

[Bur69a] Carolyn W. Burns. Particle size and sedimentation in the feeding behavior of two species of Daphnia. Limnology and Oceanography, 14(3):392–402, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Burns:1969:RBF

[Bur69b] Carolyn W. Burns. Relation between filtering rate, temperature, and body size in four species of Daphnia. Limnology and Oceanography, 14(5):693–700, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Button:1969:ECA

[But69] D. K. Button. Effect of clay on the availability of dilute organic nutrients to steady-state heterotrophic populations. Limnology and Oceanography, 14(1):95–100, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Buzas:1969:FSD

[Buz69] Martin A. Buzas. Foraminiferal species densities and environmental variables in an estuary. Limnology and Oceanography, 14(3):411–422, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Baird:1968:MDZ

[BW68a] I. E. Baird and R. G. Wetzel. A method for the determination of zero thickness activity of ^{14}C labeled benthic diatoms in sand. Limnology and Oceanography, 13(2):379–382, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1968:SVS

[BW68b] Nora B. Barnes and Adrian M. Wenner. Seasonal variation in the sand crab Emerita analoga (Decapoda, Hippidae) in the Santa Barbara area of California. Limnology and Oceanography, 13(3):465–475, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Biggs:1968:CPC

[BW68c] Robert B. Biggs and Carolyn D. Wetzel. Concentration of particulate carbohydrateatthe halocline in Chesapeake Bay. Limnology and Oceanography, 13(1):169–171, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Broch:1969:MMI

[BY69] Edmund S. Broch and William Yake. A modification of Maucha’s ionic diagram to include ionic concentrations. Limnology and Oceanography, 14(6):933–935, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Bye:1965:WDC

[Bye65] John A. T. Bye. Wind-driven circulation in unstratified lakes. Limnology and Oceanography, 10(3):451–458, April 1965. CODEN LI-OCAH. ISSN 0024-3590.

Ben-Yaakov:1968:PTP

[BYK68] S. Ben-Yaakov and I. R. Kaplan. pH–temperature profiles in ocean and lakes using an In Situ probe. Limnology and Oceanography, 13 (4):688–693, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Ben-Yaakov:1969:DCS

[BYK69] S. Ben-Yaakov and I. R. Kaplan. Determination of carbonate saturation of seawater with a carbonate saturometer. Limnology and Oceanography, 14(6):874–882, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Byrne:1965:MCL

[Byr65] John V. Byrne. Morphometry of Crater Lake, Oregon. Limnology and Oceanography, 10(3):462–465, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Coachman:1966:WET

[CA66] L. K. Coachman and K. Aagaard. On the water exchange through Bering Strait. Limnology and Oceanography, 11(1):44–59, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Carriker:1961:TCN

[Car61] Melbourn R. Carriker. Thurlow Christian Nelson. Limnology and Oceanography, 6(1):79, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1965:AWM

[Car65a] James H. Carpenter. The accuracy of the Winkler method for dissolved oxygen analysis. Limnology and Oceanography, 10(1):135–140, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1965:CBI

[Car65b] James H. Carpenter. The Chesapeake Bay Institute technique for the Winkler dissolved oxygen method. Limnology and Oceanography, 10(1):141–143, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Carter:1965:ECC

[Car65c] John C. H. Carter. The ecology of the calanoid copepod Pseudocalanus minutus Krøyer in Tessiarsuk, a coastal meromictic lake of northern Labrador. Limnology and Oceanography, 10(3):345–353, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1966:NMO

[Car66] James H. Carpenter. New measurements of oxygen solubility in pure and natural water. Limnology and Oceanography, 11(2):264–277, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Carpelan:1968:BRW

[Car68a] Lars H. Carpelan. Book review: Weatherley, A. H. [Ed.] 1967. Australian Inland Waters and Their Fauna. Eleven studies. Australian National University Press, Canberra. xv + 287 p. $A10.50. Limnology and Oceanography, 13(2):395, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1968:SIA

[Car68b] Edward J. Carpenter. A simple, inexpensive algal chemostat. Limnology and Oceanography, 13(4):720–721, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Castenholz:1960:SCA

[Cas60] Richard W. Castenholz. Seasonal changes in the attached algae of freshwater and saline lakes in the Lower Grand Coulee, Washington. Limnology and Oceanography, 5(1):1–28, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Cassie:1962:MOE

[Cas62] R. Morrison Cassie. Microdistribution and other error components of C^{14} primary production estimates. Limnology and Oceanography, 7(2):121–130, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Cassie:1963:RBP

[Cas63a] R. Morrison Cassie. Relationship between plant pigments and gross primary production in Skeletonema costatum. Limnology and Oceanography, 8(4):433–439, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Castenholz:1963:ESV

[Cas63b] Richard W. Castenholz. An experimental study of the vertical distribution of littoral marine diatoms. Limnology and Oceanography, 8 (4):450–462, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Castagna:1967:BSD

[Cas67] Michael Castagna. A benthic sampling device for shallow water. Limnology and Oceanography, 12(2):357–359, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Costlow:1960:MDB

[CB60] John D. Costlow, Jr. and C. G. Bookhout. A method for developing brachyuran eggs In vitro. Limnology and Oceanography, 5(3):212– 215, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Clark:1967:DPM

[CB67] R. C. Clark, Jr. and M. Blumer. Distribution of n-paraffins in marine organisms and sediment. Limnology and Oceanography, 12(1):79–87, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Colby:1969:CCA

[CB69a] Peter J. Colby and Larry T. Brooke. Cisco (Coregonus artedii) mortalities in a southern Michigan lake, July 1968. Limnology and Oceanography, 14(6):958–960, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Culver:1969:FGL

[CB69b] D. A. Culver and G. J. Brunskill. Fayetteville Green Lake, New York. V. Studies of primary production and zooplankton in a meromictic marl lake. Limnology and Oceanography, 14(6):862–873, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Carsola:1962:CTS

[CC62a] Alfred J. Carsola and Elliot B. Callaway. Correction: Two shortperiod internal wave frequency spectra in the sea off southern California [Limnol. Oceanogr. 7(2) 115–120 (1962)]. Limnology and Oceanography, 7(4):490, April 1962. CODEN LIOCAH. ISSN 0024-3590. See [CC62b].

Carsola:1962:TSP

[CC62b] Alfred J. Carsola and Elliott B. Callaway. Two short-period internal wave frequency spectra in the sea off southern California. Limnology and Oceanography, 7(2):115–120, April 1962. CODEN LIOCAH. ISSN 0024-3590. See correction [CC62a].

Copeland:1964:CME

[CD64a] B. J. Copeland and Troy C. Dorris. Community metabolism in ecosystems receiving oil refinery effluents. Limnology and Oceanography, 9(3):431–447, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Copeland:1964:UCP

[CD64b] B. J. Copeland and W. R. Duffer. Use of a clear plastic dome to measure gaseous diffusion rates in natural waters. Limnology and Oceanography, 9(4):494–499, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Curl:1967:ICS

[CD67] Herbert Curl, Jr. and Earl W. Davey. Improved calibration and sample-injection systems for nondestructive analysis of permanent gases, total CO_{2}, and dissolved organic carbon in water. Limnology and Oceanography, 12(4):545–548, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Chew:1962:SRD

[CDD62] Frank Chew, K. L. Drennan, and W. J. Demoran. Some results of drift bottle studies off the Mississippi Delta. Limnology and Oceanography, 7(2):252–257, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Collier:1968:TS

[CFC68] Albert Collier, Gale M. Finlayson, and Edwin W. Cake. On the transparency of the sea. Limnology and Oceanography, 13(2):391– 394, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Crandall:1962:ESC

[CG62] Catherine A. Crandall and Clarence J. Goodnight. Effects of sublethal concentrations of several toxicants on growth of the common guppy, Lebistes reticulatus. Limnology and Oceanography, 7(2):233– 239, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Carr:1965:CBF

[CH65] John F. Carr and Jarl K. Hiltunen. Changes in the bottom fauna of western Lake Erie from 1930 to 1961. Limnology and Oceanography, 10(4):551–569, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Cowell:1968:DDS

[CH68] Bruce C. Cowell and Patrick L. Hudson. A device for determining sedimentation rates in reservoirs. Limnology and Oceanography, 13 (1):196–198, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Chave:1962:FIM

[Cha62] Keith E. Chave. Factors influencing the mineralogy of carbonate sediments. Limnology and Oceanography, 7(2):218–223, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Charlier:1965:CPO

[Cha65] Roger H. Charlier. La connaissance des profondeurs oceaniques. Limnology and Oceanography, 10(2):303, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Chapman:1966:BRC

[Cha66a] W. M. Chapman. Book review: Crutchfield, J. A. [ed.]. 1965. The Fisheries Problems in Resource Management. University of Washington Press, Seattle 98105. 136 p. $5.00. Limnology and Oceanography, 11(2):313, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Charlier:1966:BRB

[Cha66b] Roger H. Charlier. Book review: Bourcart, J. 19518. Problèmes de Géologie sous-marine. Le précontinent — Le littoral et sa protection-La stratigraphie sous-marine. Masson and Cie, Paris. 125 p. $1.80. Limnology and Oceanography, 11(2):316–317, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Charlier:1966:GOE

[Cha66c] Roger H. Charlier. Growth of oceanographic education in the United States. Limnology and Oceanography, 11(4):636–640, April 1966. CO-DEN LIOCAH. ISSN 0024-3590.

Chaston:1969:ACV

[Cha69] I. Chaston. Anaerobiosis in Cyclops varicans. Limnology and Oceanography, 14(2):298–301, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Chew:1967:CSV

[Che67] Frank Chew. On the cross-stream variation of the k-factor for geomagnetic electrokinetograph data from the Florida Current off Miami. Limnology and Oceanography, 12(1):73–78, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Chesher:1968:TMP

[Che68] Richard H. Chesher. Transport of marine plankton through the Panama Canal. Limnology and Oceanography, 13(2):387–388, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Chitwood:1960:FLN

[Chi60] B. G. Chitwood. Free living nematodes and other small invertebrates of Puget Sound beaches. Limnology and Oceanography, 5(1):112–114, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Clarke:1965:MDC

[CK65] George L. Clarke and Mahlon G. Kelly. Measurements of diurnal changes in bioluminescence from the sea surface to 2,000 meters using a new photometric device. Limnology and Oceanography, 10(S1): R54–R66, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Craig:1962:ULI

[CL62] R. E. Craig and R. G. Lawrie. An underwater light intensity meter. Limnology and Oceanography, 7(2):259–261, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Clark:1962:OFN

[Cla62] R. B. Clark. Observations on the food of Nephtys. Limnology and Oceanography, 7(3):380–385, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Cline:1969:SDH

[Cli69] Joel D. Cline. Spectrophotometric determination of hydrogen sulfide in natural waters. Limnology and Oceanography, 14(3):454–458, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Clutter:1965:SCD

[Clu65] Robert I. Clutter. Self-closing device for sampling plankton near the sea bottom. Limnology and Oceanography, 10(2):293–296, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Carlucci:1969:NMB

[CM69] A. F. Carlucci and P. M. McNally. Nitrification by marine bacteria in low concentrations of substrate and oxygen. Limnology and Oceanography, 14(5):736–739, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Copeland:1964:CSO

[CMD64] B. J. Copeland, K. W. Minter, and Troy C. Dorris. Chlorophyll-a and suspended organic matter in oil refinery effluent holding ponds. Limnology and Oceanography, 9(4):500–506, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Cochrane:1963:OOP

[Coc63] John D. Cochrane. Oceanographic observations of the Pacific: 1955, the NORPAC Atlas prepared by the NORPAC Committee. Limnology and Oceanography, 8(3):366–367, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Cole:1961:SCC

[Col61] Gerald A. Cole. Some Calanoid copepods from Arizona with notes on congeneric occurrences of Diaptomus species. Limnology and Oceanography, 6(4):432–442, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Colinvaux:1964:SSS

[Col64] Paul A. Colinvaux. Sampling stiff sediments of an ice-covered lake. Limnology and Oceanography, 9(2):262–264, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Coles:1969:QEF

[Col69] Stephen L. Coles. Quantitative estimates of feeding and respiration for three scleractinian corals. Limnology and Oceanography, 14(6): 949–953, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Comita:1965:OUD

[Com65] Gabriel W. Comita. Oxygen uptake in Diaptomus siciloides Lilljeborg. Limnology and Oceanography, 10(3):466–468, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

Comita:1968:OCD

[Com68] Gabriel W. Comita. Oxygen consumption in Diaptomus. Limnology and Oceanography, 13(1):51–57, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Connell:1964:RPC

[Con64] J. H. Connell. Readings in population and community ecology. Limnology and Oceanography, 9(3):466, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Conover:1966:AOM

[Con66a] Robert J. Conover. Assimilation of organic matter by zooplankton. Limnology and Oceanography, 11(3):338–345, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Conover:1966:FAA

[Con66b] Robert J. Conover. Factors affecting the assimilation of organic matter by zooplankton and the question of superfluous feeding. Limnology and Oceanography, 11(3):346–354, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Correll:1965:PPM

[Cor65] David L. Correll. Pelagic phosphorus metabolism in Antarctic waters. Limnology and Oceanography, 10(3):364–370, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Cory:1967:DEL

[Cor67] Robert L. Cory. Distribution and ecology of lancelets (order Amphioxi) over the continental shelf of the southeastern United States. Limnology and Oceanography, 12(4):650–656, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Coulter:1968:TSD

[Cou68] G. W. Coulter. Thermal stratification in the deep hypolimnion of Lake Tanganyika. Limnology and Oceanography, 13(2):385–387, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Coull:1969:HCM

[Cou69] Bruce C. Coull. Hydrographic control of meiobenthos in Bermuda. Limnology and Oceanography, 14(6):953–957, November 1969. CO-DEN LIOCAH. ISSN 0024-3590.

Cowell:1967:CCM

[Cow67] Bruce C. Cowell. The Copepoda and Cladocera of a Missouri River reservoir: a comparison of sampling in the reservoir and the discharge. Limnology and Oceanography, 12(1):125–136, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Crisp:1961:IBB

[CP61] D. J. Crisp and Bhupendra Patel. The interaction between breeding and growth rate in the barnacle Elminius modestus Darwin. Limnology and Oceanography, 6(2):105–115, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Carey:1968:MSM

[CP68a] Andrew G. Carey, Jr. and Roger R. Paul. A modification of the Smith-McIntyre grab for simultaneous collection of sediment and bottom water. Limnology and Oceanography, 13(3):545–549, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Culberson:1968:EPC

[CP68b] C. Culberson and R. M. Pytkowicx. Effect of pressure on carbonic acid, boric acid, and the pH in seawater. Limnology and Oceanography, 13(3):403–417, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Clark:1963:MCP

[CS63] William J. Clark and William F. Sigler. Method of concentrating phytoplankton samples using membranefilters. Limnology and Oceanography, 8(1):127–129, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Curl:1965:VPA

[CS65] Herbert Curl, Jr. and Lawrence F. Small. Variations in photosynthetic assimilation ratios in natural, marine phytoplankton communities. Limnology and Oceanography, 10(S1):R67–R73, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Carlucci:1966:BSI

[CS66] A. F. Carlucci and S. B. Silbernagel. Bioassay of seawater. III. Distribution of vitamin B_{1}2 in the northeast Pacific Ocean. Limnology and Oceanography, 11(4):642–646, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Cross:1967:CIS

[CS67] Ford A. Cross and Lawrence F. Small. Copepod indicators of surface water movements off the Oregon coast. Limnology and Oceanography, 12(1):60–73, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Carlucci:1969:NRS

[CS69a] A. F. Carlucci and Hazel R. Schubert. Nitrate reduction in seawater of the deep nitrite maximum off Peru. Limnology and Oceanography, 14(2):187–193, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Chamberlain:1969:BSP

[CS69b] William Chamberlain and Joseph Shapiro. On the biological significance of phosphate analysis; comparison of standard and new methods with a bioassay. Limnology and Oceanography, 14(6):921–927, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Clark:1969:TEF

[CS69c] Shirley M. Clark and George R. Snyder. Timing and extent of a flow reversal in the Lower Columbia River. Limnology and Oceanography, 14(6):960–965, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Colton:1961:EGB

[CT61] John B. Colton, Jr. and Robert F. Temple. The enigma of Georges Bank spawning. Limnology and Oceanography, 6(3):280–291, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Cole:1965:SSC

[CU65] Gerald A. Cole and James C. Underhill. The summer standing crop of sublittoral and profundal benthos in Lake Itasca, Minnesota. Limnology and Oceanography, 10(4):591–597, April 1965. CODEN LI-OCAH. ISSN 0024-3590.

Curl:1960:OVF

[Cur60] Herbert Curl, Jr. Oceanography vs. fisheries. Limnology and Oceanography, 5(1):106, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Curl:1962:EDS

[Cur62] Herbert Curl, Jr. Effect of divalent sulfur and vitamin B_{12} in controlling the distribution of Skeletonema costatum. Limnology and Oceanography, 7(3):422–424, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Cushing:1967:CTC

[Cus67] C. E. Cushing. Concentration and transport of ^{32}P and ^{65}Zn by Columbia River plankton. Limnology and Oceanography, 12(2):330– 332, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Connors:1968:PEC

[CW68] Donald N. Connors and Peter K. Weyl. The partial equivalent conductances of salts in seawater and the density/conductance relationship. Limnology and Oceanography, 13(1):39–50, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Cole:1967:UMT

[CWB67] Gerald A. Cole, Melbourne C. Whiteside, and Robert J. Brown. Unusual monomixis in two saline Arizona ponds. Limnology and Oceanography, 12(4):584–591, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Daisley:1969:MSV

[Dai69] K. W. Daisley. Monthly survey of vitamin B_{1}2 concentrations in some waters of the English Lake District. Limnology and Oceanography, 14(2):224–228, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Davis:1962:EMN

[Dav62] Charles C. Davis. Ecological and morphological notes on Hydrophylita aquivalans (Math. and Crosby). Limnology and Oceanography, 7(3):390–392, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Davis:1964:EEL

[Dav64] Charles C. Davis. Evidence for the eutrophication of Lake Erie from phytoplankton records. Limnology and Oceanography, 9(3):275–283, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Dawson:1960:HMC

[Daw60] William A. Dawson. Home-made counting chambers for the inverted microscope. Limnology and Oceanography, 5(3):235–236, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Dutton:1962:HFL

[DB62] John A. Dutton and Reid A. Bryson. Heat flux in Lake Mendota. Limnology and Oceanography, 7(1):80–97, April 1962. CODEN LI-OCAH. ISSN 0024-3590.

Doe:1965:MSP

[DB65] L. A. E. Doe and John Brooke. A moored stable platform for air–sea interaction studies. Limnology and Oceanography, 10(S1):R79–R86, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Davies:1969:LSS

[DB69] I. E. Davies and E. G. Barham. An In Situ surge–temperature recorder. Limnology and Oceanography, 14(4):638–641, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Dorris:1962:LMM

[DC62] Troy C. Dorris and B. J. Copeland. Limnology of the Middle Mississippi River. III. Mayfly populations in relation to navigation waterlevel control. Limnology and Oceanography, 7(2):240–247, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Dorris:1963:LMM

[DCL63] Troy C. Dorris, B. J. Copeland, and Gerald J. Lauer. Limnology of the Middle Mississippi River. IV. Physical and chemical limnology of river and chute. Limnology and Oceanography, 8(1):79–88, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1961:SPN

[DD61] Richard C. Dugdale and Vera A. Dugdale. Sources of phosphorus and nitrogen for lakes on Afognak Island. Limnology and Oceanography, 6(1):13–23, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1962:NML

[DD62] Vera A. Dugdale and Richard C. Dugdale. Nitrogen metabolism in lakes II. Role of nitrogen fixation in Sanctuary Lake, Pennsylvania. Limnology and Oceanography, 7(2):170–177, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1965:NML

[DD65] Vera A. Dugdale and Richard C. Dugdale. Nitrogen metabolism in lakes III. Tracer studies of the assimilation of inorganic nitrogen sources. Limnology and Oceanography, 10(1):53–57, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

Duffer:1966:PPS

[DD66] William R. Duffer and Troy C. Dorris. Primary productivity in a southern Great Plains stream. Limnology and Oceanography, 11(2): 143–151, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Davis:1969:PCU

[DD69] Ronald B. Davis and Roger W. Doyle. A piston corer for upper sediment in lakes. Limnology and Oceanography, 14(4):643–648, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Deane:1963:LMO

[Dea63] Roger E. Deane. Limnological and meteorological observation towers in the Great Lakes. Limnology and Oceanography, 8(1):9–15, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1967:BRP

[Dee67] E. S. Deevey. Book review: Patten, D. W. 1966. The Biblical flood and the ice epoch: a study in scientific history. Pacific Meridian Publ. Co., Seattle. 336 p. $7.50. Limnology and Oceanography, 12(4):724– 725, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1969:BRM

[Dee69] E. S. Deevey. Book review: Margalef, R. 1968. Perspectives in ecological theory. Univ. Chicago Press, Chicago, Ill. 111 p. $4.50. Limnology and Oceanography, 14(2):313–315, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Deffeyes:1965:CEG

[Def65] Kenneth S. Deffeyes. Carbonate equilibria: a graphic and algebraic approach. Limnology and Oceanography, 10(3):412–426, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Dendy:1963:OBE

[Den63] J. S. Dendy. Observations on bryozoan ecology in farm ponds. Limnology and Oceanography, 8(4):478–482, April 1963. CODEN LI-OCAH. ISSN 0024-3590.

Denoyelles:1968:SOF

[Den68] Frank Denoyelles, Jr. A stained-organism filter technique for concentrating phytoplankton. Limnology and Oceanography, 13(3):562–565, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Desrosieres:1969:SMP

[Des69] Roger Desrosieres. Surface macrophytoplankton of the Pacific Ocean along the equator. Limnology and Oceanography, 14(4):626–632, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1967:UNR

[DG67] R. C. Dugdale and J. J. Goering. Uptake of new and regenerated forms of nitrogen in primary productivity. Limnology and Oceanography, 12(2):196–206, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Dera:1968:LFF

[DG68] Jerzy Dera and Howard R. Gordon. Light field fluctuations in the photic zone. Limnology and Oceanography, 13(4):697–699, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1964:HNF

[DGR64] Richard C. Dugdale, John J. Goering, and John H. Ryther. High nitrogen fixation rates in the Sargasso Sea and the Arabian Sea. Limnology and Oceanography, 9(4):507–510, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Dean:1964:BRR

[DH64] David Dean and Harold H. Haskin. Benthic repopulation of the Raritan River estuary following pollution abatement. Limnology and Oceanography, 9(4):551–563, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Dickie:1968:BRC

[Dic68] L. M. Dickie. Book review: Cushing, D. H. 1968. Fisheries biology: a study in population dynamics. University of Wisconsin Press, Madison. 200 p. $7.50. Limnology and Oceanography, 13(4):729–730, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Dickman:1969:SEL

[Dic69] Mike Dickman. Some effects of lake renewal on phytoplankton productivity and species composition. Limnology and Oceanography, 14 (5):660–666, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Dillon:1964:FTS

[Dil64] William P. Dillon. Flotation technique for separating fecal pellets and small marine organisms from sand. Limnology and Oceanography, 9 (4):601–602, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Doty:1965:IMP

[DJKMS65] Maxwell S. Doty, Harry R. Jitts, Olga J. Koblentz-Mishke, and Yatsuke Saijo. Intercalibration of marine plankton primary productivity techniques. Limnology and Oceanography, 10(2):282–286, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Duxbury:1968:LCS

[DM68] Alyn C. Duxbury and Noel B. McGary. Local changes of salinity and nutrients off the mouth of the Columbia River. Limnology and Oceanography, 13(4):626–636, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Dorris:1961:PSD

[Dor61] Troy C. Dorris. A plankton sampler for deep river waters. Limnology and Oceanography, 6(3):366–367, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Doyle:1968:IEO

[Doy68] Roger W. Doyle. Isolation of enzymes and other organic anions from freshwater sediment. Limnology and Oceanography, 13(3):518–522, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Dyer:1969:SWC

[DR69] K. R. Dyer and K. Ramamoorthy. Salinity and water circulation in the vellar estuary. Limnology and Oceanography, 14(1):4–15, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Driver:1964:SSS

[Dri64] E. A. Driver. A simple speedometer for small boats. Limnology and Oceanography, 9(2):264–265, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Driscoll:1967:AES

[Dri67] Egbert G. Driscoll. Attached epifauna–substrate relations. Limnology and Oceanography, 12(4):633–641, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1964:DNI

[DS64] Edward S. Deevey, Jr. and Minze Stuiver. Distribution of natural isotopes of carbon in Linsley Pond and other New England lakes. Limnology and Oceanography, 9(1):1–11, April 1964. CODEN LI-OCAH. ISSN 0024-3590.

Duthie:1968:CAM

[DS68] H. C. Duthie and J. Shaw. A compressed air motor shaker for use in small incubators. Limnology and Oceanography, 13(4):719–720, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1964:CGF

[DT64] Anne N. Dodson and William H. Thomas. Concentratingplankton in a gentle fashion. Limnology and Oceanography, 9(3):455–456, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Duff:1965:TCG

[DT65] Sheila Duff and John M. Teal. Temperature change and gas exchange in Nova Scotia and Georgia salt-marsh muds. Limnology and Oceanography, 10(1):67–73, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1967:NLS

[Dug67] R. C. Dugdale. Nutrient limitation in the sea: dynamics, identification, and significance. Limnology and Oceanography, 12(4):685–695, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1968:BRG

[Dug68] Richard C. Dugdale. Book review: Golterman, H. L. [Ed.]. 1967. Chemical environment in the aquatic habitat. North Holland Publishing Co., Amsterdam. 322 p. $8.40. Limnology and Oceanography, 13(4):730, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Dumont:1969:QMS

[Dum69] Henri J. Dumont. A quantitative method for the study of periphyton. Limnology and Oceanography, 14(2):303–307, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Dunn:1967:DFP

[Dun67] I. G. Dunn. Diurnal fluctuations of physicochemical conditions in a shallow tropical pond. Limnology and Oceanography, 12(3):151–154, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Duxbury:1962:ADT

[Dux62] Alyn C. Duxbury. Averaged dynamic topographies of the Gulf of Mexico. Limnology and Oceanography, 7(3):428–431, April 1962. CO-DEN LIOCAH. ISSN 0024-3590.

Duxbury:1963:CUG

[Dux63] Alyn C. Duxbury. Calibration and use of a galvanic type oxygen electrode in field work. Limnology and Oceanography, 8(4):483–485, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1960:LDB

[DW60] Richard C. Dugdale and James T. Wallace. Light and dark bottle experiments in Alaska. Limnology and Oceanography, 5(3):230–231, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Dawnport:1965:NRB

[DW65] Demorest Dawnport and Maryanna Warmuth. Notes on the relationship between the freshwater mussel Anodonta implicata Say and the alewife Pomolbus pseudoharengus (Wilson). Limnology and Oceanography, 10(S1):R74–R78, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Duedall:1967:PEV

[DW67] Iver W. Duedall and Peter K. Weyl. The partial equivalent volumes of salts in seawater. Limnology and Oceanography, 12(1):52–59, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Dukes:1968:TSE

[DWW68] Thomas W. Dukes, James N. Willis, and Douglas A. Wolfe. A technique for studying the exchange of trace elements between estuarine sediments and water. Limnology and Oceanography, 13(3):541–545, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1965:SFS

[EA65] W. T. Edmondson and George C. Anderson. Some features of saline lakes in central Washington. Limnology and Oceanography, 10(S1): R87–R96, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1969:SNR

[ECS69] R. W. Eppley, J. L. Coatsworth, and Lucia Solórzano. Studies of nitrate reductase in marine phytoplankton. Limnology and Oceanography, 14(2):194–205, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1966:BRJ

[Edm66a] W. T. Edmondson. Book review: Jørgensen, C. B. 1966. Biology of Suspension Feeding. Pergamon, New York. 357 p. Limnology and Oceanography, 11(4):649–650, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1966:BRS

[Edm66b] W. T. Edmondson. Book review: Schwoerbel, J. 1966. Methoden der Hydrohiologie. Kosmos, Stuttgart. 207 p. DM29.50. Limnology and Oceanography, 11(4):649, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1967:BRD

[Edm67] W. T. Edmondson. Book review: Dussart, Bernard. 1966. Limnologie: l’Étude des Eaux Continentales. Paris, Gauthier-Villars. Xxiv + 677 p., illus. $27.60. Limnology and Oceanography, 12(3):181–182, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1968:BRG

[Edm68a] W. T. Edmondson. Book review: Goldman, C. R. [Ed.]. 1966. Primary Productivity in Aquatic Environments. Univ. Calif. Press, Berkeley. 464 p. $7.00. Limnology and Oceanography, 13(1):206–207, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1968:BRIa

[Edm68b] W. T. Edmondson. Book review: Illies, J. [Ed.]. 1967. Limnofauna Europea. Fischer, Stuttgart. xv + 474 p. Limnology and Oceanography, 13(1):208, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1968:BRIb

[Edm68c] W. T. Edmondson. Book review: Illies, J. [Ed.]. 1967. Limnofauna Europea. Fischer, Stuttgart. xv + 474 p. Limnology and Oceanography, 13(1):208, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1969:BRR

[Edm69] W. T. Edmondson. Book review: Reeves, C. C., Jr. 1968. Introduction to paleolimnology. Developments in sedimentology No. 11. American Elsevier, New York. 228 p. $18.00. Limnology and Oceanography, 14(4):650–651, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Egerton:1964:BJA

[Ege64] Frank N. Egerton. Birge and Juday archives. Limnology and Oceanography, 9(1):155, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Eltringham:1961:MRW

[EH61] S. K. Eltringham and A. R. Hockley. Migration and reproduction of the wood-boring isopod, Limnoria, in Southampton Water. Limnology and Oceanography, 6(4):467–481, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Emery:1962:ITC

[EHR62] K. O. Emery, Jobst Hülsemann, and K. S. Rodolfo. Influence of turbidity currents upon basin waters. Limnology and Oceanography, 7(4):439–446, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Einarsson:1968:SSL

[EL68] E. Einarsson and A. B. Lowe. Seiches and set-up on Lake Winnipeg. Limnology and Oceanography, 13(2):257–271, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Ellis:1966:APH

[Ell66] D. V. Ellis. Aerial photography from helicopter as a technique for intertidal surveys. Limnology and Oceanography, 11(2):299–301, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1963:RAC

[EM63] Richard W. Eppley and Frank M. MaciasR. Role of the alga Chlamydomonas mundana in anaerobic waste stabilization lagoons. Limnology and Oceanography, 8(4):411–416, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Eaton:1966:ENP

[EM66] John W. Eaton and Brian Moss. The estimation of numbers and pigment content in epipelic algal populations. Limnology and Oceanography, 11(4):584–595, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Emery:1961:SMM

[Eme61] K. O. Emery. A simple method of measuring beach profiles. Limnology and Oceanography, 6(1):90–93, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Emery:1968:POC

[Eme68] Alan R. Emery. Preliminary observations on coral reef plankton. Limnology and Oceanography, 13(2):293–303, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Emery:1965:GBS

[EMT65] K. O. Emery, Arthur S. Merrill, and James V. A. Trumbull. Geology and biology of the sea floor as deduced from simultaneous photographs and samples. Limnology and Oceanography, 10(1):1–21, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Enright:1963:ECS

[Enr63] J. T. Enright. Estimates of the compressibility of some marine crustaceans. Limnology and Oceanography, 8(4):382–387, April 1963. CO-DEN LIOCAH. ISSN 0024-3590.

Eppley:1968:IME

[Epp68] Richard W. Eppley. An incubation method for estimating the carbon content of phytoplankton in natural samples. Limnology and Oceanography, 13(4):574–582, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1969:HSC

[ERM69] Richard W. Eppley, Jane N. Rogers, and James J. McCarthy. Halfsaturation constants for uptake of nitrate and ammonium by marine phytoplankton. Limnology and Oceanography, 14(6):912–920, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Everest:1967:MBC

[Eve67] Fred H. Everest. Midget Bentzel Current speed tube for ecological investigations. Limnology and Oceanography, 12(3):179–180, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Emery:1965:SPD

[EWR65] K. O. Emery, R. L. Wigley, and Meyer Rubin. A submerged peat deposit off the Atlantic coast of the United States. Limnology and Oceanography, 10(S1):R97–R102, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Franklin:1961:BSS

[FA61] W. R. Franklin and D. V. Anderson. A bottom sediment sampler. Limnology and Oceanography, 6(2):233–234, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Fager:1968:SBE

[Fag68] Edward W. Fager. A sand-bottom epifaunal community of invertebrates in shallow water. Limnology and Oceanography, 13(3):448–464, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Fager:1969:PSB

[Fag69] E. W. Fager. Production of stream benthos: a critique of the method of assessment proposed by Hynes and Coleman (1968). Limnology and Oceanography, 14(5):766–770, September 1969. CODEN LI-OCAH. ISSN 0024-3590. See [HC68b].

Faller:1969:GLC

[Fal69] Alan J. Faller. The generation of Langmuir circulations by the eddy pressure of surface waves. Limnology and Oceanography, 14(4):504– 513, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Fell:1960:IYB

[FAMR60] Jack W. Fell, Donald G. Ahearn, Samuel P. Meyers, and Frank J. Roth, Jr. Isolation of yeasts from Biscayne Bay, Florida and adjacent benthic areas. Limnology and Oceanography, 5(4):366–371, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Farris:1960:FAH

[Far60] David A. Farris. Failure of an anchovy to hatch with continued growth of the larva. Limnology and Oceanography, 5(1):107, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Fee:1968:ESD

[FB68] Everett J. Fee and Roger W. Bachmann. An empirical study of the defant method of seiche analysis. Limnology and Oceanography, 13 (4):665–669, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Freudenthal:1968:UPE

[FBS68] Hugo D. Freudenthal, Walter Blogoslowski, and Roy Stoecker. The use of the pneumatic ejector pump in oceanography. Limnology and Oceanography, 13(4):706–708, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Fleminger:1965:ATN

[FC65] Abraham Fleminger and Robert I. Clutter. Avoidance of towed nets by zooplankton. Limnology and Oceanography, 10(1):96–104, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Frankenberg:1967:PTS

[FCJ67] Dirk Frankenberg, Stephen Lee Coles, and R. E. Johannes. The potential trophic significance of Callianassa major fecal pellets. Limnology and Oceanography, 12(1):113–120, April 1967. CODEN LI-OCAH. ISSN 0024-3590.

Fremling:1963:MDD

[FE63] Calvin R. Fremling and John J. Evans. A method for determining the dissolved-oxygen concentration near the mud–water interface. Limnology and Oceanography, 8(3):363–364, April 1963. CODEN LI-OCAH. ISSN 0024-3590.

Fee:1969:NME

[Fee69] Everett J. Fee. A numerical model for the estimation of photosynthetic production, integrated over time and depth, in natural waters. Limnology and Oceanography, 14(6):906–911, November 1969. CO-DEN LIOCAH. ISSN 0024-3590.

Felsher:1964:CAD

[Fel64] Murray Felsher. A core aligner designed to recover oriented recent marine cores. Limnology and Oceanography, 9(4):603–605, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Fitzgerald:1967:EWS

[FF67] G. P. Fitzgerald and S. L. Faust. Effect of water sample preservation methods on the release of phosphorus from algae. Limnology and Oceanography, 12(2):332–334, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Fager:1966:EUE

[FFF^{+}66] E. W. Fager, A. O. Flechsig, R. F. Ford, R. I. Clutter, and R. J. Ghelardi. Equipment for use in ecological studies using scuba. Limnology and Oceanography, 11(4):503–509, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Fisher:1969:DDH

[FH69] Kenneth D. Fisher and Gregory Hugo. Distribution of Dolichodorus heterocephalus (Nematoda) in the littoral zone of Lake Champlain. Limnology and Oceanography, 14(4):617–620, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Fitzgerald:1969:FLE

[Fit69] George P. Fitzgerald. Field and laboratory evaluations of bioassays for nitrogen and phosphorus with algae and aquatic weeds. Limnology and Oceanography, 14(2):206–212, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Faber:1966:NCG

[FJ66] Daniel J. Faber and E. G. Jermolajev. A new copepod genus in the plankton of the Great Lakes. Limnology and Oceanography, 11(2): 301–303, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Fowler:1966:MC

[FK66] Gerald A. Fowler and L. D. Kulm. A multiple corer. Limnology and Oceanography, 11(4):630–633, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Finucane:1961:MVD

[FM61] John H. Finucane and Billie Z. May. Modified van Dorn water sampler. Limnology and Oceanography, 6(1):85–87, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Fukai:1962:AAV

[FM62] Rinnosuke Fukai and W. Wayne Meinke. Activation analyses of vanadium, arsenic, molybdenum, tungsten, rhenium, and gold in marine organisms. Limnology and Oceanography, 7(2):186–200, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Fuss:1965:SWO

[FO65] C. M. Fuss, Jr. and L. H. Ogren. A shallow water observation chamrer. Limnology and Oceanography, 10(2):290, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Fott:1968:BRG

[Fot68] B. Fott. Book review: Golubić, S. 1967. Algenvegetation der Felsen. Die Binnengewässer, v. 23. E. Schweizerbart’sche Verlagsbuchhandlung, Stuttgart, W. Germany. 183 p. DM5. Limnology and Oceanography, 13(4):726–727, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Fournier:1968:OPO

[Fou68] Robert O. Fournier. Observations of particulate organic carbon in the Mediterranean Sea and their relevance to the deep-living coccolithophorid Cyclococcolithus fragilis. Limnology and Oceanography, 13 (4):693–697, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Fox:1966:BRW

[Fox66] Marian C. Fox. Book review: Wilimovsky, J. J., and J. N. Wolfe [Eds.]. 1966. Environment of the Cape Thompson Region, Alaska. U. S. Atomic Energy Commission, Division of Technical Information, Oak Ridge, Tennessee. xvi + 1250 p. + 7 separate maps. $9.50. Limnology and Oceanography, 11(4):650, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Frolander:1962:BS

[FP62] H. F. Frolander and Ivan Pratt. A bottom skimmer. Limnology and Oceanography, 7(1):104–106, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Frassetto:1966:FEF

[Fra66] R. Frassetto. Further evidence of “fishbite” damage on deep-sea mooring lines in the Mediterranean. Limnology and Oceanography, 11(3):435–437, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Frey:1960:RMI

[Fre60a] David C. Frey. Die Rolle der Mikroorganismen im Stoffkreislauf der Seen. (German) [The role of microorganisms in the matter cycle of lakes]. Limnology and Oceanography, 5(3):346–347, April 1960. CO-DEN LIOCAH. ISSN 0024-3590.

Frey:1960:FWB

[Fre60b] David G. Frey. Fresh-water biology. Limnology and Oceanography, 5 (4):435–436, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Fremling:1961:SPS

[Fre61a] Calvin R. Fremling. Screened pail for sifting bottom-fauna samples. Limnology and Oceanography, 6(1):96, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Frey:1961:BPW

[Fre61b] David G. Frey. The biology of polluted waters. Limnology and Oceanography, 6(1):98, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Frey:1965:GCC

[Fre65] David G. Frey. Gynandromorphism in the chydorid cladocera. Limnology and Oceanography, 10(S1):R103–R114, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Frey:1966:BRL

[Fre66] David G. Frey. Book review: Liepolt, R. [Ed.]. 1965. Limnologie der Donau. E. Schweizerbart’sche Verlagsbuchhandlung (Nägele und Obermiller), Stuttgart. No. 1, 57 p. Limnology and Oceanography, 11(4):648–649, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Frey:1968:BRH

[Fre68] David G. Frey. Book review: Hrbáček, J. [Ed.]. 1966. Hydrobiological Studies I. Academia, Prague. 408 p. 70 Kčs. (Also available through E. Schweizerbart’sche Verlagsbuchhandlung, Stuttgart. DM 50). Limnology and Oceanography, 13(2):395–397, April 1968. CO-DEN LIOCAH. ISSN 0024-3590.

Frey:1969:BRH

[Fre69] David G. Frey. Book review: Hutchinson, G. E. 1967. A treatise on limnology. Volume II. Introduction to lake biology and the limnoplankton. John Wiley & Sons, New York, London, and Sydney. xi + 1115 p. $39.00. Limnology and Oceanography, 14(3):472–474, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Frankenberg:1967:CMA

[FS67] Dirk Frankenberg and K. L. Smith, Jr. Coprophagy in marine animals. Limnology and Oceanography, 12(3):443–450, April 1967. CO-DEN LIOCAH. ISSN 0024-3590.

Fanning:1969:IMS

[FS69] Kent A. Fanning and David R. Schink. Interaction of marine sediments with dissolved silica. Limnology and Oceanography, 14(1): 59–68, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Fiadeiro:1967:HS

[FSS67] M. Fiadeiro, L. Solórzano, and J. D. H. Strickland. Hydroxylamine in seawater. Limnology and Oceanography, 12(4):555–556, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Fuglister:1965:NMT

[FV65] F. C. Fuglister and A. D. Voorhis. A new method of tracking the Gulf Stream. Limnology and Oceanography, 10(S1):R115–R124, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Forster:1967:CTA

[FZ67] William O. Forster and Harry Zeitlin. Cobalt-60 tracer analysis of the nitroso-r method for the determination of cobalt in seawater. Limnology and Oceanography, 12(2):359–361, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Garrett:1965:CSF

[Gar65] William D. Garrett. Collection of slick-forming materials from the sea surface. Limnology and Oceanography, 10(4):602–605, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Gast:1966:DBS

[Gas66] James A. Gast. A drift bottle study in northern California. Limnology and Oceanography, 11(3):415–417, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Grzenda:1960:QMC

[GB60] Alfred R. Grzenda and Morris L. Brehmer. A quantitative method for the collection and measurement of stream periphyton. Limnology and Oceanography, 5(3):190–194, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Gillespie:1966:QSM

[GB66] D. M. Gillespie and C. J. D. Brown. A quantitative sampler for macroinvertebrates associated with aquatic macrophytes. Limnology and Oceanography, 11(3):404–406, April 1966. CODEN LIOCAH. ISSN 0024-3590.

G:1963:HCR

[GC63a] G and W. Coulter. Hydrological changes in relation to biological production in southern Lake Tanganyika. Limnology and Oceanography, 8(4):463–477, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Guilllard:1963:MCR

[GC63b] Robert R. L. Guilllard and Vivienne Cassie. Minimum cyanocobalamin requirements of some marine centric diatoms. Limnology and Oceanography, 8(2):161–165, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Ghassemi:1968:PYO

[GC68] Masood Ghassemi and R. F. Christman. Properties of the yellow organic acids of natural waters. Limnology and Oceanography, 13(4): 583–597, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Gray:1963:CSW

[GCV63] I. E. Gray and M. J. Cerame-Vivas. The circulation of surface waters in Raleigh Bay, North Carolina. Limnology and Oceanography, 8(3): 330–337, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Goering:1966:ERD

[GD66] John J. Goering and Vera A. Dugdale. Estimates of the rates of denitrification in a subarctic lake. Limnology and Oceanography, 11 (1):113–117, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Goering:1964:CDV

[GDM64] J. J. Goering, R. C. Dugdale, and D. W. Menzel. Cyclic diurnal variations in the uptake of ammonia and nitrate by photosynthetic organisms in the Sargasso Sea. Limnology and Oceanography, 9(3): 448–451, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Goering:1966:ELS

[GDM66] John J. Goering, Richard C. Dugdale, and David W. Menzel. Estimates of In Situ rates of nitrogen uptake by Trichodesmium sp. in the Tropical Atlantic Ocean. Limnology and Oceanography, 11(4): 614–620, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Gunnerson:1962:SSP

[GE62] C. G. Gunnerson and K. O. Emery. Suspended sediment and plankton over San Pedro Basin, California. Limnology and Oceanography, 7(1):14–20, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Geiger:1964:ODS

[Gei64] Stephen R. Geiger. Ophiophteus dubius from Scandinavian and North American waters. Limnology and Oceanography, 9(2):258–259, April 1964. CODEN LIOCAH. ISSN 0024-3590.

George:1964:OMA

[Geo64] J. D. George. Organic matter available to the polychaete Cirriformia tentaculata (Montagu) living in an intertidal mud flat. Limnology and Oceanography, 9(3):453–455, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Gucluer:1964:RMS

[GG64] Sevket M. Gucluer and M. Grant Gross. Recent marine sediments in Saanich Inlet, a stagnant marine basin. Limnology and Oceanography, 9(3):359–376, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Gieskes:1969:ETP

[Gie69] Joris M. Gieskes. Effect of temperature on the pH of seawater. Limnology and Oceanography, 14(5):679–685, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Gilmartin:1967:CIP

[Gil67] Malvern Gilmartin. Changes in inorganic phosphate concentration occurring during seawater sample storage. Limnology and Oceanography, 12(2):325–328, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Gillespie:1969:PSF

[Gil69] D. M. Gillespie. Population studies of four species of molluscs in the Madison River, Yellowstone National Park. Limnology and Oceanography, 14(1):101–114, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Gjessing:1964:FIW

[Gje64] Egil T. Gjessing. Ferrous iron in water. Limnology and Oceanography, 9(2):272–274, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Gerloff:1966:TAM

[GK66] G. C. Gerloff and P. H. Krombholz. Tissue analysis as a measure of nutrient availability for the growth of angiosperm aquatic plants. Limnology and Oceanography, 11(4):529–537, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Grzenda:1964:WPI

[GLN64] Alfred R. Grzenda, Gerald J. Lauer, and H. Page Nicholson. Water pollution by insecticides in an agricultural river basin. II. The zooplankton, bottom fauna, and fish. Limnology and Oceanography, 9 (3):318–323, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1967:TAD

[GMH67] Charles R. Goldman, David T. Mason, and John E. Hobbie. Two Antarctic desert lakes. Limnology and Oceanography, 12(2):295–310, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1963:LII

[GMW63] Charles R. Goldman, David T. Mason, and Brian J. B. Wood. Light injury and inhibition in Antarctic freshwater phytoplankton. Limnology and Oceanography, 8(3):313–322, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Goming:1964:NFT

[GN64] John J. Goming and John C. Neess. Nitrogen fixation in two Wisconsin lakes. Limnology and Oceanography, 9(4):530–539, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Gifford:1961:CCI

[GO61] Cameron E. Gifford and Eugene P. Odum. Chlorophyll-a content of intertidal zones on a rocky seashore. Limnology and Oceanography, 6(1):83–85, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Goldberg:1962:ECS

[Gol62a] Edward D. Goldberg. Elemental composition of some pelagic fishes. Limnology and Oceanography, 7(S1):lxxii–lxxv, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Goldberg:1962:ESV

[Gol62b] Edward D. Goldberg. Errata — supplement to volume VII of Limnology and Oceanography. Limnology and Oceanography, 7(S1):xciii, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1962:MSN

[Gol62c] Charles R. Goldman. A method of studying nutrient limiting factors In Situ in water columns isolated by polyethylene film. Limnology and Oceanography, 7(1):99–101, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Gold:1964:MAV

[Gol64] Kenneth Gold. A microbiological assay for vitamin B_{12} in seawater using radiocarbon. Limnology and Oceanography, 9(3):343–347, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Goldberg:1965:OMS

[Gol65] Edward D. Goldberg. An observation on marine sedimentation rates during the Pleistocene. Limnology and Oceanography, 10(S1):R125– R128, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Gold:1968:SFA

[Gol68] Kenneth Gold. Some factors affecting the stability of thiamine. Limnology and Oceanography, 13(1):185–188, April 1968. CODEN LI-OCAH. ISSN 0024-3590.

Gorham:1960:CDS

[Gor60] Eville Gorham. Chlorophyll derivatives in surface muds from the English lakes. Limnology and Oceanography, 5(1):29–33, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1964:MCA

[Gor64] Eville Gorham. Morphometric control of annual heat budgets in temperate lakes. Limnology and Oceanography, 9(4):525–529, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1966:EDF

[Gor66] Donald C. Gordon, Jr. The effects of the deposit feeding polychaete Pectinaria gouldii on the intertidal sediments of Barnstable Harbor. Limnology and Oceanography, 11(3):327–332, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1968:JCF

[Gor68] Eville Gorham. Journal coverage in the field of limnology. Limnology and Oceanography, 13(2):366–369, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Griffiths:1969:OSL

[GPE69] Mary Griffiths, Pamela S. Perrott, and W. T. Edmondson. Oscillaxanthin in the sediment of Lake Washington. Limnology and Oceanography, 14(3):317–326, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Graham:1966:SDO

[Gra66] Joseph J. Graham. Secchi disc observations and extinction coefficients in the central and eastern north Pacific Ocean. Limnology and Oceanography, 11(2):184–190, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Graefe:1967:DRL

[Gra67] Volker Graefe. Digital recording of In Situ depth–temperature– salinity data. Limnology and Oceanography, 12(4):718–721, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Gold:1966:TDD

[GRB66] Kenneth Gold, Oswald A. Roels, and Harvey Bank. Temperature dependent destruction of thiamine in seawater. Limnology and Oceanography, 11(3):410–413, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Gross:1965:GSB

[Gro65a] M. Ghant Gross. Geochemistry of sediments. Brief survey. Limnology and Oceanography, 10(4):612–613, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Gross:1965:MGG

[Gro65b] M. Grant Gross. Marine geology of the Gulf of California. Limnology and Oceanography, 10(2):303–304, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Gross:1969:BRB

[Gro69] M. Grant Gross. Book review: Bouma, A. H. 1969. Methods for the study of sedimentary structures. John Wiley and Sons, New York. xvi + 458 p. $19.95. Limnology and Oceanography, 14(6):966, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1965:IOR

[GS65] Eville Gorham and Dalway J. Swaine. The influence of oxidizing and reducing conditions upon the distribution of some elements in lake sediments. Limnology and Oceanography, 10(2):268–279, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Gessner:1967:MDA

[GS67] Fritz Gessner and Reimer Simonsen. Marine diatoms in the Amazon? Limnology and Oceanography, 12(4):709–711, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Gunter:1961:SRE

[Gun61] Gordon Gunter. Some relations of estuarine organisms to salinity. Limnology and Oceanography, 6(2):182–190, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Gushing:1961:PSS

[Gus61] Colbert E. Gushing, Jr. A plankton sub-sampler. Limnology and Oceanography, 6(4):489–490, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Gutknecht:1961:MRZ

[Gut61] John Gutknecht. Mechanism of radioactive zinc uptake by Ulva lactuca. Limnology and Oceanography, 6(4):426–431, April 1961. CO-DEN LIOCAH. ISSN 0024-3590.

Gutknecht:1963:UBM

[Gut63] John Gutknecht. ^{65}Zn uptake by benthic marine algae. Limnology and Oceanography, 8(1):31–38, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Gutknecht:1965:URC

[Gut65] John Gutknecht. Uptake and retention of cesium 137 and zinc 65 by seaweeds. Limnology and Oceanography, 10(1):58–66, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Graham:1966:NDD

[GV66] Joseph J. Graham and George B. Vaughan. A new depressor design. Limnology and Oceanography, 11(1):130–135, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Gates:1960:TGM

[GW60] Jean A. Gates and William B. Wilson. The toxicity of Gonyaulax monilata Howell to Mugil cephalus. Limnology and Oceanography, 5 (3):171–174, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Hahow:1963:ATM

[Hah63] K. Hahow. Acclimation to temperature in the marine copepod, Calanus finmarchicus (Gunner). Limnology and Oceanography, 8(1): 1–8, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:1964:PPI

[Ham64] R. D. Hamilton. Photochemical processes in the inorganic nitrogen cycle of the sea. Limnology and Oceanography, 9(1):107–111, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Hair:1968:IAS

[HAM68] Malcolm E. Hair, James E. Alexander, and John J. A. Mclaughlin. An inexpensive apparatus for sampling shallow water. Limnology and Oceanography, 13(4):718–719, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:1969:EAP

[Ham69a] Andrew L. Hamilton. On estimating annual production. Limnology and Oceanography, 14(5):771–781, September 1969. CODEN LI-OCAH. ISSN 0024-3590. See reply [Hyn69].

Hamilton:1969:NLS

[Ham69b] D. H. Hamilton, Jr. Nutrient limitation of summer phytoplankton growth in Cayuga Lake. Limnology and Oceanography, 14(4):579– 590, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Haq:1967:NPM

[Haq67] S. M. Haq. Nutritional, physiology of Metridia lucens and M. longa from the Gulf of Maine. Limnology and Oceanography, 12(1):40–51, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Harvey:1966:LVP

[Har66a] George W. Harvey. A low velocity plankton siphon. Limnology and Oceanography, 11(4):646–647, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Harvey:1966:MCS

[Har66b] George W. Harvey. Microlayer collection from the sea surface: a new method and initial results. Limnology and Oceanography, 11(4): 608–613, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Harriss:1967:SCI

[Har67] Robert C. Harriss. Silica and chloride in interstitial waters of river and lake sediments. Limnology and Oceanography, 12(1):8–12, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Harder:1968:RPO

[Har68] Wilhelm Harder. Reactions of plankton organisms to water stratification. Limnology and Oceanography, 13(1):156–168, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1969:SOU

[Har69] Barry T. Hargrave. Similarity of oxygen uptake by benthic communities. Limnology and Oceanography, 14(5):801–805, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Hasler:1968:BRC

[Has68a] Arthur D. Hasler. Book review: Campbell, T. H., and R. O. Sylvester [Eds.]. 1968. Water resources and public policy. University of Washington Press, Seattle. 253 p. $9.50. Limnology and Oceanography, 13 (4):728–729, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Hastenrath:1968:ELS

[Has68b] Stefan L. Hastenrath. Estimates of the latent and sensible heat flux for the Caribbean Sea and the Gulf of Mexico. Limnology and Oceanography, 13(2):322–331, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Hayes:1964:LNA

[Hay64] F. R. Hayes. Limnology in North America. Limnology and Oceanography, 9(1):156–158, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Haynes:1968:BRS

[Hay68] H. B. N. Haynes. Book review: Studies on Loch Lomond 2. The Fauna of the River Endrick. Univ. Glasgow, Blackie and Son, Glasgow. 194 p. 27s. Limnology and Oceanography, 13(1):207–208, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Hazelwood:1962:TPE

[Haz62] Donald H. Hazelwood. Temperature and photoperiod effects on cyclomorpihosis in Daphnia. Limnology and Oceanography, 7(2):230–232, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Hazelwood:1966:ITE

[Haz66] Donald H. Hazelwood. Illumination and turbulence effects on relative growth in Daphnia. Limnology and Oceanography, 11(2):212–216, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Haines:1961:ESW

[HB61] Donald A. Haines and Reid A. Bryson. An empirical study of wind factor in Lake Mendota. Limnology and Oceanography, 6(3):356–364, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Hansen:1962:RCM

[HB62] Palle G. Hansen and Eric G. Barham. Resonant cavity measurements of the effects of “red water” plankton on the attenuation of underwater sound. Limnology and Oceanography, 7(1):8–13, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Hartman:1966:MCI

[HB66] Richard T. Hartman and Doris L. Brown. Methane as a constituent of the internal atmosphere of vascular hydrophytes. Limnology and Oceanography, 11(1):109–112, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Harp:1967:DTP

[HC67] George L. Harp and Robert S. Campbell. The distribution of Tendipes plumosus (Linné) in mineral acid water. Limnology and Oceanography, 12(2):260–263, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Hicks:1968:GDN

[HC68a] Sonja E. Hicks and Francis G. Carey. Glucose determination in natural waters. Limnology and Oceanography, 13(2):361–363, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1968:SMA

[HC68b] H. B. N. Hynes and Mary J. Coleman. A simple method of assessing the annual production of stream benthos. Limnology and Oceanography, 13(4):569–573, October 1968. CODEN LIOCAH. ISSN 0024-3590. See critique [Fag69].

Hobbie:1969:RCB

[HC69] John E. Hobbie and Claude C. Crawford. Respiration corrections for bacterial uptake of dissolved organic compounds in natural waters. Limnology and Oceanography, 14(4):528–532, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Howmiller:1969:CAS

[HD69] Richard Howmiller and Keith Dahnke. Chemical analysis of salt from Tagus Crater Lake, Isabela, Galapagos. Limnology and Oceanography, 14(4):602–604, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Healey:1967:SDC

[Hea67] M. C. Healey. The seasonal and diel changes in distribution of Diaptomus leptopus in a small eutrophic lake. Limnology and Oceanography, 12(1):34–39, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Hedgpeth:1960:SSC

[Hed60] Joel W. Hedgpeth. The sea off Southern California, a modem habitat of petroleum. Limnology and Oceanography, 5(3):347–348, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Hunkins:1960:BGO

[HEHM60] Kenneth L. Hunkins, Maurice Ewing, Bruce C. Heezen, and Robert J. Menzies. Biological and geological observations on the first photographs of the Arctic Ocean deep-sea floor. Limnology and Oceanography, 5(3):154–161, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Hellebust:1965:ESO

[Hel65] J. A. Hellebust. Excretion of some organic compounds by marine phytoplankton. Limnology and Oceanography, 10(2):192–206, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Hepher:1962:PPF

[Hep62] B. Hepher. Primary production in fishponds and its application to fertilization experiments. Limnology and Oceanography, 7(2):131– 136, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Heron:1961:SVP

[Her61] J. Heron. The seasonal variation of phosphate, silicate, and nitrate in waters of the English Lake District. Limnology and Oceanography, 6(3):338–346, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Heron:1962:DPW

[Her62] J. Heron. Determination of phosphate in water after storage in polyethylene. Limnology and Oceanography, 7(3):316–321, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Herman:1963:PFE

[Her63a] Sidney S. Herman. Planktonic fish eggs and larvae of Narragansett Bay. Limnology and Oceanography, 8(1):103–109, April 1963. CO-DEN LIOCAH. ISSN 0024-3590.

Herman:1963:VMO

[Her63b] Sidney S. Herman. Vertical migration of the opossum shrimp, Neomysis americana Smith. Limnology and Oceanography, 8(2):228–238, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Heubach:1969:NAS

[Heu69] William Heubach. Neomysis awatschensis in the Sacramento–San Joaquin River Estuary. Limnology and Oceanography, 14(4):533–546, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1968:PEZ

[HG68] Barry T. Hargrave and Glen H. Geen. Phosphorus excretion by zooplankton. Limnology and Oceanography, 13(2):332–342, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Hales:1969:CTT

[HG69] Donald C. Hales and Arden R. Gaufin. Comparison of two types of stream insect drift nets. Limnology and Oceanography, 14(3):459– 461, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Holm-Hansen:1968:DPO

[HH68] Osmund Holm-Hansen. Determination of particulate organic nitrogen. Limnology and Oceanography, 13(1):175–178, April 1968. CO-DEN LIOCAH. ISSN 0024-3590.

Holm-Hansen:1969:DMB

[HH69] Osmund Holm-Hansen. Determination of microbial biomass in ocean profiles. Limnology and Oceanography, 14(5):740–747, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Holm-Hansen:1966:MAT

[HHB66] Osmund Holm-Hansen and Charles R. Booth. The measurement of adenosine triphosphate in the ocean and its ecological significance. Limnology and Oceanography, 11(4):510–519, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:1967:ATC

[HHH67] Robert D. Hamilton and Osmund Holm-Hansen. Adenosine triphosphate content of marine bacteria. Limnology and Oceanography, 12 (2):319–324, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Holm-Hansen:1966:DAB

[HHSW66] O. Holm-Hansen, J. D. H. Strickland, and P. M. Williams. A detailed analysis of biologically important substances in a profile off southern California. Limnology and Oceanography, 11(4):548–561, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Hickman:1969:MDP

[Hic69] M. Hickman. Methods for determining the primary productivity of epipelic and epipsammic algal associations. Limnology and Oceanography, 14(6):936–941, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Hilsenhoff:1969:ASD

[Hil69a] William L. Hilsenhoff. An artificial substrate device for sampling benthic stream invertebrates. Limnology and Oceanography, 14(3): 465–471, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Hiltunen:1969:DOW

[Hil69b] Jarl K. Hiltunen. Distribution of oligochaetes in western Lake Erie, 1961. Limnology and Oceanography, 14(2):260–264, March 1969. CO-DEN LIOCAH. ISSN 0024-3590.

Hirata:1963:ASA

[Hir63] Hachiro Hirata. An automatic solution adder. Limnology and Oceanography, 8(2):299–300, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Hassenteufel:1963:IGS

[HJK63] W. Hassenteufel, R. Jagitsch, and F. F. Koczy. Impregnation of glass surface against sorption of phospiiate traces. Limnology and Oceanography, 8(2):152–156, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Hassan:1964:USM

[HL64] El Sayed Mohamed Hassan and S. Loeb. The use of semipermeable membranes to conserve fertilizer in artificially fertilized bays. Limnology and Oceanography, 9(2):259–260, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Hlll:1967:NTW

[Hll67] Hibbert Hlll. A note on temperatures and water conditions beneath lake ice in spring. Limnology and Oceanography, 12(4):550–552, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Hughes:1962:OBA

[HM62] John T. Hughes and George C. Matthiessen. Observations on the biology of the American lobster, Homarus americanus. Limnology and Oceanography, 7(3):414–421, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Hill:1964:DAU

[HM64] Eddie P. Hill and Richard Y. Morita. Dehydrogenase activity under hydrostatic pressure by isolated mitochondria obtained from Allomyces macrogynus. Limnology and Oceanography, 9(2):243–248, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Haight:1966:SPD

[HM66] Janet J. Haight and Richard Y. Morita. Some physiological differences in Vibrio marinus grown at environmental and optimal temperatures. Limnology and Oceanography, 11(4):470–474, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Hobson:1969:DCC

[HM69] Louis A. Hobson and David W. Menzel. The distribution and chemical composition of organic particulate matter in the sea and sediments off the east coast of South America. Limnology and Oceanography, 14(1):159–163, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Haven:1965:AHI

[HMA65] Dexter S. Haven and R. Morales-Alamo. Apparatus for holding the individual oysters under equal water flows. Limnology and Oceanography, 10(4):605–606, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Haven:1966:ABO

[HMA66] Dexter S. Haven and Reinaldo Morales-Alamo. Aspects of biodeposition by oysters and other invertebrate filter feeders. Limnology and Oceanography, 11(4):487–498, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Hoffman:1961:ERT

[HO61a] Dale A. Hoffman and John R. Olive. The effects of rotenone and toxaphene upon plankton of two Colorado reservoirs. Limnology and Oceanography, 6(2):219–222, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Hopwood:1961:SMA

[HO61b] Alfred J. Hopwood and John R. Olive. Simple mechanical aid to limnological sampling. Limnology and Oceanography, 6(1):87–88, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1961:STL

[Hob61] John E. Hobbie. Summer temperatures in Lake Schrader, Alaska. Limnology and Oceanography, 6(3):326–329, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Hobson:1966:SIC

[Hob66] Louis A. Hobson. Some influences of the Columbia River effluent on marine phytoplankton during January 1961. Limnology and Oceanography, 11(2):223–234, April 1966. CODEN LIOCAH. ISSN 0024-

Hobson:1967:SVD

[Hob67] Louis A. Hobson. The seasonal and vertical distribution of suspended particulate matter in an area of the northeast Pacific Ocean. Limnology and Oceanography, 12(4):642–649, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Hoese:1960:BCB

[Hoe60] H. Dickson Hoese. Biotic changes in a bay associated with the end of a drought. Limnology and Oceanography, 5(3):326–336, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Holmes:1962:MPA

[Hol62] R. W. Holmes. Marine phytoplankton — areal surveys. Limnology and Oceanography, 7(S1):xxvii–xxviii, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Holland:1968:CMS

[Hol68] Ruth E. Holland. Correlation of Melosira species with trophic conditions in Lake Michigan. Limnology and Oceanography, 13(3):555–557, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Holland:1969:SFL

[Hol69] Ruth E. Holland. Seasonal fluctuations of Lake Michigan diatoms. Limnology and Oceanography, 14(3):423–436, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Hopper:1960:RMZ

[Hop60] Arthur F. Hopper. The resistance of marine zooplankton of the Caribbean and South Atlantic to changes in salinity. Limnology and Oceanography, 5(1):43–47, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Hopkins:1961:O

[Hop61a] Sewell H. Hopkins. Oysters. Limnology and Oceanography, 6(1): 100–101, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Hopkins:1961:NCM

[Hop61b] Thomas L. Hopkins. Natural coloring matter as an indicator of inshore water masses. Limnology and Oceanography, 6(4):484–486, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Hopkins:1962:ZS

[Hop62] Thomas L. Hopkins. A zooplankton subsampler. Limnology and Oceanography, 7(3):424–426, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Hopkins:1969:ZSC

[Hop69] Thomas L. Hopkins. Zooplankton standing crop in the Arctic basin. Limnology and Oceanography, 14(1):80–85, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Horrer:1962:POT

[Hor62] Paul L. Horrer. Physical oceanography of the test area. Limnology and Oceanography, 7(S1):vii–xxvi, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Hulburt:1963:DPS

[HR63] Edward M. Hulburt and Janet Rodman. Distribution of phytoplankton species with respect to salinity between the coast of southern New England and Bermuda. Limnology and Oceanography, 8(2):263–269, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Hansen:1966:NDE

[HR66] Donald V. Hansen and Maurice Rattray, Jr. New dimensions in estuary classification. Limnology and Oceanography, 11(3):319–326, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Hammer:1968:SSC

[HS68] U. Theodore Hammer and William W. Sawchyn. Seasonal succession and congeneric associations of Diaptomus spp. (Copepoda) in some Saskatchewan ponds. Limnology and Oceanography, 13(3):476–484, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Howmiller:1969:HWS

[HS69] Richard P. Howmiller and William E. Sloey. A horizontal water sampler for investigation of stratified waters. Limnology and Oceanography, 14(2):291–292, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Hsiao:1965:KSA

[Hsi65] Sidney C. Hsiao. Kinetic studies on alkaline phosphatase from echinoplutei. Limnology and Oceanography, 10(S1):R129–R136, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Hellebust:1967:EEC

[HT67] Johan A. Hellebust and John Terborgh. Effects of environmental conditions on the rate of photosynthesis and some photosynthetic enzymes in Dunaliella tertiolecta Butcher. Limnology and Oceanography, 12(4):559–567, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Howe:1969:SOD

[HT69] M. R. Howe and R. I. Tait. Some observations of the diurnal heat wave in the ocean. Limnology and Oceanography, 14(1):16–22, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Hughes:1969:SCM

[Hug69] Peter Hughes. Submarine cable measurements of tidal currents in the Irish Sea. Limnology and Oceanography, 14(2):269–278, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Hulburt:1962:PSS

[Hul62] Edward M. Hulburt. Phytoplankton in the southwestern Sargasso Sea and North Equatorial Current, February 1961. Limnology and Oceanography, 7(3):307–315, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Hunkins:1965:TSS

[Hun65a] Kenneth L. Hunkins. Tide and storm surge observations in the Chukchi Sea. Limnology and Oceanography, 10(1):29–39, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Huntsman:1965:EMS

[Hun65b] A. G. Huntsman. The ectology of Margaree salmon. Limnology and Oceanography, 10(S1):R137–R147, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1965:BBU

[HW65] John E. Hobbie and Richard T. Wrights. Bioassay with bacterial uptake kinetics: glucosein freshwater. Limnology and Oceanography, 10(3):471–474, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Hall:1967:ZFL

[HW67] Donald J. Hall and George G. Waterman. Zooplankton of the finger lakes. Limnology and Oceanography, 12(4):542–544, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Holmes:1967:RWJ

[HWE67] R. W. Holmes, P. M. Williams, and R. W. Eppley. Red water in La Jolla Bay, 1964–1966. Limnology and Oceanography, 12(4):503–512, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1968:BRM

[Hyn68] H. B. N. Hynes. Book review: Maitland, P. S. 1966. Studies on Loch Lomond 2. The Fauna of the River Endrick. Univ. Glasgow, Blackie and Son, Glasgow. 194 p. 27s. Limnology and Oceanography, 13(1): 208, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1969:RH

[Hyn69] H. B. N. Hynes. Reply [to Hamilton]. Limnology and Oceanography, 14(5):782, September 1969. CODEN LIOCAH. ISSN 0024-3590. See [Ham69a].

Higaki:1962:NHW

[HZ62] Shiro Higaki and Harry Zeitlin. Nitrate in Hawaiian waters. Limnology and Oceanography, 7(1):103, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Iovino:1969:RLC

[IB69] A. J. Iovino and W. H. Bradley. The role of larval Chironomidae in the production of lacustrine copropel in Mud Lake, Marion County, Florida. Limnology and Oceanography, 14(6):898–905, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Ichiye:1961:HSA

[IJ61] Takashi Ichiye and Meredith L. Jones. On the hydrography of the St. Andrew Bay System, Florida. Limnology and Oceanography, 6 (3):302–311, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Ivanoff:1961:CSI

[IJW61] Alexandre Ivanoff, Nils Jerlov, and Talbot H. Waterman. A comparative study of irradiance, beam transmittance and scattering in the sea near Bermuda. Limnology and Oceanography, 6(2):129–148, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Ichiye:1966:PMD

[IP66] Takashi Ichiye and Noel B. Plutchak. Photodensitometric measurement of dye concentration in the ocean. Limnology and Oceanography, 11(3):364–370, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Iselin:1969:BRW

[Ise69] C. O’D. Iselin. Book review: Williams, J., J. J. Higginson, and J. D. Rohrbough. 1968. Sea and air. The naval environment. U.S. Naval Institute, Annapolis, Maryland. 338 p. $11.50. Limnology and Oceanography, 14(4):650, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Jeffrey:1967:PCM

[JA67] S. W. Jeffrey and M. B. Allen. A paper chromatographic method for the separation of phytoplankton pigments at sea. Limnology and Oceanography, 12(4):533–537, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Jacobs:1961:LCM

[Jac61] Jürgen Jacobs. Laboratory cultivation of the marine copepod Pseudodiaptomus coronatus Williams. Limnology and Oceanography, 6 (4):443–446, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Jacobs:1965:SEM

[Jac65] W. C. Jacobs. Symposium — environmental measurements — valid data and logical interpretation. Limnology and Oceanography, 10(4): 613–614, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Jannasch:1967:GMB

[Jan67] Holger W. Jannasch. Growth of marine bacteria at limiting concentrations of organic carbon in seawater. Limnology and Oceanography, 12(2):264–271, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Jarvis:1967:ASA

[Jar67] N. L. Jarvis. Adsorption of surface-active material at the sea–air interface. Limnology and Oceanography, 12(2):213–221, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Jawed:1969:BNN

[Jaw69] Mohammad Jawed. Body nitrogen and nitrogenous excretion in Neomysis rayii Murdoch and Euphausia pacifica Hansen. Limnology and Oceanography, 14(5):748–754, September 1969. CODEN LI-OCAH. ISSN 0024-3590.

Johnson:1964:SSC

[JCG64] Robert F. Johnson, J. J. Carroll, and L. J. Greenfield. Some sources of carbonate in molluscan shell formation. Limnology and Oceanography, 9(3):377–384, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Jeffrles:1960:WOA

[Jef60] Harry P. Jeffrles. Winter occurrences of Anguilla rostrata elvers in New England and Middle Atlantic estuaries. Limnology and Oceanography, 5(3):338–340, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Jeffries:1962:ECR

[Jef62a] Harry P. Jeffries. Environmental characteristics of Raritan Bay, a polluted estuary. Limnology and Oceanography, 7(1):21–31, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Jeffries:1962:STA

[Jef62b] Harry P. Jeffries. Succession of two acartia species in estuaries. Limnology and Oceanography, 7(3):354–362, April 1962. CODEN LI-OCAH. ISSN 0024-3590.

Jeffries:1964:CSE

[Jef64] Harry P. Jeffries. Comparative studies on estuarine zooplankton. Limnology and Oceanography, 9(3):348–358, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Jeffrey:1968:PPP

[Jef68] S. W. Jeffrey. Photosynthetic pigments of the phytoplankton of some coral reef waters. Limnology and Oceanography, 13(2):350–355, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Jeffries:1969:SCT

[Jef69] H. Perry Jeffries. Seasonal composition of temperate plankton communities: free amino acids. Limnology and Oceanography, 14(1): 41–52, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Jarvis:1967:SCC

[JGST67] N. L. Jarvis, W. D. Garrett, M. A. Scheiman, and C. O. Timmons. Surface chemical characterization of surface-active material in seawater. Limnology and Oceanography, 12(1):88–96, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Jitts:1964:TSL

[Jit64] H. R. Jitts. A twin six-liter plastic water sampler. Limnology and Oceanography, 9(3):452, April 1964. CODEN LIOCAH. ISSN 0024-

Jerde:1966:MES

[JL66] Charles W. Jerde and Reuben Lasker. Molting of euphausiid shrimps: shipboard observations. Limnology and Oceanography, 11(1):120– 124, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1968:MCS

[JM68] M. G. Johnson and D. H. Matheson. Macroinvertebrate communities of the sediments of Hamilton Bay and adjacent Lake Ontario. Limnology and Oceanography, 13(1):99–111, April 1968. CODEN LI-OCAH. ISSN 0024-3590.

Joeris:1964:HSC

[Joe64] Leonard S. Joeris. A horizontal sampler for collection of water samples near the botttom. Limnology and Oceanography, 9(4):595–598, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1963:ZCH

[Joh63a] Martin W. Johnson. Zooplankton collections from the high polar basin with special reference to the copepoda. Limnology and Oceanography, 8(1):89–102, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Johnston:1963:EGM

[Joh63b] R. Johnston. Effects of gibberellins on marine algae in mixed cultures. Limnology and Oceanography, 8(2):270–275, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Johannes:1964:URD

[Joh64a] R. E. Johannes. Uptake and release of dissolved organic phosphorus by representatives of a coastal marine ecosystem. Limnology and Oceanography, 9(2):224–234, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Johannes:1964:URP

[Joh64b] R. E. Johannes. Uptake and release of phosphorus by a benthic marine amphipod. Limnology and Oceanography, 9(2):235–242, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Johannes:1965:IMP

[Joh65a] R. E. Johannes. Influence of marine protozoa on nutrient regeneration. Limnology and Oceanography, 10(3):434–442, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1965:TVI

[Joh65b] Ralph Gordon Johnson. Temperature variation in the infaunal environment of a sand flat. Limnology and Oceanography, 10(1):114–120, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1966:MIB

[Joh66] Phyllis T. Johnson. Mass mortality in a bivalve mollusc. Limnology and Oceanography, 11(3):429–431, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Johannes:1967:EOA

[Joh67a] R. E. Johannes. Ecology of organic aggregates in the vicinity of a coral reef. Limnology and Oceanography, 12(2):189–195, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1967:SOH

[Joh67b] Martin W. Johnson. Some observations on the hatching of Tortanus discaudatus eggs subjected to low temperatures. Limnology and Oceanography, 12(3):405–410, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1967:SIW

[Joh67c] Ralph Gordon Johnson. Salinity of interstitial water in a sandy beach. Limnology and Oceanography, 12(1):1–7, April 1967. CODEN LI-OCAH. ISSN 0024-3590.

Jones:1960:CSS

[Jon60a] Galen E. Jones. Caulobacter sp. in sea water. Limnology and Oceanography, 5(4):432–433, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Jones:1960:ANR

[Jon60b] Raymond F. Jones. The accumulation of nitrosyl ruthenium by fine particles and marine organisms. Limnology and Oceanography, 5(3): 312–325, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Jones:1967:GEC

[Jon67a] Galen E. Jones. Growth of Escherichia coli in heat- and coppertreated synthetic seawater. Limnology and Oceanography, 12(3):167– 172, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Jones:1967:PAS

[Jon67b] Galen E. Jones. Precipitates from autoclaved seawater. Limnology and Oceanography, 12(3):165–167, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Jossi:1966:IOM

[Jos66] Jack W. Jossi. The ICITA one-meter plankton net: description and evaluation. Limnology and Oceanography, 11(4):640–642, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Joyner:1968:BRA

[Joy68] T. Joyner. Book review: Angino, E. E., and G. K. Billings. 1967. Methods in Geochemistry and Geophysics. 7. Atomic Absorption Spectrometry in Geology. Elsevier, Amsterdam, London, New York. Limnology and Oceanography, 13(1):205–206, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Jitts:1961:DZT

[JS61] H. R. Jitts and B. D. Scott. The determination of zero-thickness activity in Geiger counting of C^{14} solutions used in marine productivity studies. Limnology and Oceanography, 6(2):116–123, April 1961. CO-DEN LIOCAH. ISSN 0024-3590.

Johannes:1966:CNV

[JS66] R. E. Johannes and Masako Satomi. Composition and nutritive value of fecal pellets of a marine crustacean. Limnology and Oceanography, 11(2):191–197, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1968:CDM

[JSP68] Roy M. Johnson, Rose Mary Schwent, and Wesley Press. The characteristics and distribution of marine bacteria isolated from the Indian Ocean. Limnology and Oceanography, 13(4):656–664, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Kahl:1963:TSP

[Kah63] M. Philip Kahl, Jr. Technique for sampling population density of small shallow-water fish. Limnology and Oceanography, 8(2):302–304, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Kane:1967:OAS

[Kan67] Jasmine E. Kane. Organic aggregates in surface waters of the Ligurian Sea. Limnology and Oceanography, 12(2):287–294, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Katz:1965:FRP

[Kat65] Max Katz. Fish and river pollution. Limnology and Oceanography, 10(4):612–613, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Kutzbach:1963:FMS

[KB63] John E. Kutzbach and Reid A. Bryson. Field measurement of the speed of propagation of wind waves as a function of wavelength. Limnology and Oceanography, 8(2):157–160, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Kevern:1965:PPE

[KB65] N. R. Kevern and R. C. Ball. Primary productivity and energy relationships in artificial streams. Limnology and Oceanography, 10(1): 74–87, April 1965. CODEN LIOCAH. ISSN 0024-3590.

King:1967:CEP

[KB67] Darrell L. King and Robert C. Ball. Comparative energetics of a polluted stream. Limnology and Oceanography, 12(1):27–33, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Khailov:1969:RDO

[KB69] K. M. Khailov and Z. P. Burlakova. Release of dissolved organic matter by marine seaweeds and distribution of their total organic production to inshore communities. Limnology and Oceanography, 14(4):521–527, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Ketchum:1964:PWW

[KC64] Ketchum and Nathaniel Corwin. The persistence of “winter” water on the continental shelf south of Long Island, New York. Limnology and Oceanography, 9(4):467–475, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Ketchum:1965:CPP

[KC65] Bostwick H. Ketchum and Nathaniel Corwin. The cycle of phosphorus in a plankton bloom in the Gulf of Mainea. Limnology and Oceanography, 10(S1):R148–R161, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Kibby:1968:TCO

[KDB68] H. V. Kibby, J. R. Donaldson, and C. E. Bond. Temperature and current observations in Crater Lake, Oregon. Limnology and Oceanography, 13(2):363–366, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Kester:1967:PAS

[KDCP67] Dana R. Kester, Iver W. Duedall, Donald N. Connors, and Ricardo M. Pytkowicz. Preparation of artificial seawater. Limnology and Oceanography, 12(3):176–179, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Knowles:1962:PMR

[KEB62] G. Knowles, R. W. Edwards, and R. Briggs. Polarographic measurement of the rate of respiration of natural sediments. Limnology and Oceanography, 7(4):481–484, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Kesteven:1960:PMB

[Kes60] G. L. Kesteven. Perspectives in marine biology. Limnology and Oceanography, 5(3):237–242, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Klots:1961:EHP

[Klo61] Cornelius E. Klots. Effect of hydrostatic pressure upon the solubility of gases. Limnology and Oceanography, 6(3):365–366, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Klotz:1963:VSD

[Klo63] Irving M. Klotz. Variation of solubility with depth in the ocean: a thermodynamic analysis. Limnology and Oceanography, 8(2):149– 151, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Kemmerer:1969:MGP

[KN69] Andrew J. Kemmerer and John M. Neuhold. A method for gross primary productivity measurements. Limnology and Oceanography, 14(4):607–610, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1967:BRC

[Koh67] Alan J. Kohn. Book review: Cromie, W. J. 1966. The Living World of the Sea. Prentice Hall, Englewood Cliffs, New Jersey. 343 p. $6.95. Limnology and Oceanography, 12(3):183, April 1967. CODEN LI-OCAH. ISSN 0024-3590.

Kohn:1968:BRC

[Koh68] Alan J. Kohn. Book review: Cheng, T. C. 1967. Marine Molluscs as Hosts for Symbioses. Advances in marine biology, v. 5. Academic Press, Berkeley Square, London. 424 p. 100s. $17.50. Limnology and Oceanography, 13(3):566, July 1968. CODEN LIOCAH. ISSN 0024-

Kormondy:1965:ULZ

[Kor65] Edward J. Kormondy. Uptake and loss of zinc-65 in the dragonfly plathemis lydia. Limnology and Oceanography, 10(3):427–433, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Kormondy:1968:WLC

[Kor68] Edward J. Kormondy. Weight loss of cellulose and aquatic macrophytes in a Carolina bay. Limnology and Oceanography, 13(3):522– 526, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Kester:1967:DAD

[KP67] Dana R. Kester and Ricardo M. Pytkowicx. Determination of the apparent dissociation constants of phosphoric acid in seawater. Limnology and Oceanography, 12(2):243–252, April 1967. CODEN LI-OCAH. ISSN 0024-3590.

Kester:1968:MSA

[KP68] Dana R. Kester and Ricardo M. Pytkowicx. Magnesium sulfate association at 25c in synthetic seawater. Limnology and Oceanography, 13(4):670–674, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Kester:1969:SMC

[KP69] Dana R. Kester and Ricardo M. Pytkowicx. Sodium, magnesium, and calcium sulfate ion-pairs in seawater at 25c. Limnology and Oceanography, 14(5):686–692, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Kornicker:1962:FCP

[KS62] Louis S. Kornicker and Donald F. Squires. Floating corals: a possible source of erroneous distribution data. Limnology and Oceanography, 7(4):447–452, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Kuenzler:1961:PBM

[Kue61a] Edward J. Kuenzler. Phosphorus budget of a mussel population. Limnology and Oceanography, 6(4):400–415, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Kuenzler:1961:SEF

[Kue61b] Edward J. Kuenzler. Structure and energy flow of a mussel population in a Georgia salt marsh. Limnology and Oceanography, 6(2): 191–204, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Kurenkov:1966:IVA

[Kur66] I. I. Kurenkov. The influence of volcanic ashfall on biological processes in a lake. Limnology and Oceanography, 11(3):426–429, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Kuznetsov:1968:RSR

[Kuz68] S. I. Kuznetsov. Recent studies on the role of microorganisms in the cycling of substances in lakes. Limnology and Oceanography, 13(2): 211–224, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Kevern:1966:UAS

[KWV66] N. R. Kevern, J. L. Wilhm, and G. M. Van Dyne. Use of artificial substrata to estimate the productivity of periphyton. Limnology and Oceanography, 11(4):499–502, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Kim:1968:DMS

[KZ68] Young Sik Kim and Harry Zeitlin. The determination of molybdenum in seawater. Limnology and Oceanography, 13(3):534–537, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Lackey:1961:BSE

[Lac61] James B. Lackey. Bottom sampling and environmental niches. Limnology and Oceanography, 6(3):271–279, April 1961. CODEN LI-OCAH. ISSN 0024-3590.

LaFond:1963:DTS

[LaF63] E. C. LaFond. Detailed temperature structures of the sea off baja California. Limnology and Oceanography, 8(4):417–425, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Lagler:1968:BRG

[Lag68] Karl F. Lagler. Book review: Gerking, S. D. [Ed.]. 1967. The Biological Basis of Freshwater Fish Production. Wiley and Sons, New York. xiv + 495 p. Limnology and Oceanography, 13(2):398, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Lammerst:1962:DGS

[Lam62] William T. Lammerst. Density gradient separation of plankton and clay from river water. Limnology and Oceanography, 7(2):224–229, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Lammers:1967:PCA

[Lam67] William Tuthill Lammers. Photosynthesis by Chlorella after densitygradient centrifugation. Limnology and Oceanography, 12(1):148– 150, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Lance:1963:STS

[Lan63] Joan Lance. The salinity tolerance of some estuarine planktonic copepods. Limnology and Oceanography, 8(4):440–449, April 1963. CO-DEN LIOCAH. ISSN 0024-3590.

LaRow:1968:PDR

[LaR68a] Edward J. LaRow. A persistent diurnal rhythm in Chaoborus larvae. I. The nature of the rhythmicity. Limnology and Oceanography, 13 (2):250–256, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Larsen:1968:BRM

[Lar68b] L. H. Larsen. Book review: Macmillan, D. H. 1966. Tides. American Elsevier, New York. 240 p. $9.50. Limnology and Oceanography, 13 (1):207–208, April 1968. CODEN LIOCAH. ISSN 0024-3590.

LaRow:1969:PDR

[LaR69] Edward J. LaRow. A persistent diurnal rhythm in Chaoborus larvae. II. Ecological significance. Limnology and Oceanography, 14(2):213– 218, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Lauff:1962:PAS

[Lau62] George H. Lauff. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 7(1):110–114, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Lauff:1963:PAS

[Lau63] George H. Lauff. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 8(1):130–132, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1962:VDP

[LB62] D. A. Livingstone and J. C. Boykin. Vertical distribution of phosphorus in linsley pond mud. Limnology and Oceanography, 7(1):57–62, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Lathbury:1960:SOC

[LBL60] Alison Lathbury, Reid Bryson, and Bernhard Lettau. Some observations of currents in the hypolimnion of Lake Mendota. Limnology and Oceanography, 5(4):409–413, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Lind:1969:CUL

[LC69] Owen T. Lind and Robert S. Campbell. Comments on the use of liquid scintillation for routine determination of ^{14}C activity in production studies. Limnology and Oceanography, 14(5):787–789, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Lauff:1961:MSB

[LCEP61] George H. Lauff, Kenneth W. Cummins, Clyde H. Eriksen, and Michael Parker. A method for sorting bottom fauna samples by elutriation. Limnology and Oceanography, 6(4):462–466, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Lee:1961:OIW

[Lee61] Owen S. Lee. Observations on internal waves in shallow water. Limnology and Oceanography, 6(3):312–321, April 1961. CODEN LI-OCAH. ISSN 0024-3590.

Lewis:1967:ESC

[Lew67] Alan G. Lewis. An enrichment solution for culturing the early developmental stages of the planktonic marine copepod Euchaeta japonica Marukawa. Limnology and Oceanography, 12(1):147–148, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Lewin:1969:BRL

[Lew69a] Ralph A. Lewin. Book review: Levring, T., H. A. HOPPE, and O. J. SCHMID. 1969. Marine algae. A survey of research and utilization. Botanica Marina Handbooks, v. 1. Cram, deGruyter and Co., Hamburg. 421 p. DM140. Limnology and Oceanography, 14(6):967–968, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1969:FAC

[Lew69b] R. W. Lewis. The fatty acid composition of Arctic marine phytoplankton and zooplankton with special reference to minor acids. Limnology and Oceanography, 14(1):35–40, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Likens:1962:MRW

[LH62] Gene E. Likens and Arthur D. Hasler. Movements of radiosodium (Na^{24}) within an ice-covered lake. Limnology and Oceanography, 7 (1):48–56, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Lidz:1965:SEF

[Lid65] Louis Lidz. Sedimentary environment and Foraminiferal parameters: Nantucket Bay, Massachusetts. Limnology and Oceanography, 10(3): 392–402, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Lidz:1966:PFW

[Lid66] Louis Lidz. Planktonic Foraminifera in the water column of the mainland shelf off Newport Beach, California. Limnology and Oceanography, 11(2):257–263, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Lighthart:1963:SRB

[Lig63] Bruce Lighthart. Sulfate-reducing bacteria in San Vicente Reservoir, San Diego Countty, California. Limnology and Oceanography, 8(3): 349–351, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1965:SHW

[Liv65] D. A. Livingstone. Sedimentation and the history of water level change in Lake Tanganyika. Limnology and Oceanography, 10(4): 607–610, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1967:UFT

[Liv67] D. A. Livingstone. The use of filament tape in raising long cores from soft sediment. Limnology and Oceanography, 12(2):346–348, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Likens:1969:MAA

[LJ69] Gene E. Likens and Noye M. Johnson. Measurement and analysis of the annual heat budget for the sediments in two Wisconsin lakes. Limnology and Oceanography, 14(1):115–135, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Little:1964:DPM

[LM64] Frank J. Little, Jr. and Bill Mullins. Diving plate modification of Blake (beam) trawl, for deep-sea sampling. Limnology and Oceanography, 9(1):148–150, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Lear:1962:BRR

[LO62a] Donald W. Lear, Jr. and Carl H. Oppenheimer, Jr. Biological removal of radioisotopes ^{90}Sr and ^{90}Y from sea water by marine microorganisms. Limnology and Oceanography, 7(S1):xliv–lxii, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Lear:1962:CMC

[LO62b] Donald W. Lear, Jr. and Carl H. Oppenheimer, Jr. Consumption of microorganisms by the copepod tigriopus californicus. Limnology and Oceanography, 7(S1):lxiii–lxv, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Longhurst:1968:DLP

[Lon68] Alan R. Longhurst. Distribution of the larvae of Pleuroncodes planipes in the California Current. Limnology and Oceanography, 13(1): 143–155, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1963:DVP

[Lor63] Carl J. Lorenzen. Diurnal variation in photosynthetic activity of natural phytoplankton populations. Limnology and Oceanography, 8 (1):56–62, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1965:NCP

[Lor65] Carl J. Lorenzen. A note on the chlorophyll and phaeophytin content of the chlorophyll maximum. Limnology and Oceanography, 10(3): 482–483, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1967:DCP

[Lor67] Carl J. Lorenzen. Determination of chlorophyll and pheo-pigments: spectrophotometric equations. Limnology and Oceanography, 12(2): 343–346, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1968:CCR

[Lor68] Carl J. Lorenzen. Carbon/chlorophyll relationships in an upwelling area. Limnology and Oceanography, 13(1):202–204, April 1968. CO-DEN LIOCAH. ISSN 0024-3590.

Lovett:1968:VTG

[Lov68a] J. R. Lovett. Vertical temperature gradient variations related to current shear and turbulence. Limnology and Oceanography, 13(1): 127–142, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Lovett:1968:DSS

[Lov68b] Jack R. Lovett. Determination of salinity from simultaneous measurements of sound velocity, temperature, and pressure. Limnology and Oceanography, 13(3):557–559, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Lie:1965:DCS

[LP65] Ulf Lie and Mario M. Pamatmat. Digging characteristics and sampling efficiency of the 0.1 m^{2} van Veen grab. Limnology and Oceanography, 10(3):379–384, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Longhurst:1967:SUF

[LS67] Alan R. Longhurst and Don L. R. Seibert. Skill in the use of Folsom’s plankton sample splitter. Limnology and Oceanography, 12(2):334– 335, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Levandowsky:1969:CPV

[LS69] Michael Levandowsky and John Sibert. A criticism of Patten and Van Dyne’s productivity model. Limnology and Oceanography, 14 (2):311–312, March 1969. CODEN LIOCAH. ISSN 0024-3590. See reply [Pat69].

Lasker:1965:MES

[LT65] Reuben Lasker and Gail H. Theilacker. Maintenance of euphausiid shrimps in the laboratory. Limnology and Oceanography, 10(2):287– 288, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Ludlam:1967:SCL

[Lud67] Stuurt D. Ludlam. Sedimentation in Cayuga Lake, New York. Limnology and Oceanography, 12(4):618–632, April 1967. CODEN LI-OCAH. ISSN 0024-3590.

Ludlam:1969:FGL

[Lud69] Stuart D. Ludlam. Fayetteville Green Lake, New York. III. The laminated sediments. Limnology and Oceanography, 14(6):848–857, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Lund:1962:CCC

[Lun62] J. W. G. Lund. Concerning a counting chamber for nannoplankton described previously. Limnology and Oceanography, 7(2):261–262, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Lyman:1969:RSC

[Lym69] John Lyman. Redefinition of salinity and chlorinity. Limnology and Oceanography, 14(6):928–929, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Lynts:1966:VFS

[Lyn66] George W. Lynts. Variation of Foraminiferal standing crop over short lateral distances in Buttonwood Sound, Florida Bay. Limnology and Oceanography, 11(4):562–566, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Laevastu:1965:NOC

[LZS65] Taivo Laevastu, Harry Zeitlin, and Moon Ki Song. Notes on oxygen consumption in seawater. Limnology and Oceanography, 10(1):144– 146, April 1965. CODEN LIOCAH. ISSN 0024-3590.

MacPhee:1961:BAP

[Mac61] Craig MacPhee. Bioassay of algal production in chemically altered waters. Limnology and Oceanography, 6(4):416–422, April 1961. CO-DEN LIOCAH. ISSN 0024-3590.

Maciolek:1963:PFA

[Mac63] John A. Maciolek. Pressure-filtration apparatus. Limnology and Oceanography, 8(2):301–302, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Macintyre:1964:BDQ

[Mac64] R. J. Macintyre. A box dredge for quantitative sampling of benthic organisms. Limnology and Oceanography, 9(3):460–461, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Macisaac:1967:ADT

[Mac67] Jane J. Macisaac. Ammonia determinations by two methods in the northeast equatorial Pacific Ocean. Limnology and Oceanography, 12 (4):552–554, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Mackereth:1969:SCS

[Mac69] F. J. H. Mackereth. A short core sampler for subaqueous deposits. Limnology and Oceanography, 14(1):145–151, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Maddux:1966:LM

[Mad66] William S. Maddux. A 4π light meter. Limnology and Oceanography, 11(1):136–137, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Mairs:1966:TAA

[Mai66] Donald F. Mairs. A total alkalinity atlas for Maine lake waters. Limnology and Oceanography, 11(1):68–72, April 1966. CODEN LI-OCAH. ISSN 0024-3590.

Maitland:1969:SCS

[Mai69] Peter S. Maitland. A simple corer for sampling sand and finer sediments in shallow water. Limnology and Oceanography, 14(1):151–156, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Mangum:1964:SSM

[Man64] Charlotte Preston Mangum. Studies on speciation in maldanid polychaetes of the North American Atlantic Coast. II. Distribution and competitive interaction of five sympatric species. Limnology and Oceanography, 9(1):12–26, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Mandelli:1967:EPA

[Man67a] Enrique F. Mandelli. Enhanced photosynthetic assimilation ratios in Antarctic polar front (convergence) diatoms. Limnology and Oceanography, 12(3):484–491, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Mandelli:1967:ILO

[Man67b] Enrique F. Mandelli. Index to Limnology and Oceanography. Limnology and Oceanography, 12(3):729–733, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Manny:1969:RBO

[Man69] Bruce A. Manny. The relationship between organic nitrogen and the carotenoid to chlorophyll-a ratio in five freshwater phytoplankton species. Limnology and Oceanography, 14(1):69–79, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Marshall:1960:SNR

[Mar60] Nelson Marshall. Studies of the Niantic River, Connecticut with special reference to the bay scallop, Aequipecten irradians. Limnology and Oceanography, 5(1):86–105, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Margalef:1961:CSP

[Mar61] Ramon Margalef. Communication of structure in planktonic populations. Limnology and Oceanography, 6(2):124–128, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Martin:1965:PZR

[Mar65] John H. Martin. Phytoplankton-zooplankton relationships in Narragansett Bay. Limnology and Oceanography, 10(2):185–191, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Marshall:1966:OVD

[Mar66] Harold G. Marshall. Observations on the vertical distribution of coccolithophores in the northwestern Sargasso Sea. Limnology and Oceanography, 11(3):432–435, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Marshall:1967:RSE

[Mar67a] J. S. Marshall. Radiation stress in exploited Daphnia populations. Limnology and Oceanography, 12(3):154–158, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Martin:1967:NPD

[Mar67b] John W. Martin. New plastic drift card. Limnology and Oceanography, 12(4):706–707, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Marshall:1968:CNS

[Mar68a] Harold G. Marshall. Coccolithophores in the northwest Sargasso Sea. Limnology and Oceanography, 13(2):370–376, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Martin:1968:PZR

[Mar68b] John H. Martin. Phytoplankton-zooplankton relationships in Narragansett Bay. III. Seasonal changes in zooplankton excretion rates in relation to phytoplankton abundance. Limnology and Oceanography, 13(1):63–71, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Matthiessen:1960:IZP

[Mat60a] George C. Matthiessen. Intertidal zonation in populations of Mya arenaria. Limnology and Oceanography, 5(4):381–388, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Matthiessen:1960:OES

[Mat60b] George C. Matthiessen. Observations on the ecology of the soft clam, Mya arenaria, in a salt pond. Limnology and Oceanography, 5(3): 291–300, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Mauchline:1965:WE

[Mau65] J. Mauchline. Wandering euphausiids. Limnology and Oceanography, 10(2):300, April 1965. CODEN LIOCAH. ISSN 0024-3590.

May:1960:SCC

[May60] Billie Z. May. Stabilization of the carbohydrate content of sea water samples. Limnology and Oceanography, 5(3):342–343, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Mazeika:1967:TDE

[Maz67] P. A. Mazeika. Thermal domes in the eastern tropical Atlantic Ocean. Limnology and Oceanography, 12(4):537–539, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Mcleod:1965:GES

[MB65] G. C. Mcleod and F. Bobblis, Jr. A graphite electrode system for measuring dissolved oxygen. Limnology and Oceanography, 10(1): 146–149, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1967:LCG

[MB67] Michael M. Mullin and Elaine R. Brooks. Laboratory culture, growth rate, and feeding behavior of a planktonic marine copepod. Limnology and Oceanography, 12(4):657–666, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Menzel:1965:MTP

[MC65] David W. Menzel and Nathaniel Corwin. The measurement of total phosphorus in seawater based on the liberation of organically bound fractions by persulfate oxidation. Limnology and Oceanography, 10 (2):280–282, April 1965. CODEN LIOCAH. ISSN 0024-3590.

McAllister:1961:DFM

[McA61a] C. D. McAllister. Decontamination of filters in the ^{14}C method of measuring marine photosynthesis. Limnology and Oceanography, 6 (4):447–450, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Mcallister:1961:OVP

[Mca61b] C. D. Mcallister. Observations on the variation of planktonic photosynthesis with light intensity, using both the O_{2} and C^{14} methods. Limnology and Oceanography, 6(4):483–484, April 1961. CODEN LI-OCAH. ISSN 0024-3590.

Mcallister:1963:MDV

[Mca63] C. D. Mcallister. Measurement of diurnal variation in productivity at ocean station “p”. Limnology and Oceanography, 8(2):289–292, April 1963. CODEN LIOCAH. ISSN 0024-3590.

McFadden:1965:SEE

[MCA65] James T. McFadden, Edwin L. Cooper, and John K. Andersen. Some effects of environment on egg production in brown trout (Salmo trutta). Limnology and Oceanography, 10(1):88–95, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

McConnell:1962:PRC

[McC62] William J. McConnell. Productivity relations in carboy microcosms. Limnology and Oceanography, 7(3):335–343, April 1962. CODEN LIOCAH. ISSN 0024-3590.

McConnell:1965:RHG

[McC65] William J. McConnell. Relationship of herbivore growth to rate of gross photosynthesis in microcosms. Limnology and Oceanography, 10(4):539–543, April 1965. CODEN LIOCAH. ISSN 0024-3590.

McCauley:1966:BEO

[McC66] Rita N. McCauley. The biological effects of oil pollution in a river. Limnology and Oceanography, 11(4):475–486, April 1966. CODEN LIOCAH. ISSN 0024-3590.

McConnell:1968:LEO

[McC68] William J. McConnell. Limnological effects of organic extracts of litter in a southwestern impoundment. Limnology and Oceanography, 13(2):343–349, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Moshiri:1969:REE

[MCC69] Gerald A. Moshiri, Kenneth W. Cummins, and Robert R. Costa. Respiratory energy expenditure by the predaceous zooplankter Leptodora kindtii (Focke) (Crustacea: Cladocera). Limnology and Oceanography, 14(4):475–484, July 1969. CODEN LIOCAH. ISSN 0024-3590.

McGill:1964:OPD

[MCK64] David A. McGill, Nathaniel Corwin, and Bostwick H. Ketchum. Organic phosphorus in the deep water of the western North Atlantic. Limnology and Oceanography, 9(1):27–34, April 1964. CODEN LI-OCAH. ISSN 0024-3590.

McLaren:1965:SRB

[McL65] Ian A. McLaren. Some relationships between temperature and egg size, body size, development rate, and fecundity, of the copepod pseudocalanus. Limnology and Oceanography, 10(4):528–538, April 1965. CODEN LIOCAH. ISSN 0024-3590.

McMahon:1967:ILA

[McM67] J. W. McMahon. The influence of light and acid on the measurement of ferrous iron in lake water. Limnology and Oceanography, 12(3): 437–442, April 1967. CODEN LIOCAH. ISSN 0024-3590.

McMahon:1969:ADV

[McM69] J. W. McMahon. The annual and diurnal variation in the vertical distribution of acid-soluble ferrous and total iron in a small dimictic lake. Limnology and Oceanography, 14(3):357–367, May 1969. CO-DEN LIOCAH. ISSN 0024-3590.

McNabb:1960:EFP

[McN60] Clarence D. McNabb. Enumeration of freshwater phytoplankton concentrated on the membrane filter. Limnology and Oceanography, 5 (1):57–61, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Malone:1967:TSE

[MD67] Philip G. Malone and J. Robert Dodd. Temperature and salinity effects on calcification rate in Mytilus edulis and its paleoecological implications. Limnology and Oceanography, 12(3):432–436, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Merna:1962:QSO

[Mer62] James W. Merna. Quantitative sampling with the orange-peel dredge. Limnology and Oceanography, 7(3):432–433, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Mcgowan:1964:MHF

[MF64] John A. Mcgowan and Vernie J. Fraundorf. A modified heavy fraction zooplankton sorter. Limnology and Oceanography, 9(1):152–155, April 1964. CODEN LIOCAH. ISSN 0024-3590.

McGowan:1966:RBS

[MF66] John A. McGowan and Vernie J. Fraundorf. The relationship between size of net used and estimates of zooplankton diversity. Limnology and Oceanography, 11(4):456–469, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Martin:1962:UAR

[MG62] DeCourcey Martin and Edward D. Goldberg. Uptake and assimilation of radiostrontium by Pacific mackerel. Limnology and Oceanography, 7(S1):lxxvi–lxxxi, April 1962. CODEN LIOCAH. ISSN 0024-

Menzel:1966:DOD

[MG66] David W. Menzel and John J. Goering. The distribution of organic detritus in the ocean. Limnology and Oceanography, 11(3):333–337, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Merlini:1965:SMC

[MGPB65] Margaret Merlini, F. Girardi, R. Pietra, and A. Brazzelli. The stable manganese content of molluscs from Lake Maggiore determined by activation analysis. Limnology and Oceanography, 10(3):371–378, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Mclntire:1964:PPL

[MGPW64] C. David Mclntire, Robert L. Garrison, Harry K. Phinney, and Charles E. Warren. Primary production in laboratory streams. Limnology and Oceanography, 9(1):92–102, April 1964. CODEN LI-OCAH. ISSN 0024-3590.

McNaught:1961:SSF

[MH61] Donald C. McNaught and Arthur D. Hasler. Surface schooling and feeding behavior in the white bass, Roccus chrysops (Rafinesque), in Lake Mendota. Limnology and Oceanography, 6(1):53–60, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Marshall:1962:DMT

[MH62] Nelson Marshall and Steacy D. Hicks. Drift of medusae and their distribution in relation to the hydrography of the Niantic River, Connecticut. Limnology and Oceanography, 7(2):268–269, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Morita:1964:TEG

[MH64] Richard Y. Morita and Roger D. Haight. Temperature effects on the growth of an obligate psychrophilic marine bacterium. Limnology and Oceanography, 9(1):103–106, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Millero:1969:PMV

[Mil69] Frank J. Millero. The partial molal volumes of ions in seawater. Limnology and Oceanography, 14(3):376–385, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Maddux:1964:SIT

[MJ64] William S. Maddux and Raymond F. Jones. Some interactions of temperature, light intensity, and nutrient concentration during the continuous culture of nitzschia closterium and tetraselmis sp. Limnology and Oceanography, 9(1):79–86, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Maddux:1965:LSP

[MK65] William S. Maddux and John W. Kanwisher. An In Situ particle counter. Limnology and Oceanography, 10(S1):R162–R168, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Mulligan:1968:AEP

[MK68] Hugh F. Mulligan and John M. Kingsbury. Application of an electronic particle counter in analyzing natural populations of phytoplankton. Limnology and Oceanography, 13(3):499–506, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Mishima:1963:ERL

[MO63] Jiro Mishima and Eugene P. Odum. Excretion rate of ^{65}Zn by littorina irrorata in relation to temperature and body size. Limnology and Oceanography, 8(1):39–44, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Mohammad:1961:NRB

[Moh61] Murad-B. M. Mohammad. A new record for Balanus balanoides (L.). Limnology and Oceanography, 6(4):488, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Mokyevsky:1960:GZM

[Mok60] O. B. Mokyevsky. Geographical zonation of marine littoral types. Limnology and Oceanography, 5(4):389–396, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Moore:1963:RMF

[Moo63] Johnes K. Moore. Refinement of a method for filtering and preserving marine phytoplankton on a mersbrane filter. Limnology and Oceanography, 8(2):304–305, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Morris:1960:TSS

[Mor60] Robert W. Morris. Temperature, salinity, and southern limits of three species of Pacific cottid fishes. Limnology and Oceanography, 5 (3):175–179, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Moss:1967:NEC

[Mos67a] Brian Moss. A note on the estimation of chlorophyll-a in freshwater algal communities. Limnology and Oceanography, 12(2):340–342, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Moss:1967:SME

[Mos67b] Brian Moss. A spectrophotometric method for the estimation of percentage degradation of chlorophylls to pheo-pigments in extracts of algae. Limnology and Oceanography, 12(2):335–340, April 1967. CO-DEN LIOCAH. ISSN 0024-3590.

Moss:1969:LAG

[Mos69] Brian Moss. Limitation of algal growth in some Central African waters. Limnology and Oceanography, 14(4):591–601, July 1969. CO-DEN LIOCAH. ISSN 0024-3590.

Marvin:1965:SAN

[MP65] Kenneth T. Marvin and Raphael R. Proctor, Jr. Stabilizing the ammonia=-nitrogen content of estuarine and coastal waters by freezing. Limnology and Oceanography, 10(2):288–289, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

McAllister:1961:MPP

[MPSS61] C. D. McAllister, T. R. Parsons, K. Stephens, and J. D. H. Strickland. Measurements of primary production in coastal sea water using a large-volume plastic sphere. Limnology and Oceanography, 6(3):237– 258, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Menzel:1964:CPO

[MR64] David W. Menzel and John H. Ryther. The composition of particulate organic matter in the western North Atlantic. Limnology and Oceanography, 9(2):179–186, April 1964. CODEN LIOCAH. ISSN 0024-3590.

McMahon:1965:FRD

[MR65] J. W. McMahon and F. H. Rigler. Feeding rate of Daphnia magna Straus in different foods labeled with radioactive phosphorus. Limnology and Oceanography, 10(1):105–113, April 1965. CODEN LI-OCAH. ISSN 0024-3590.

Merifield:1966:SMS

[MR66] Paul M. Merifield and Donald M. Rosencrantz. A simple method for surveying a small area underwater. Limnology and Oceanography, 11 (3):408–409, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Menzies:1968:LLU

[MR68] Robert J. Menzies and Gilbert T. Rowe. The lubs, a large undisturbed bottom sampler. Limnology and Oceanography, 13(4):708– 714, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Mcallister:1961:LAU

[MS61] C. D. Mcallister and J. D. H. Strickland. Light attenuators for use in phytoplankton photosynthesis studies. Limnology and Oceanography, 6(2):226–228, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Menzel:1962:OAS

[MS62a] David W. Menzel and Jane P. Spaeth. Occurrence of ammonia in Sargasso Sea waters and in rain water at Bermuda. Limnology and Oceanography, 7(2):159–162, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Menzel:1962:OIS

[MS62b] David W. Menzel and Jane P. Spaeth. Occurrence of iron in the Sargasso Sea off Bermuda. Limnology and Oceanography, 7(2):155– 158, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Menzel:1962:OVB

[MS62c] David W. Menzel and Jane P. Spaeth. Occurrence of vitamin B_{12} in the Sargasso Sea. Limnology and Oceanography, 7(2):151–154, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Megard:1966:MLL

[MS66] Robert O. Megard and Joseph Shapiro. A mobile limnological laboratory. Limnology and Oceanography, 11(3):420–422, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1966:RBC

[MSE66] M. M. Mullin, P. R. Sloan, and R. W. Eppley. Relationship between carbon content, cell volume, and area in phytoplankton. Limnology and Oceanography, 11(2):307–311, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Mauchline:1964:ART

[MT64] John Mauchline and Angela M. Taylor. The accumulation of radionuclides by the thornback ray, Raia clavata L., in the Irish Sea. Limnology and Oceanography, 9(3):303–310, April 1964. CODEN LI-OCAH. ISSN 0024-3590.

Mauchline:1964:RB

[MTR64] John Mauchline, Angela M. Taylor, and Eric B. Ritson. The radioecology of a beach. Limnology and Oceanography, 9(2):187–194, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1963:SFA

[Mul63] Michael M. Mullin. Some factors affecting the feeding of marine copepods of the genus Calanus. Limnology and Oceanography, 8(2): 239–250, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Mulford:1964:IIC

[Mul64] Richard A. Mulford. Investigations of inner continental shelf waters off lower Chesapeake Bay. Part V, Seasonality of the diatom genus Chaetoceros. Limnology and Oceanography, 9(3):385–390, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Muller:1965:ASD

[Mül65a] K. Müller. An automatic stream drift sampler. Limnology and Oceanography, 10(3):483–485, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1965:SFPa

[Mul65b] Michael M. Mullin. Size fractionation of particulate organic carbon in the surface waters of the western Indian Ocean. Limnology and Oceanography, 10(3):459–462, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1965:SFPb

[Mul65c] Michael M. Mullin. Size fractionation of particulate organic carbon in the surface waters of the western Indian Ocean: addendum. Limnology and Oceanography, 10(4):610–611, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Mundie:1964:SCE

[Mun64] J. H. Mundie. A sampler for catching emerging insects and drifting materials in streams. Limnology and Oceanography, 9(3):456–459, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1962:TUN

[Mur62] D. G. Murphy. Three undescribed nematodes from the coast of Oregon. Limnology and Oceanography, 7(3):386–389, April 1962. CO-DEN LIOCAH. ISSN 0024-3590.

Menzel:1964:MDO

[MV64] David W. Menzel and Ralph F. Vaccaro. The measurement of dissolved organic and particulate carbon in seawater. Limnology and Oceanography, 9(1):138–142, April 1964. CODEN LIOCAH. ISSN 0024-3590.

McIntire:1969:LMS

[MW69] C. David McIntire and Barry L. Wulff. A laboratory method for the study of marine benthic diatoms. Limnology and Oceanography, 14 (5):667–678, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Noble:1961:PPF

[NA61] Vincent E. Noble and John C. Ayers. A portable photocell fluorometer for dilution measurements in natural waters. Limnology and Oceanography, 6(4):457–461, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Nalewajko:1966:PEV

[Nal66] C. Nalewajko. Photosynthesis and excretion in various planktonic algae. Limnology and Oceanography, 11(1):1–10, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Natarajan:1966:BDT

[ND66] K. V. Natarajan and R. C. Dugdale. Bioassay and distribution of thiamine in the sea. Limnology and Oceanography, 11(4):621–629, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Neess:1962:NML

[NDDG62] John C. Neess, Richard C. Dugdale, Vera A. Dugdale, and John J. Goering. Nitrogen metabolism in lakes I. measurement of nitrogen fixation with ^{15}N. Limnology and Oceanography, 7(2):163–169, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Newhouse:1967:SDF

[NDT67] Jan Newhouse, Maxwell S. Doty, and Roy T. Tsuda. Some diurnal features of a neritic surface plankton population. Limnology and Oceanography, 12(2):207–212, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Neshyba:1967:PLS

[Nes67] Steve Neshyba. Pulsed light stimulation of marine bioluminescence In Situ. Limnology and Oceanography, 12(2):222–235, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Neumann:1966:OCE

[Neu66] A. Conrad Neumann. Observations on coastal erosion in Bermuda and measurements of the boring rate of the sponge, Cliona lampa. Limnology and Oceanography, 11(1):92–108, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Nicholson:1964:WPI

[NGL^{+}64] H. Page Nicholson, Alfred R. Grzenda, Gerald J. Lauer, William S. Cox, and John I. Teasley. Water pollution by insecticides in an agricultural river basin. I. Occurrence of insecticides in river and treated municipal water. Limnology and Oceanography, 9(3):310–317, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Nichols:1960:REG

[Nic60] David Nichols. Russian-English glossary of hydrobiology. Limnology and Oceanography, 5(3):242–243, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1965:DAA

[Nie65] E. Steemann Nielsen. On the determination of the activity in 14C-ampoules for measuring primary production. Limnology and Oceanography, 10(S1):R247–R252, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Niskin:1964:RTM

[Nis64] Shale J. Niskin. A reversing-thermometer mechanism for attachment to oceanographic devices. Limnology and Oceanography, 9(4):591– 594, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Nixon:1969:SM

[Nix69] Scott W. Nixon. A synthetic microcosm. Limnology and Oceanography, 14(1):142–145, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Noble:1966:OFO

[Nob66] Vincent E. Noble. Observations of the fall overturn of Lake Michigan. Limnology and Oceanography, 11(3):413–415, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Northcote:1964:UHF

[Nor64] T. G. Northcote. Use of a high-frequency echo sounder to record distribution and migration of chaoborus larvae. Limnology and Oceanography, 9(1):87–91, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Norris:1965:OCI

[Nor65] Kenneth S. Norris. Oceanic coordinate index. Limnology and Oceanography, 10(4):616, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Norris:1969:PPF

[Nor69] Dean R. Norris. Possible phagotrophic feeding in ceratium lunula schimper. Limnology and Oceanography, 14(3):448–449, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Nriagu:1968:SMS

[Nri68] Jerome O. Nriagu. Sulfur metabolism and sedimentary environment: Lake Mendota, Wisconsin. Limnology and Oceanography, 13(3):430– 439, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Nelson:1962:RDP

[NS62] Daniel J. Nelson and Donald C. Scott. Role of detritus in the productivity of a rock-outcrop community in a Piedmont stream. Limnology and Oceanography, 7(3):396–413, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Nutting:1961:DSI

[Nut61] Wm B. Nutting. A device for successive ice penetration. Limnology and Oceanography, 6(2):234–235, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Ogawa:1969:INH

[OC69] Roann E. Ogawa and John F. Carr. The influence of nitrogen on heterocyst production in blue-green algae. Limnology and Oceanography, 14(3):342–351, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Ochinero:1967:BRS

[Och67] Robert V. Ochinero. Book review: Stewart, H. B., Jr. 1966. Deep Challenge. D. Van Nostrand Company, Inc., Princeton, New Jersey. 202 p. $5.95. Limnology and Oceanography, 12(3):183, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Olsen:1967:WBD

[OCO67] Sigurd Olsen, Diptiman Chakravarti, and Paul R. Olson. Water, bottom deposits, and zooplankton of Fern Lake, Washington. Limnology and Oceanography, 12(3):392–404, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Oppenheimer:1963:MDO

[OCV63] C. H. Oppenheimer, E. F. Corcoran, and J. Vanarman. Method for the determination of organic carbon in seawater. Limnology and Oceanography, 8(4):487–488, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Odum:1968:ESF

[Odu68] William E. Odum. The ecological significance of fine particle selection by the striped mullet Mugil cephalus. Limnology and Oceanography, 13(1):92–98, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1968:PDN

[OH68] Akira Otsuki and Takahisa Hanya. On the production of dissolved nitrogen-rich organic matter. Limnology and Oceanography, 13(1): 183–185, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Okubo:1969:DCS

[OK69] Akira Okubo and Michael J. Karweit. Diffusion from a continuous source in a uniform shear flow. Limnology and Oceanography, 14(4): 514–520, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Olive:1967:VDC

[OM67] John H. Olive and John H. Morrison. Variations in distribution of 14C in cell extracts of phytoplankton living under natural conditions. Limnology and Oceanography, 12(3):383–391, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Osterberg:1964:ZER

[OPP64] Charles Osterberg, June Pattullo, and William Pearcy. Zinc 65 in euphausiids as related to Columbia River water off the Oregon coast. Limnology and Oceanography, 9(2):249–257, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Orians:1968:BRD

[Ori68] Gordon H. Orians. Book review: Dunbar, M. J. 1968. Ecological Development in Polar Regions, A Study in Evolution. Concepts in modern biology. Prentice-Hall, Englewood Cliffs, N. J. 119 p. $4.95. Limnology and Oceanography, 13(3):566–568, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Ormerod:1966:SMD

[Orm66] J. G. Ormerod. A simple method for the detection of oxidized manganese in particles on membrane filters. Limnology and Oceanography, 11(4):635–636, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Orton:1962:FEL

[Ort62] Grace L. Orton. Fish eggs and larvae in the plankton samples. Limnology and Oceanography, 7(S1):xxxii–xxxviii, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Osterberg:1962:CSE

[Ost62] Charles Osterberg. ^{65}Zn content of salps and euphaushds. Limnology and Oceanography, 7(4):478–479, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Owen:1963:TIS

[Owe63] R. W. Owen, Jr. Trailing incubator for studies of C^{14} uptake by phytoplankton. Limnology and Oceanography, 8(2):297–298, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Owen:1969:BRB

[Owe69] David M. Owen. Book review: Balder, A. P. 1968. The complete manual of skin diving. Macmillan, New York. 302 p. $6.95. Limnology and Oceanography, 14(1):171, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Powers:1960:WTS

[PA60] Charles F. Powers and John C. Ayers. Water transport studies in the Straits of Mackinac region of Lake Huron. Limnology and Oceanography, 5(1):81–86, April 1960. CODEN LIOCAH. ISSN 0024-3590. See comment [Rod61].

Paasche:1968:ETL

[Paa68] E. Paasche. The effect of temperature, light intensity, and photoperiod on coccolith formation. Limnology and Oceanography, 13(1): 178–181, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Paine:1966:EBC

[Pai66] Robert T. Paine. Endothermy in bomb calorimetry. Limnology and Oceanography, 11(1):126–129, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Paine:1967:BRM

[Pai67] Robert T. Paine. Book review: MacArthur, R. H., and J. H. Connell. 1967. The biology of populations. Wiley, New York. 216 p. $5.95. Limnology and Oceanography, 12(4):727, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Paine:1968:BRN

[Pai68] Robert T. Paine. Book review: Nicol, J. A. C. 1967. The Biology of Marine Animals, 2nd Ed. Wiley and Sons, New York. 699 p. $15.95. Limnology and Oceanography, 13(1):207, April 1968. CODEN LI-OCAH. ISSN 0024-3590.

Paine:1969:BRK

[Pai69] R. T. Paine. Book review: Kormondy, E. J. 1969. Concepts of ecology. Prentice-Hall, Inc., Englewood Cliffs, N. J. 209 p. Hardbound, $4.95; paperback, $2.95. Limnology and Oceanography, 14(6):968– 969, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Pamatmat:1965:CFA

[Pam65] Mario M. Pamatmat. A continuous-flow apparatus for measuring metabolism of benthic communities. Limnology and Oceanography, 10(3):486–489, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Parker:1960:CBS

[Par60] Richard A. Parker. Competition between Simocephalus vetulus and Cyclops viridis. Limnology and Oceanography, 5(3):180–189, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Parker:1961:CBE

[Par61] Richard A. Parker. Competition between Eucyclops agilis and Daphnia pulex. Limnology and Oceanography, 6(3):299–301, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1965:ATD

[Par65] T. R. Parsons. An automated technique for determining the growth rate of chain-forming phytoplankton. Limnology and Oceanography, 10(4):598–602, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Park:1966:BRMb

[Par66a] Kilho Park. Book review: Miyake, Y. 1965. Elements of Geochemistry. Maruzen Co., Tokyo. 475 p. $12.00. Limnology and Oceanography, 11(3):444–445, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Park:1966:BRMa

[Par66b] Kilho Park. Book review: Miyake, Y., and T. Koyama [ed.]. 1964. Recent Researches in the Fields of Hydrosphere, Atmosphere, and Nuclear Geochemistry. Ken Sugawara Festival Volume. Maruzen Co., Tokyo. 404 p. $10.00. Limnology and Oceanography, 11(3):444–445, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Park:1966:CRP

[Par66c] Kilho Park. Columbia River plume identification by specific alkalinity. Limnology and Oceanography, 11(1):118–120, April 1966. CO-DEN LIOCAH. ISSN 0024-3590.

Park:1967:BRS

[Par67a] Kilho Park. Book review: Stumm, W. [Chairman], and R. F. Gould [Ed.]. 1967. Equilibrium concepts in natural water systems. Advances in Chemistry, Ser. 67. Am. Chem. Soc., Washington, D.C. 344 p. $8.50. Limnology and Oceanography, 12(4):726–727, April 1967. CO-DEN LIOCAH. ISSN 0024-3590.

Park:1967:NRP

[Par67b] Kilho Park. Nutrient regeneration and preformed nutrients off Oregon. Limnology and Oceanography, 12(2):353–357, April 1967. CO-DEN LIOCAH. ISSN 0024-3590.

Parker:1967:BNS

[Par67c] Bruce C. Parker. Biodialystat: new sampler for dissolved organic matter. Limnology and Oceanography, 12(4):722–723, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1968:BRM

[Par68] T. R. Parsons. Book review: Martin, D. F. 1968. Marine chemistry. Volume 1: Analytical methods. Marcel Dekker, Inc. New York. 280 p. $5.75. Limnology and Oceanography, 13(4):726, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Park:1969:OCS

[Par69] P. Kilho Park. Oceanic CO_{2} system: an evaluation of ten methods of investigation. Limnology and Oceanography, 14(2):179–186, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Patten:1961:PER

[Pat61a] Bernard Patten. Plankton energetics of Raritan Bay. Limnology and Oceanography, 6(4):369–387, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Patten:1961:NFC

[Pat61b] Bernard C. Patten. Negentropy flow in communities of plankton. Limnology and Oceanography, 6(1):26–30, April 1961. CODEN LI-OCAH. ISSN 0024-3590.

Patten:1962:IME

[Pat62] Bernard C. Patten. Improved method for estimating stability in plankton. Limnology and Oceanography, 7(2):266–268, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Patten:1969:RLS

[Pat69] Bernard C. Patten. Reply [to Levandowsky and Sibert]. Limnology and Oceanography, 14(2):313–315, March 1969. CODEN LIOCAH. ISSN 0024-3590. See [LS69].

Pautzke:1965:FW

[Pau65] Clarence F. Pautzke. Fish and wildlife. Limnology and Oceanography, 10(2):304, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Pamatmat:1969:OCS

[PB69] Mario M. Pamatmat and Karl Banse. Oxygen consumption by the seabed. II. In Situ measurements to a depth of 180 m. Limnology and Oceanography, 14(2):250–259, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Pattullo:1969:OHC

[PBK69] June G. Pattullo, Wayne V. Burt, and Sally A. Kulm. Oceanic heat content off Oregon: its variations and their causes. Limnology and Oceanography, 14(2):279–287, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Pattullo:1965:PAS

[PD65] June Pattullo and Warren Denner. Processes affecting seawater ciiaracteristics along the Oregon coast. Limnology and Oceanography, 10(3):443–450, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Pennak:1962:QZS

[Pen62] Robert W. Pennak. Quantitative zooplankton sampling in littoral vegetation areas. Limnology and Oceanography, 7(4):487–489, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Pennak:1969:CSM

[Pen69] Robert W. Pennak. Colorado semidrainage mountain lakes. Limnology and Oceanography, 14(5):720–725, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Pamatmat:1968:IMS

[PF68] Mario M. Pamatmat and Douglas Fenton. An instrument for measuring subtidal benthic metabolism In Situ. Limnology and Oceanography, 13(3):537–540, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Pearcy:1969:OSS

[PF69] W. G. Pearcy and Carl A. Forss. The oceanic shrimp Sergestes similis off the Oregon coast. Limnology and Oceanography, 14(5):755–765, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Pfanstiehl:1962:ASB

[Pfa62] Alfred Pfanstiehl. An aid for small boat maneuvers. Limnology and Oceanography, 7(3):431–432, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Pickard:1960:SOT

[PG60] G. L. Pickard and L. F. Giovando. Some observations of turbidity in British Columbia inlets. Limnology and Oceanography, 5(3):162–170, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Pikey:1963:TER

[PG63] Orrin H. Pikey and H. G. Goodell. Trace elements in recent mollusk shells. Limnology and Oceanography, 8(2):137–148, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Pearson:1960:WQU

[PH60] Erman A. Pearson and George A. Holt. Water quality and upwelling at Grays Harbor entrance. Limnology and Oceanography, 5(1):48–56, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Parker:1962:SPE

[PH62] Richard A. Parker and Donald H. Hazelwood. Some possible effects of trace elements on fresh-water microcrustacean populations. Limnology and Oceanography, 7(3):344–347, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Park:1963:RFM

[PH63] Kilho Park and Donald V. Hood. Radiometric field measurement of CO_{2} exchange from the atmosphere to the sea. Limnology and Oceanography, 8(2):287–289, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Pilkey:1966:EIE

[PH66] Orrin H. Pilkey and Robert C. Harris. The effect of intertidal environment on the composition of calcareous skeletal material. Limnology and Oceanography, 11(3):381–385, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Parker:1969:SDC

[PH69] Michael Parker and Arthur D. Hasler. Studies on the distribution of cobalt in lakes. Limnology and Oceanography, 14(2):229–241, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Phleger:1965:PMF

[Phl65] Fred B. Phleger. Patterns of marsh Foraminifera, Galveston Bay, Texas. Limnology and Oceanography, 10(S1):R169–R184, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Pierce:1964:CCO

[Pie64] Madelene E. Pierce. Careers in conservation: Opportunities in natural resources. Limnology and Oceanography, 9(1):156, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1963:CVS

[PK63] Ricardo M. Pytkowicz and I. M. Klotz. Comment on the variation of solubility with depth. Limnology and Oceanography, 8(4):486, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1967:RCP

[PK67] R. M. Pytkowicz and D. R. Kester. Relative calcium phosphate saturation in two regions of the north Pacific Ocean. Limnology and Oceanography, 12(4):714–718, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Pearson:1969:DSD

[PK69] William D. Pearson and Robert H. Kramer. A drift sampler driven by a waterwheel. Limnology and Oceanography, 14(3):462–465, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1966:RPM

[PKB66] R. M. Pytkowicz, D. R. Kester, and B. C. Burgener. Reproducibility of pH measurements in seawater. Limnology and Oceanography, 11 (3):417–419, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Platt:1969:CEE

[Pla69] Trevor Platt. The concept of energy efficiency in primary production. Limnology and Oceanography, 14(5):653–659, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Parker:1968:SST

[PLH68] Bruce C. Parker, George Leeper, and William Hurni. Sampler for studies of thin horizontal layers. Limnology and Oceanography, 13 (1):172–175, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Proctor:1962:RGS

[PLPP62] Charles M. Proctor, Andrew R. Lowrey, Emanuel Papadopulos, and John R. Penning, Jr. Response of gamma scintillation detectors for field survey use. Limnology and Oceanography, 7(3):273–279, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Phinney:1965:ETM

[PM65] Hawy K. Phinney and C. David McIntire. Effect of temperature on metabolism of periphyton communities developed in laboratory streams. Limnology and Oceanography, 10(3):341–345, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1963:EPS

[PMM63] Lawrence R. Pomeroy, Henry M. Mathews, and Hong Shik Min. Excretion of phosphate and soluble organic phosphorus compounds by zooplankton. Limnology and Oceanography, 8(1):50–55, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Pullen:1968:NSI

[PMR68] E. J. Pullen, C. R. Mock, and R. D. Ringo. A net for sampling the intertidal zone of an estuary. Limnology and Oceanography, 13(1): 200–202, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Perkins:1966:BTR

[PNH66] R. W. Perkins, J. L. Nelson, and W. L. Haushild. Behavior and transport of radionuclides in the Columbia River between Hanford and Vancouver, Washington. Limnology and Oceanography, 11(2): 235–248, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Pearcy:1968:ZMA

[PO68] William G. Pearcy and Charles L. Osterberg. Zinc-65 and manganese-54 in albacore Thunnus alalunga from the west coast of North America. Limnology and Oceanography, 13(3):490–498, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Patterson:1963:DCD

[PP63] J. Patterson and T. R. Parsons. Distribution of chlorophyll-a and degradation products in various marine materials. Limnology and Oceanography, 8(3):355–356, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Park:1967:BRF

[PP67a] Kilho Park and Daniel Panshin. Book review: Fairbridge, R. W. [Ed.]. 1966. Encyclopedia of Oceanography. Reinhold, New York. 1021 p. $25.00. Limnology and Oceanography, 12(2):364–365, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Park:1967:BRW

[PP67b] Kilho Park and Daniel Panshin. Book review: Wadachi, K. [Ed.]. 1960. Dictionary of the Oceans [In Japanese]. Tokyo-Do, Tokyo. 671 p. $5.28. Limnology and Oceanography, 12(2):364, April 1967. CO-DEN LIOCAH. ISSN 0024-3590.

Proctor:1962:GSP

[PPF62] Charles M. Proctor, Emanuel Papadopulos, and Ralph H. Firminhac. Gamma scintillation probe for field use and measurements of radiation background in Puget Sound. Limnology and Oceanography, 7(3):280–286, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Park:1962:CPI

[PPW62] Kilho Park, June G. Pattullo, and Bruce Wyatt. Chemical properties as indicators of upwelling along the Oregon coast. Limnology and Oceanography, 7(3):435–437, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Prakash:1968:IHS

[PR68] A. Prakash and M. A. Rashid. Influence of humic substances on the growth of marine phytoplankton: dinoflagellates. Limnology and Oceanography, 13(4):598–606, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Pratt:1965:WSD

[Pra65] David M. Pratt. The winter–spring diatom flowering in Narragansett Bay. Limnology and Oceanography, 10(2):173–184, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

Pratt:1966:CBS

[Pra66a] David M. Pratt. Competition between Skeletonema costatum and Olisthodiscus luteus in Narragansett Bay and in culture. Limnology and Oceanography, 11(4):447–455, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Pratt:1966:GSG

[Pra66b] R. M. Pratt. The Gulf Stream as a graded river. Limnology and Oceanography, 11(1):60–67, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Proctor:1962:SNC

[Pro62] Raphael R. Proctor, Jr. Stabilization of the nitrite content of sea water by freezing. Limnology and Oceanography, 7(4):479–480, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Pomero:1965:EPB

[PSG65] Lawrence R. Pomero, E. E. Smith, and Carol M. Grant. The exchange of phosphate between estuarine water and sediments. Limnology and Oceanography, 10(2):167–172, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Paquette:1961:ECB

[PSS61] Robert G. Paquette, Eugene L. Scott, and Paul N. Sund. An enlarged Clarke–Bumpus plankton sampler. Limnology and Oceanography, 6 (2):230–233, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Pillsbury:1969:RLC

[PST69] Dale Pillsbury, Robert L. Smith, and Ronald C. Tipper. A reliable low-cost mooring system for oceanographic instrumentation. Limnology and Oceanography, 14(2):307–311, March 1969. CODEN LI-OCAH. ISSN 0024-3590.

Packard:1968:RBS

[PT68] Theodore T. Packard and Peter B. Taylor. The relationship between succinate dehydrogenase activity and oxygen consumption in the brine shrimp Artemia salina. Limnology and Oceanography, 13 (3):552–555, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Pearcy:1969:OCE

[PTL69] William G. Pearcy, Gail H. Theilacker, and Reuben Lasker. Oxygen consumption of Euphausia pacifica: the lack of a diel rhythm or lightdark effect, with a comparison of experimental techniques. Limnology and Oceanography, 14(2):219–223, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Patten:1968:FPE

[PV68] Bernard C. Patten and George M. Van Dyne. Factorial productivity experiments in a shallow estuary: energetics of individual plankton species in mixed populations. Limnology and Oceanography, 13(2): 309–314, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Paine:1969:EGS

[PV69] Robert T. Paine and Robert L. Vadas. The effects of grazing by sea urchins, strongylocentrotus spp., on benthic algal populations. Limnology and Oceanography, 14(5):710–719, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Park:1969:ABC

[PWY69] P. Kilho Park, George R. Webster, and Roy Yamamoto. Alkalinity budget of the Columbia River. Limnology and Oceanography, 14(4): 559–567, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1962:EGD

[Pyt62] Ricardo Marcos Pytkowicz. Effect of gravity on distribution of salts in sea water. Limnology and Oceanography, 7(3):434–435, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1963:GPS

[Pyt63] Ricardo Marcos Pytkowicz. Gravity and the properties of sea water. Limnology and Oceanography, 8(2):286–287, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1965:CCS

[Pyt65] R. M. Pytkowicz. Calcium carbonate saturation in the ocean. Limnology and Oceanography, 10(2):220–225, April 1965. CODEN LI-OCAH. ISSN 0024-3590.

Robertson:1966:CSL

[RA66] Andrew Robertson and Wayne P. Alley. A comparative study of Lake Michigan macrobenthos. Limnology and Oceanography, 11(4): 576–583, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Roth:1962:ETY

[RAF^{+}62] Frank J. Roth, Jr., Donald G. Ahearn, Jack W. Fell, Samuel P. Meyers, and Sally A. Meyer. Ecology and taxonomy of yeasts isolated from various marine substrates. Limnology and Oceanography, 7(2): 178–185, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Ragotzkie:1962:EAS

[Rag62] Robert A. Ragotzkie. Effect of air stability on the development of wind waves on lakes. Limnology and Oceanography, 7(2):248–251, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Ramsey:1960:DOS

[Ram60] W. L. Ramsey. Dissolved oxygen in sweetwater lake. Limnology and Oceanography, 5(1):34–42, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Ramsey:1962:BGD

[Ram62a] W. L. Ramsey. Bubble growth from dissolved oxygen near the sea surface. Limnology and Oceanography, 7(1):1–7, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Ramsey:1962:DOS

[Ram62b] W. L. Ramsey. Dissolved oxygen in shallow near-shore water and its relation to possible bubble formation. Limnology and Oceanography, 7(4):453–461, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Rao:1968:NFG

[Rao68] Desiraju B. Rao. A note on the free gravitational oscillations in a rectangular basin of variable depth. Limnology and Oceanography, 13(2):369–370, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Rao:1969:AJB

[Rao69] D. V. Subba Rao. Asterionella japonica bloom and discoloration off Waltair, Bay of Bengal. Limnology and Oceanography, 14(4):632–634, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Rawson:1960:LCT

[Raw60] D. S. Rawson. A limnological comparison of twelve large lakes in northern Saskatchewan. Limnology and Oceanography, 5(3):195–211, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Ray:1968:BRB

[Ray68] Dixie Lee Ray. Book review: Bardach, J. 1968. Harvest of the Sea. Harper and Row, New York. 301 p. $6.95. Limnology and Oceanography, 13(2):398–399, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Raymont:1961:FIC

[RC61] J. E. G. Raymont and Robert J. Conover. Further investigations on the carbohydrate content of marine zooplankton. Limnology and Oceanography, 6(2):154–164, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Reeburgh:1964:DCU

[RC64] William S. Reeburgh and James H. Carpenter. Determination of chlorinity using a differential potentiometric end pointy. Limnology and Oceanography, 9(4):589–591, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Richards:1965:SCD

[RCBA65] Francis A. Richards, Joel D. Cline, William W. Broenkow, and Larry P. Atkinson. Some consequences of the decomposition of organic matter in Lake Nitinat, an anoxic fjord. Limnology and Oceanography, 10(S1):R185–R201, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Redfield:1961:TSL

[Red61] Alfred C. Redfield. The tidal system of Lake Maracaibo, Venezuela. Limnology and Oceanography, 6(1):1–12, April 1961. CODEN LI-OCAH. ISSN 0024-3590.

Redfield:1962:PPT

[Red62] Alfred C. Redfield. A practical portable tide gage. Limnology and Oceanography, 7(2):262–265, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Reeburgh:1967:IIW

[Ree67] W. S. Reeburgh. An improved interstitial water sampler. Limnology and Oceanography, 12(3):163–165, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Reeburgh:1969:OGC

[Ree69] William S. Reeburgh. Observations of gases in Chesapeake Bay sediments. Limnology and Oceanography, 14(3):368–375, May 1969. CO-DEN LIOCAH. ISSN 0024-3590.

Reid:1960:GSP

[Rei60] R. O. Reid. The Gulf Stream: a physical and dynamical description. Limnology and Oceanography, 5(1):111–112, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Reid:1962:CPP

[Rei62] Joseph L. Reid, Jr. On circulation, phosphate-phosphorus content, and zooplankton volumes in the upper part of the Pacific Ocean. Limnology and Oceanography, 7(3):287–306, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Reid:1963:EWC

[Rei63] Joseph L. Reid, Jr. Equatorial waters and currents at 150^{o}W in July– August 1952. Limnology and Oceanography, 8(3):366, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1960:WMV

[RH60] J. H. Ryther and E. M. Hulburt. On winter mixing and the vertical distribution of phytoplankton. Limnology and Oceanography, 5(3): 337–338, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Rona:1962:AAM

[RHMB62] Elizabeth Rona, Donald W. Hood, Lowell Muse, and Benjamin Buglio. Activation analysis of manganese and zinc in sea water. Limnology and Oceanography, 7(2):201–206, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1966:POP

[RHP^{+}66] John H. Ryther, John R. Hall, Allan K. Pease, Andrew Bakun, and Mark M. Jones. Primary organic production in relation to the chemistry and hydrography of the western Indian Ocean. Limnology and Oceanography, 11(3):371–380, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Rice:1967:BRP

[Ric67] T. R. Rice. Book review: Polikarpov, G. G. 1966. Radioecology of Aquatic Organisms [Transl. from Russian], Reinhold, New York. xxviii + 314 p. $16.50. Limnology and Oceanography, 12(2):364, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Richards:1969:BRF

[Ric69] Francis A. Richards. Book review: Forbes, L. [compiler]. 1968. Oceanography in print, a selected list of educational resources. Town of Falmouth Public Schools in Cooperation with Woods Hole Oceanographic Institution, U.S. Bureau of Commercial Fisheries, and Marine Biological Laboratory. Falmouth, Mass. 58 p. $0.75. Limnology and Oceanography, 14(1):171–173, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Rigler:1961:URI

[Rig61] F. H. Rigler. The uptake and release of inorganic phosphorus by Daphnia magna Straus. Limnology and Oceanography, 6(2):165–174, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Rigler:1964:PFT

[Rig64] F. H. Rigler. The phosphorus fractions and the turnover time of inorganic phosphorus in different types of lakes. Limnology and Oceanography, 9(4):511–518, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Rigler:1968:FOI

[Rig68] F. H. Rigler. Further observations inconsistent with the hypothesis that the molybdenum blue method measures orthophosphate in lake water. Limnology and Oceanography, 13(1):7–13, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Riley:1963:OAS

[Ril63] Gordon A. Riley. Organic aggregates in seawater and the dynamics of their formation and utilization. Limnology and Oceanography, 8 (4):372–381, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Riley:1965:MMR

[Ril65] Gordon A. Riley. A mathematical model of regional variations in plankton. Limnology and Oceanography, 10(S1):R202–R215, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Ritterberg:1965:PAA

[Rit65] S. C. Ritterberg. Principles and applications in aquatic microbiology. Limnology and Oceanography, 10(4):614–615, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Richards:1962:WCV

[RK62] A. F. Richards and G. H. Keller. Water content variability in a silty clay core from off Nova Scotia. Limnology and Oceanography, 7(3): 426–427, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Ragotzkie:1964:HBT

[RL64] Robert A. Ragotzkie and Gene E. Likens. The heat balance of two Antarctic lakes. Limnology and Oceanography, 9(3):412–425, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1965:CTD

[RM65] J. H. Ryther and D. W. Menzel. Comparison of the ^{14}C-technique with direct measurement of photosynthetic carbon fixation. Limnology and Oceanography, 10(3):490–492, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Rowe:1968:OTB

[RM68] Gilbert T. Rowe and Robert J. Menzies. Orientation in two bathyal, benthic decapods, Munida valida Smith and Parapagurus pilosimanus Smith. Limnology and Oceanography, 13(3):549–552, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1961:DVS

[RMV61] J. H. Ryther, D. W. Menzel, and R. F. Vaccaro. Diurnal variations in some chemical and biological properties of the Sargasso Sea. Limnology and Oceanography, 6(2):149–153, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Rochna:1963:MDL

[Roc63] A. G. Rochna. Microbiology of detritus of lakes. Limnology and Oceanography, 8(4):388–393, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Rodgers:1961:CWT

[Rod61] G. K. Rodgers. Comment on “Water transport in Lake Huron”. Limnology and Oceanography, 6(4):490–491, April 1961. CODEN LIOCAH. ISSN 0024-3590. See [PA60].

Ropes:1963:IMG

[Rop63] John W. Ropes. The incidence of Malacobdella grossa in hard clams from Nantucket Sound, Massachusetts. Limnology and Oceanography, 8(3):353–355, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Roth:1968:BLD

[Rot68] James C. Roth. Benthic and limnetic distribution of three Chaoborus species in a southern Michigan lake (Diptera, Chaoboridae). Limnology and Oceanography, 13(2):242–249, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Royce:1968:BRR

[Roy68] W. F. Royce. Book review: Ricker, W. E. [Ed.]. 1968. Methods for assessment of fish production in fresh waters. IBP Handbook No. 3. F. A. Davis, Philadelphia, Pennsylvania. 328 p. $6.50. Limnology and Oceanography, 13(4):730–731, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Robertson:1968:DOM

[RPA68] Andrew Robertson, Charles F. Powers, and Robert F. Anderson. Direct observations on Mysis relicta from a submarine. Limnology and Oceanography, 13(4):700–702, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Renfro:1969:IPS

[RPG69] William C. Renfro, Roger V. Phelps, and Thomas F. Guthrie. Influence of preservation and storage on radiozinc and total zinc concentrations in fish. Limnology and Oceanography, 14(1):168–170, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Richards:1965:EDP

[RR65] Norman J. Richards and Sarah W. Richards. Effect of decalcification procedures on the dry weights of benthic invertebrates. Limnology and Oceanography, 10(3):469–471, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Ragotzkie:1967:IRS

[RR67] Robert A. Ragotzkie and Evon P. Ruzecki. An ice record of standing waves. Limnology and Oceanography, 12(2):328–329, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Richman:1969:FCH

[RR69] Sumner Richman and Jane N. Rogers. The feeding of Calanus helgolandicus on synchronously growing populations of the marine diatom Ditylum brightwellii. Limnology and Oceanography, 14(5):701– 709, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Reif:1966:SPS

[RT66] Charles B. Reif and Donald W. Tappa. Selective predation: smelt and cladocerans in Harveys Lake. Limnology and Oceanography, 11 (3):437–438, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Runnells:1965:CMD

[Run65] Donald D. Runnells. A conductometric method for determining the alkalinity of freshwater. Limnology and Oceanography, 10(2):296–298, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Ragotzkie:1961:PCC

[RV61] Robert A. Ragotzkie and Jacob Verduin. Phytoplankton communities, CO_{2} and O_{2} changes. Limnology and Oceanography, 6(3): 367–368, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Riley:1965:OAS

[RVW65] Gordon A. Riley, Denise Van Hemert, and Peter J. Wanngersky. Organic aggregates in surface and deep waters of the Sargasso Sea. Limnology and Oceanography, 10(3):354–363, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Rich:1969:SSU

[RW69] Peter H. Rich and Robert G. Wetzel. A simple, sensitive underwater photometer. Limnology and Oceanography, 14(4):611–613, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Riley:1964:OAT

[RWV64] Gordon A. Riley, Peter J. Wangersky, and Denise Van Hemert. Organic aggregates in tropical and subtropical surface waters of the north Atlantic Ocean. Limnology and Oceanography, 9(4):546–550, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Saila:1961:SWF

[Sai61] Saul B. Saila. A study of winter flounder movements. Limnology and Oceanography, 6(3):292–298, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1960:BSB

[San60] Howard L. Sanders. Benthic studies in Buzzards Bay III. The structure of the soft-bottom community. Limnology and Oceanography, 5 (3):138–153, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Sandercock:1967:SSM

[San67] Gail A. Sandercock. A study of selected mechanisms for the coexistence of Diaptomus spp. in Clarke Lake, Ontario. Limnology and Oceanography, 12(1):97–112, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Shulman:1961:VVW

[SB61a] Mark D. Shulman and Reid A. Bryson. The vertical variation of winddriven currents in Lake Mendota. Limnology and Oceanography, 6(3): 347–355, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Steele:1961:RBP

[SB61b] J. H. Steele and I. E. Baird. Relations between primary production, chlorophyll and particulate carbon. Limnology and Oceanography, 6 (1):68–78, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Steele:1962:CCR

[SB62a] J. H. Steele and I. E. Baird. Carbon-chlorophyll relations in cultures. Limnology and Oceanography, 7(1):101–102, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Steele:1962:FRB

[SB62b] J. H. Steele and I. E. Baird. Further relations between primary production, chlorophyll, and particulate carbon. Limnology and Oceanography, 7(1):42–47, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1965:EOF

[SB65a] Theodore J. Smayda and Brenda J. Boleyn. Experimental observations on the flotation of marine diatoms. I. Thalassiosira cf. nana, Thalassiosira rotula and Nitzschia seriata. Limnology and Oceanography, 10(4):499–509, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Steele:1965:CCP

[SB65b] J. H. Steele and I. E. Baird. The chlorophyll-a content of particulate organic matter in the northern North Sea. Limnology and Oceanography, 10(2):261–267, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1966:EOFa

[SB66a] Theodore J. Smayda and Brenda J. Boleyn. Experimental observations on the flotation of marine diatoms. II. Skeletonema costatum and Rhizosolenia setigera. Limnology and Oceanography, 11(1):18– 34, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1966:EOFb

[SB66b] Theodore J. Smayda and Brenda J. Boleyn. Experimental observations on the flotation of marine diatoms. III. Bacteriastrum hyalinum and Chaetoceros lauderi. Limnology and Oceanography, 11(1):35–43, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Stockner:1967:SDA

[SB67] John G. Stockner and Woodruff W. Benson. The succession of diatom assemblages in the recent sediments of Lake Washington. Limnology and Oceanography, 12(4):513–532, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Steele:1968:PES

[SB68] J. H. Steele and I. E. Baird. Production ecology of a sandy beach. Limnology and Oceanography, 13(1):14–25, April 1968. CODEN LI-OCAH. ISSN 0024-3590.

Sanders:1968:RTS

[SC68a] Herman O. Sanders and Oliver B. Cope. The relative toxicities of several pesticides to naiads of three species of stoneflies. Limnology and Oceanography, 13(1):112–117, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Small:1968:RCP

[SC68b] Lawrence F. Small and Herbert Curl, Jr. The relative contribution of particulate chlorophyll and river tripton to the extinction of light off the coast of Oregon. Limnology and Oceanography, 13(1):84–91, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Schaefer:1962:NS

[Sch62] Milner B. Schaefer. Ntroduction and summary. Limnology and Oceanography, 7(S1):iii–vi, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Schiffman:1965:EMS

[Sch65] Arnold Schiffman. Energy measurements in the swash–surf zone. Limnology and Oceanography, 10(2):255–260, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Scott:1965:PPC

[Sco65] George T. Scott. Physiology and pharmacology of color change in the sand flounder Scopthalamus aquosus. Limnology and Oceanography, 10(S1):R230–R246, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Scrimger:1960:TVE

[Scr60] J. A. Scrimger. Temperature variations in Esquimalt Lagoon — a small landlocked body of water subject to tidal flushing. Limnology and Oceanography, 5(4):414–424, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Strand:1969:FFS

[SCV69] John A. Strand, Joseph T. Cummins, and Burton E. Vaughan. A fast-flow sealed disk filter system for marine aquaria. Limnology and Oceanography, 14(3):444–448, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Sears:1965:BACa

[Sea65a] Mary Sears. Bibliography of Alfred C. Redfield. Limnology and Oceanography, 10(S1):R5–R8, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Sears:1965:BACb

[Sea65b] Mary Sears. Bibliography of Alfred C. Redfield. Limnology and Oceanography, 10(S1):R9–R14, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1967:OFS

[SEP67] R. W. Sheldon, T. P. T. Evelyn, and T. R. Parsons. On the occurrence and formation of small particles in seawater. Limnology and Oceanography, 12(3):367–375, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Sette:1967:BRM

[Set67] Oscar E. Sette. Book review: Miller, Robert C. 1966. The Sea. Random House, New York. 316 p. $15.00. Limnology and Oceanography, 12(2):365, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Seliger:1969:TPR

[SFM69] H. H. Seliger, W. G. Fastie, and W. D. Mcelroy. Towable photometer for rapid area mapping of concentrations of bioluminescent marine dinoflagellates. Limnology and Oceanography, 14(5):806–813, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Sperling:1969:BCT

[SG69] Jon A. Sperling and Ralph Grunewald. Batch culturing of thermophilic benthic algae and phosphorus uptake in a laboratory stream model. Limnology and Oceanography, 14(6):944–949, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1962:SIF

[SGMH62] H. L. Sanders, E. M. Goudsmit, E. L. Mills, and G. E. Hampson. A study of the intertidal fauna of Barnstable Harbor, Massachusetts. Limnology and Oceanography, 7(1):63–79, April 1962. CODEN LI-OCAH. ISSN 0024-3590.

Stross:1960:SLI

[SH60] Raymond G. Stross and A. D. Hasler. Some lime-induced changes in lake metabolism. Limnology and Oceanography, 5(3):265–272, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Small:1967:RVM

[SH67] Lawrence F. Small and James F. Hebard. Respiration of a vertically migrating marine crustacean Euphausia pacifica Hansen. Limnology and Oceanography, 12(2):272–280, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Shapiro:1960:CMM

[Sha60] Joseph Shapiro. The cause of a metalimnetic minimum of dissolved oxygen. Limnology and Oceanography, 5(3):216–227, April 1960. CO-DEN LIOCAH. ISSN 0024-3590.

Shapiro:1966:MFI

[Sha66] Joseph Shapiro. On the measurement of ferrous iron in natural waters. Limnology and Oceanography, 11(2):293–298, April 1966. CO-DEN LIOCAH. ISSN 0024-3590.

Sharp:1969:BGA

[Sha69] Jonathan H. Sharp. Blue-green algae and carbonates — Schizothrix calcicola and algal stromatolites from Bermuda. Limnology and Oceanography, 14(4):568–578, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Sherman:1963:PCD

[She63] Kenneth Sherman. Pontellid copepod distribution in relation to surface water types in the central north Pacific. Limnology and Oceanography, 8(2):214–227, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Sherman:1964:PCO

[She64] Kenneth Sherman. Pontellid copepod occurrence in the central south Pacific. Limnology and Oceanography, 9(4):476–484, April 1964. CO-DEN LIOCAH. ISSN 0024-3590.

Sheldon:1968:SER

[She68] R. W. Sheldon. Sedimentation in the estuary of the River Crouch, Essex, England. Limnology and Oceanography, 13(1):72–83, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Strickland:1969:UDT

[SHHEL69] J. D. H. Strickland, O. Holm-Hansen, R. W. Eppley, and R. J. Linn. The use of a deep tank in plankton ecology. I. Studies of the growth and composition of phytoplankton crops at low nutrient levels. Limnology and Oceanography, 14(1):23–34, January 1969. CODEN LI-OCAH. ISSN 0024-3590.

Shimada:1962:RLL

[Shi62] Bell M. Shimada. Results of long-line fishing. Limnology and Oceanography, 7(S1):xlii–xliii, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Shonting:1964:SOS

[Sho64] David H. Shonting. Some observations of short-term heat transfer through the surface layers of the ocean. Limnology and Oceanography, 9(4):576–588, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Sieburth:1963:SFZ

[Sie63] John McN. Sieburth. A simple form of the ZoBell bacteriological sampler for shallow watery. Limnology and Oceanography, 8(4):489– 492, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Silker:1964:VEC

[Sil64] W. B. Silker. Variations in elemental concentrations in the Columbia River. Limnology and Oceanography, 9(4):540–545, April 1964. CO-DEN LIOCAH. ISSN 0024-3590.

Swinnerton:1968:SGC

[SLC68] J. W. Swinnerton, V. J. Linnenbom, and C. H. Cheek. A sensitive gas chromatographic method for determining carbon monoxide in seawater. Limnology and Oceanography, 13(1):193–195, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Small:1963:EWD

[Sma63] Lawrence F. Small. Effect of wind on the distribution of chlorophylla in Clear Lake, Iowa. Limnology and Oceanography, 8(4):426–432, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1969:SMS

[Sma69] Theodore J. Smayda. Some measurements of the sinking rate of fecal pellets. Limnology and Oceanography, 14(4):621–625, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1965:SFD

[SMH65] H. L. Sanders, P. C. Mangelsdorf, Jr., and G. R. Hampson. Salinity and faunal distribution in the Pocasset River, Massachusetts. Limnology and Oceanography, 10(S1):R216–R229, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Smith:1962:TCC

[Smi62] Stanford H. Smith. Temperature correction in conductivity measurements. Limnology and Oceanography, 7(3):330–334, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Smith:1968:OCN

[Smi68] Raymond C. Smith. The optical characterization of natural waters by means of an “extinction coefficient”. Limnology and Oceanography, 13(3):423–429, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Scott:1969:MLC

[SMSW69] Jon T. Scott, Glenn E. Myer, Ronald Stewart, and Eric G. Walther. On the mechanism of Langmuir circulations and their role in epilimnion mixing. Limnology and Oceanography, 14(4):493–503, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Scarola:1968:FPS

[SN68] John F. Scarola and Anthony J. Novotny. Folsom plankton splitter modified for enumeration of entomostraca. Limnology and Oceanography, 13(1):195–196, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Schoettger:1961:ATF

[SO61] Richard A. Schoettger and John R. Oliver. Accumulation of toxaphene by fish-food organisms. Limnology and Oceanography, 6 (2):216–219, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Soli:1964:SIP

[Sol64] Giorgio Soli. A system for isolating phytoplankton organisms in unialgal and bacteria-free culture. Limnology and Oceanography, 9(2): 265–268, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Soli:1966:BCP

[Sol66] Giorgio Soli. Bioluminescent cycle of photosynthetic dinoflagellates. Limnology and Oceanography, 11(3):355–363, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Solorzano:1969:DAN

[Sol69] Lucia Solórzano. Determination of ammonia in natural waters by the phenolhypochlorite method. Limnology and Oceanography, 14 (5):799–801, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Smith:1960:FYG

[SP60] Lloyd L. Smith, Jr. and Richard L. Pycha. First-year growth of the walleye, Stizostedion vitreum vitreum (Mitchill), and associated factors in the Red Lakes, Minnesota. Limnology and Oceanography, 5(3):281–290, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Stone:1969:BSS

[SP69] Julian Stone and T. E. Pochapsky. Brillouin scattering by seawater. Limnology and Oceanography, 14(5):783–787, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Sparks:1966:BRG

[Spa66] Albert K. Sparks. Book review: Galtsoff, P. S. 1964. The American Oyster Crassostrea virginica Gmelin. Fishery Bulletin, v. 64. United States Government Printing Office, Washington, D. C. iii + 480 p, $2.75. Limnology and Oceanography, 11(2):312, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Smith:1965:DRS

[SPCF65] Raymond C. Smith, K. C. Pillai, Tsaihwa J. Chow, and Theodore R. Folsom. Determination of rubidium in seawater. Limnology and Oceanography, 10(2):226–232, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Spilhaus:1968:OLS

[Spi68] A. F. Spilhaus, Jr. Observations of light scattering in seawater. Limnology and Oceanography, 13(3):418–422, July 1968. CODEN LI-OCAH. ISSN 0024-3590.

Squire:1969:OCB

[Squ69] James L. Squire, Jr. Observations on cumulative bottom drift in Monterey Bay using seabed drifters. Limnology and Oceanography, 14(1):163–167, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Stefansson:1963:PCN

[SR63] Unnsteinn Stefánsson and Francis A. Richards. Processes contributing to the nutrient distribution off the Columbia River and Strait of Juan de Fuca. Limnology and Oceanography, 8(4):394–410, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Sreenivasan:1964:LSF

[Sre64a] A. Sreenivasan. Limnological studies and fish yield in three upland lakes of Madras State, India. Limnology and Oceanography, 9(4): 564–575, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Sreenivasan:1964:LPP

[Sre64b] A. Sreenivasan. The limnology, primary production, and fish production in a tropical pond. Limnology and Oceanography, 9(3):391–396, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Stephens:1961:UAA

[SS61] Grover C. Stephens and Robert A. Schinske. Uptake of amino acids by marine invertebrates. Limnology and Oceanography, 6(2):175–181, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Schurr:1962:MAS

[SS62a] J. M. Schurr and M. N. Stamper. Model for the accumulation of strontium and calcium by recently molted crayfish (Cambarus longulus longerostris Ort.). Limnology and Oceanography, 7(4):474–477, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Stephens:1962:UTR

[SS62b] K. Stephens and J. D. H. Strickland. Use of a thermopile radiometer for measuring the attenuation of photosynthetically active radiation in the sea. Limnology and Oceanography, 7(4):485–487, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Sladecek:1963:RBW

[SS63] Vladimír Sládeček and Alena Sládečková. Relationship between wet weight and dry weight of the periphyton. Limnology and Oceanography, 8(2):309–311, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Sherman:1968:ODB

[SS68a] Kenneth Sherman and Everett G. Schaner. Observations on the distribution and breeding of Sagitta elegans (Chaetognatha) in coastal waters of the Gulf of Maine. Limnology and Oceanography, 13(4): 618–625, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Solorzano:1968:PS

[SS68b] L. Solórzano and J. D. H. Strickland. Polyphosphate in seawate. Limnology and Oceanography, 13(3):515–518, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Sutcliffe:1968:MDA

[SS68c] William H. Sutcliffe, Jr. and Jonathan Sharp. Measurement of deoxyribonucleic acid in the ocean and its ecological significance. Limnology and Oceanography, 13(3):507–514, July 1968. CODEN LI-OCAH. ISSN 0024-3590.

Schubel:1969:MDV

[SS69a] J. R. Schubel and E. W. Schiemer. On the microscopic determination of the volume size distribution of fine-grained suspended particles. Limnology and Oceanography, 14(3):438–441, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1969:RMP

[SS69b] R. W. Sheldon and W. H. Sutcliffe, Jr. Retention of marine particles by screens and filters. Limnology and Oceanography, 14(3):441–444, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Sanford:1969:SFM

[SSG69] Gary R. Sanford, Anne Sands, and Charles R. Goldman. A settlefreeze method for concentrating phytoplankton in quantitative studies. Limnology and Oceanography, 14(5):790–794, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Seki:1968:ODM

[SSP68] H. Seki, J. Skelding, and T. R. Parsons. Observations on the decomposition of a marine sediment. Limnology and Oceanography, 13(3): 440–447, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Strickland:1961:SSM

[ST61] J. D. H. Strickland and L. D. B. Terhune. The study of in-situ marine photosynthesis using a large plastic bag. Limnology and Oceanography, 6(1):93–96, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Salsman:1965:OSD

[ST65] G. G. Salsman and W. H. Tolbert. Observations on the sand dollar, Mellita quinquiesperforata. Limnology and Oceanography, 10(1):152– 155, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Stahl:1966:ECM

[Sta66] John B. Stahl. The ecology of Chaoborus in Myers Lake, Indiana. Limnology and Oceanography, 11(2):177–183, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Steele:1962:ECP

[Ste62a] J. H. Steele. Environmental control of photosynthesis in the sea. Limnology and Oceanography, 7(2):137–150, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Stephens:1962:ITM

[Ste62b] K. Stephens. Improved tripping mechanism for plastic water samplers. Limnology and Oceanography, 7(4):484, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Stephens:1963:DLP

[Ste63] K. Stephens. Determination of low phosphate concentrations in lake and marine waters. Limnology and Oceanography, 8(3):361–362, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Stewart:1964:OEL

[Ste64] Ronald Stewart. Onthe estimation of lake depth from theperiod of the seiche. Limnology and Oceanography, 9(4):606–607, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Smith:1960:CCA

[STH60] James B. Smith, Mitsunobu Tatsumoto, and Donald W. Hood. Carbamino carboxylic acids in photosynthesis. Limnology and Oceanography, 5(4):425–431, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Strohal:1969:ICM

[STK69] Petar Strohal, Josip Tuta, and Zvonimir Kolar. Investigations of certain microconstituents in two tunicates. Limnology and Oceanography, 14(2):265–268, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Stockner:1967:OTA

[Sto67] John G. Stockner. Observations of thermophilic algal communities in Mount Rainier and Yellowstone National Parks. Limnology and Oceanography, 12(1):13–17, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Straskraba:1964:PRN

[Str64] Milan Straškraba. Preliminary results of a new method for the quantitative sorting of freshwater net plankton into main groups. Limnology and Oceanography, 9(2):268–270, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Strathmann:1967:EOC

[Str67] Richard R. Strathmann. Estimating the organic carbon content of phytoplankton from cell volume or plasma volume. Limnology and Oceanography, 12(3):411–418, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Strickland:1968:CPN

[Str68] J. D. H. Strickland. A comparison of profiles of nutrient and chlorophyll concentrations taken from discrete depths and by continuous recording. Limnology and Oceanography, 13(2):388–391, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Sturges:1969:BRT

[Stu69] Wilton Sturges. Book review: Tsuchiya, M. 1968. Upper waters of the intertropical Pacific Ocean. The Johns Hopkins Oceanographic Studies, No. 4. The Johns Hopkins Press, Baltimore, Md. 50 p. + 4 separate Figs. $8.50. Limnology and Oceanography, 14(4):650, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Stevenson:1966:IMN

[SU66] Robert A. Stevenson and Sara Lugo Ufret. Iron, manganese, and nickel in skeletons and food of the sea urchins tripneustes esculentus and echinometra lucunter. Limnology and Oceanography, 11(1):11– 17, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Sugden:1967:TWL

[Sug67] Lawson G. Sugden. A technique for weighing live aquatic invertebrates. Limnology and Oceanography, 12(4):557, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Sutcliffe:1965:GER

[Sut65] W. H. Sutcliffe, Jr. Growth estimates from ribonucleic acid content in some small organisms. Limnology and Oceanography, 10(S1):R253– R258, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Saur:1965:NCA

[SV65] J. F. T. Saur and J. Verduin. Notice from the Committee on Apparatus and Supplies. Limnology and Oceanography, 10(2):305–306, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Sherman:1961:ICP

[SW61] Kenneth Sherman and John P. Wise. Incidence of the cod parasite Lernaeocera branchialis L. in the New England area, and its possible use as an indicator of cod populations. Limnology and Oceanography, 6(1):61–67, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Swanson:1965:APS

[Swa65] George A. Swanson. Automatic plankton sampling system. Limnology and Oceanography, 10(1):149–152, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Saz:1963:AAM

[SWBL63] Arthur K. Saz, Stanley Watson, Sarah R. Brown, and Dolores L. Lowery. Antimicrobial activity of marine waters.1 I. macromolecular nature of antistapi-iylococcal factor. Limnology and Oceanography, 8(1):63–67, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Taft:1960:SSE

[Taf60] Bruce A. Taft. A statistical study of the estimation of abundance of sardine (Sardinops caerulea) eggs. Limnology and Oceanography, 5 (3):245–264, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Tailing:1960:CLF

[Tai60] J. F. Tailing. Comparative laboratory and field studies of photosynthesis by a marine planktonic diatom. Limnology and Oceanography, 5(1):62–77, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Talling:1963:OSA

[Tal63] J. F. Talling. Origin of stratification in an African Rift lake. Limnology and Oceanography, 8(1):68–78, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Taub:1969:BMF

[Tau69] Frieda B. Taub. A biological model of a freshwater community: a gnotobiotic ecosystem. Limnology and Oceanography, 14(1):136–142, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Takahashi:1968:CIB

[TBLT68] Taro Takahashi, Wallace Broecker, Yuan Hui Li, and David Thurber. Chemical and isotopic balances for a meromictic lake. Limnology and Oceanography, 13(2):272–292, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Teal:1967:REO

[TC67] J. M. Teal and F. G. Carey. Respiration of a euphausiid from the oxygen minimum layer. Limnology and Oceanography, 12(4):548–550, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Taub:1964:CDF

[TD64] Frieda B. Taub and Alexander M. Dollar. A Chlorella–Daphnia foodchain study: the design of a compatible chemically defined culture medium. Limnology and Oceanography, 9(1):61–74, April 1964. CO-DEN LIOCAH. ISSN 0024-3590.

Tubb:1965:HIP

[TD65] Richard A. Tubb and Troy C. Dorris. Herbivorous insect populations in oil refinery effluent holding pond series. Limnology and Oceanography, 10(1):121–134, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1967:POR

[TD67] Robert W. Thomas and Stephen W. Dorey. Protected oceanographic reversing thermometer comparison study. Limnology and Oceanography, 12(2):361–363, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Taub:1968:NIC

[TD68] Frieda B. Taub and Alexander M. Dollar. The nutritional inadequacy of Chlorella and Chlamydomonas as food for Daphnia pulex. Limnology and Oceanography, 13(4):607–617, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Teal:1960:TSN

[Tea60] John M. Teal. A technique for separating nematodes and small arthropods from marine muds. Limnology and Oceanography, 5(3): 341–342, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Takahashi:1968:VDO

[TeI68] Masayuki Takahashi and Shun ei Ichimura. Vertical distribution and organic matter production of photosynthetic sulfur bacteria in Japanese lakes. Limnology and Oceanography, 13(4):644–655, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Truper:1968:CPS

[TG68] Hans G. Trüper and Sebastiano Genovese. Characterization of photosynthetic sulfur bacteria causing red water in Lake Faro (Messina, Sicily). Limnology and Oceanography, 13(2):225–232, April 1968. CO-DEN LIOCAH. ISSN 0024-3590.

Tyler:1962:IMF

[TH62] John E. Tyler and Roger Howerton. Instrument for measuring the forward scattering coefficient of sea water. Limnology and Oceanography, 7(3):393–395, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Tashiro:1969:MPV

[TH69] Joseph E. Tashiro and James F. Hebard. A modification of the plankton volume gauge for use aboard ship. Limnology and Oceanography, 14(5):794–796, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Turekian:1967:VSC

[THJ67] Karl K. Turekian, Robert C. Harris, and David G. Johnson. The variations of Si, Cl, Na, Ca, Sr, Ba, Co, and Ag in the Neuse River, North Carolina. Limnology and Oceanography, 12(4):702–706, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1966:SNN

[Tho66] William H. Thomas. Surface nitrogenous nutrients and phytoplankton in the northeastern tropical Pacific Ocean. Limnology and Oceanography, 11(3):393–400, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Tibbs:1965:OGC

[Tib65] John F. Tibbs. Observations of Gigantocypris (Crustacea: Ostracoda) in the Antarctic Ocean. Limnology and Oceanography, 10(3): 480–482, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Tietjen:1968:CPP

[Tie68] John H. Tietjen. Chlorophyll and pheo-pigments in estuarine sediments. Limnology and Oceanography, 13(1):189–192, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Teal:1961:GEG

[TK61] John M. Teal and John Kanwisher. Gas exchange in a Georgia salt marsh. Limnology and Oceanography, 6(4):388–399, April 1961. CO-DEN LIOCAH. ISSN 0024-3590.

Thomas:1962:UMD

[TLH62] William H. Thomas, Donald W. Lear, Jr., and Francis T. Haxo. Uptake by the marine dinoflagellate, Gonyaulax polyedra, of radioactivity formed during an underwater nuclear test. Limnology and Oceanography, 7(S1):lxvi–lxxi, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Tindall:1964:IAT

[TM64] Donald R. Tindall and W. L. Minckley. An integrated application of three kinds of sampling techniques to stream limnology. Limnology and Oceanography, 9(2):270–272, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1966:MMP

[TO66] N. A. Thomas and R. L. O’Connell. A method for measuring primary production by stream benthos. Limnology and Oceanography, 11(3): 386–392, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Topping:1969:CCZ

[Top69] G. Topping. Concentration of cobalt and zinc from seawater using 5,7-dibromo-8-hydroxyquinoline (bromo-oxine) supported on an anion exchange resin. Limnology and Oceanography, 14(5):798–799, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Turner:1965:SCF

[TP65] Harry J. Turner, Jr. and Bryce Prindle. Some characteristics of “fishbite” damage on deep-sea mooring lines. Limnology and Oceanography, 10(S1):R259–R264, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Tolbert:1964:UC

[TPS64] W. H. Tolbert, R. H. Payne, and G. G. Salsman. An underwater crane. Limnology and Oceanography, 9(1):150–151, April 1964. CO-DEN LIOCAH. ISSN 0024-3590.

Thayer:1969:BSR

[TR69] Orla E. Thayer and Richard G. Redmond. Budget salinity recorder. Limnology and Oceanography, 14(4):641–643, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Traganza:1967:CCA

[TS67] Eugene D. Traganza and Barney J. Szabo. Calculation of calcium anomalies on the Great Bahama Bank from alkalinity and chlorinity data. Limnology and Oceanography, 12(2):281–286, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1963:TGI

[TSB63] William H. Thomas, Harold L. Scotten, and John S. Bradshaw. Thermal gradient incubators for small aquatic organisms. Limnology and Oceanography, 8(3):357–360, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Teraguchi:1966:VDM

[TT66] M. Teraguchi and M. Teraguchi. Vertical distribution and migration of Chaoborus flavicans larvae in Corbett Lake, British Columbia. Limnology and Oceanography, 11(2):164–176, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Tracy:1969:FDA

[TV69] Susan F. Tracy and J. R. Vallentyne. Fungal decomposition and amino acid analysis of Mysis relicta Lovén. Limnology and Oceanography, 14(3):352–356, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Taysi:1964:OPD

[TvU64] I. Taysi and N. van Uden. Occurrence and population densities of yeast species in an estuarine–marine area. Limnology and Oceanography, 9(1):42–45, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Teal:1966:DEN

[TW66] John M. Teal and Wolfgang Wieser. The distribution and ecology of nematodes in a Georgia salt marsh. Limnology and Oceanography, 11(2):217–222, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1961:SAE

[Tyl61a] J. E. Tyler. Sun-altitude effect on the distribution of underwater light. Limnology and Oceanography, 6(1):24–25, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1961:SPD

[Tyl61b] John E. Tyler. Scattering properties of distilled and natural waters. Limnology and Oceanography, 6(4):451–456, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1968:BRJ

[Tyl68a] John Tyler. Book review: Jerlov, N. G. 1968. Optical oceanography. American Elsevier Publ. Co., Inc., New York. 194 p. $13.50. Limnology and Oceanography, 13(4):731–732, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1968:SD

[Tyl68b] John E. Tyler. The secchi disc. Limnology and Oceanography, 13(1): 1–6, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Ukeles:1965:SMM

[Uke65] Ravenna Ukeles. A simple method for the mass culture of marine algae. Limnology and Oceanography, 10(3):492–495, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

Ukeles:1969:IDT

[US69] Ravenna Ukeles and Beatrice M. Sweeney. Influence of dinoflagellate trichocysts and other factors on the feeding of Crassostrea virginica larvae on Monochrysis lutheri. Limnology and Oceanography, 14(3): 403–410, May 1969. CODEN LIOCAH. ISSN 0024-3590.

vonArx:1965:ADT

[vA65] William S. von Arx. Absolute dynamic topography. Limnology and Oceanography, 10(S1):R265–R273, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Vaccaro:1969:RNM

[Vac69] Ralph F. Vaccaro. The response of natural microbial populations in seawater to organic enrichment. Limnology and Oceanography, 14(5): 726–735, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1963:IPS

[Val63] J. R. Vallentyne. Isolation of pyrite spherules from recent sediments. Limnology and Oceanography, 8(1):16–30, April 1963. CODEN LI-OCAH. ISSN 0024-3590.

Valentine:1966:NAM

[Val66] James W. Valentine. Numerical analysis of marine molluscan ranges on the extratropical northeastern Pacific shelf. Limnology and Oceanography, 11(2):198–211, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1967:BRC

[Val67] J. R. Vallentyne. Book review: Cowgill, U. M., C. E. Goulden, G. E. Hutchinson, R. Patrick, A. A. Racek, and M. Tsukada. 1966. The History of Laguna de Petenxil. Mem. Conn. Acad. Arts and Sci., v.

27. 126 p. Limnology and Oceanography, 12(3):182–183, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1969:DL

[Val69] J. R. Vallentyne. Definition of a limnologist. Limnology and Oceanography, 14(5):815, September 1969. CODEN LIOCAH. ISSN 0024-3590.

VanLandingham:1960:NSS

[Van60] John W. Van Landingham. A note on a stabilized starch indicator for use in iodometric and iodimetric determinations. Limnology and Oceanography, 5(3):343–345, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Verduin:1960:DTS

[VBO60] Jacob Verduin, Robert J. Beyers, and Howard T. Odum. Differential titration with strong acids or bases vs. CO_{2} water for productivity studies. Limnology and Oceanography, 5(3):228–230, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Verduin:1960:PCW

[Ver60] Jacob Verduin. Phytoplankton communities of western Lake Erie and the CO_{2} and O_{2} changes associated with them. Limnology and Oceanography, 5(4):372–380, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Verduin:1961:CPT

[Ver61] Jacob Verduin. Changes in pH and total CO_{2} in natural waters. Limnology and Oceanography, 6(1):80–83, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Verduin:1962:ZUU

[Ver62] Jacob Verduin. Zn^{65} uptake by Ulva lactuca. Limnology and Oceanography, 7(2):270–271, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Verber:1964:ICS

[Ver64] James L. Verber. Initial current studies in Lake Michigan. Limnology and Oceanography, 9(3):426–430, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Vaccaro:1968:ORG

[VHJC68] Ralph F. Vaccaro, Sonja E. Hicks, Holger W. Jannasch, and Francis G. Carey. The occurrence and role of glucose in seawater. Limnology and Oceanography, 13(2):356–360, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Vincent:1967:CRI

[Vin67] Ernest Richard Vincent. A comparison of riffle insect populations in the Gibbon River above and below the geyser basins, Yellowstone National Park. Limnology and Oceanography, 12(1):18–25, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Vishniac:1960:SRM

[Vis60] Helen S. Vishniac. Salt requirements of marine phycomycetes. Limnology and Oceanography, 5(4):362–366, April 1960. CODEN LI-OCAH. ISSN 0024-3590.

Vishniac:1961:BAT

[Vis61] Helen S. Vishniac. A biological assay for thiamine in sea water. Limnology and Oceanography, 6(1):31–35, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Vaccaro:1966:SHA

[VJ66] Ralph F. Vaccaro and Holger W. Jannasch. Studies on heterotrophic activity in seawater based on glucose assimilation. Limnology and Oceanography, 11(4):596–607, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Vaccaro:1967:VUK

[VJ67] Ralph F. Vaccaro and Holger W. Jannasch. Variations in uptake kinetics for glucose by natural populations in seawater. Limnology and Oceanography, 12(4):540–542, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Vishniac:1961:CTL

[VR61] H. S. Vishniac and G. A. Riley. Cobalamin and thiamine in Long Island Sound: patterns of distribution and ecological significance. Limnology and Oceanography, 6(1):36–41, April 1961. CODEN LI-OCAH. ISSN 0024-3590.

Veeh:1968:CSU

[VT68] H. Herbert Veeh and Karl K. Turekian. Cobalt, silver, and uranium concentrations of reef-building corals in the Pacific Ocean. Limnology and Oceanography, 13(2):304–308, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Vaccaro:1962:NMP

[VTK62] Ralph F. Vaccaro, Jon C. Thunberg, and Bostwick H. Ketchum. Nitrate measurements in the presence of high nitrite concentrations.

Limnology and Oceanography, 7(3):322–329, April 1962. CODEN LI-OCAH. ISSN 0024-3590.

VanCleve:1966:NDS

[VTK66] Richard Van Cleve, Robert Y. Ting, and Joseph C. Kent. A new device for sampling marine demersal animals for ecological study. Limnology and Oceanography, 11(3):438–443, April 1966. CODEN LIOCAH. ISSN 0024-3590.

vanUden:1963:DPD

[vUB63] N. van Uden and R. Castelo Branco. Distribution and population densities of yeast species in Pacific water, air, animals, and kelp off southern California. Limnology and Oceanography, 8(3):323–329, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1965:SDCa

[Wal65a] Gerald E. Walsh. Studies on dissolved carbohydrate in Cape Cod waters. I. General survey. Limnology and Oceanography, 10(4):570– 576, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1965:SDCb

[Wal65b] Gerald E. Walsh. Studies on dissolved carbohydrate in Cape Cod waters. II. Diurnal fluctuation in Oyster Pond. Limnology and Oceanography, 10(4):577–582, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1966:SDC

[Wal66] Gerald E. Walsh. Studies on dissolved carbohydrate in Cape Cod waters. III. Seasonal variation in Oyster Pond and Wequaquet Lake, Massachusetts. Limnology and Oceanography, 11(2):249–256, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Walker:1967:DOP

[Wal67] Bryan Walker. A diver-operated pneumatic core sampler. Limnology and Oceanography, 12(1):144–146, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1969:VDA

[Wal69] John J. Walsh. Vertical distribution of Antarctic phytoplankton. II. A comparison of phytoplankton standing crops in the Southern Ocean with that of the Florida Strait. Limnology and Oceanography, 14(1): 86–94, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Wangersky:1969:DSC

[Wan69] Peter J. Wangersky. Distribution of suspended carbonate with depth in the ocean. Limnology and Oceanography, 14(6):929–933, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Waters:1961:NCM

[Wat61] Thomas F. Waters. Notes on the chlorophyll method of estimating the photosynthetic capacity of stream periphyton. Limnology and Oceanography, 6(4):486–488, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Watson:1965:CMN

[Wat65a] Stanley W. Watson. Characteristics of a marine nitrifying bacterium, Nitrosocystis oceanus sp. N. Limnology and Oceanography, 10(S1): R274–R289, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Watt:1965:CAL

[Wat65b] W. D. Watt. A convenient apparatus for in situ primary production studies. Limnology and Oceanography, 10(2):298–300, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Waters:1969:SDL

[Wat69] Thomas F. Waters. Subsampler for dividing large samples of stream invertebrate drift. Limnology and Oceanography, 14(5):813–815, September 1969. CODEN LIOCAH. ISSN 0024-3590.

Williams:1962:MRN

[WC62] Peter M. Williams and Arthur R. Coote. Mullin and Riley nitrate determination in the presence of hydrogen sulfide. Limnology and Oceanography, 7(2):258, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1966:VDD

[WD66] Gerald E. Walsh and Jane Douglass. Vertical distribution of dissolved carbohydrate in the Sargasso Sea off Bermuda. Limnology and Oceanography, 11(3):406–408, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Watson:1963:IDA

[WDH63] D. G. Watson, J. J. Davis, and W. C. Hanson. Interspecies differences in accumulation of gamma emitters by marine organisms near the Columbia River mouth. Limnology and Oceanography, 8(2):305–309, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Whitford:1964:ASA

[WDS64] L. A. Whitford, Gary E. Dillard, and George J. Schumacher. An artificial stream apparatus for the study of lotic orcanisms. Limnology and Oceanography, 9(4):598–600, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Wang:1969:VSD

[WE69] Wun-Cheng Wang and Ralph L. Evans. Variation of silica and diatoms in a stream. Limnology and Oceanography, 14(6):941–944, November 1969. CODEN LIOCAH. ISSN 0024-3590.

Webster:1961:ULA

[Web61] Dwight A. Webster. An unusual lake of the Adirondack Mountains, New York. Limnology and Oceanography, 6(1):88–90, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Weise:1961:EUG

[Wei61] John G. Weise. The ecology of Urnatella gracilis Leidy: Phylum Endoprocta. Limnology and Oceanography, 6(2):228–230, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Weibezahn:1965:IWS

[Wei65] Franz H. Weibezahn. An improved water sampling device for small shallow pools. Limnology and Oceanography, 10(2):301–302, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Welch:1968:UMD

[Wel68a] Harold E. Welch. Use of modified diurnal curves for the measurement of metabolism in standing water. Limnology and Oceanography, 13 (4):679–687, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Wells:1968:DDP

[Wel68b] Larue Wells. Daytime distribution of Pontoporeia affinis off bottom in Lake Michigan. Limnology and Oceanography, 13(4):703–705, October 1968. CODEN LIOCAH. ISSN 0024-3590.

Weyl:1963:ECC

[Wey63] Peter K. Weyl. Environment of calcium carbonate deposition west of Andros Island, Bahamas. Limnology and Oceanography, 8(4):494, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Weyl:1964:CEC

[Wey64] Peter K. Weyl. On the change in electrical conductance of seawater with temperature. Limnology and Oceanography, 9(1):75–78, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Weyl:1965:OSD

[Wey65] Peter K. Weyl. On the oxygen supply of the deep Pacific Ocean. Limnology and Oceanography, 10(2):215–219, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Wells:1960:SUN

[WG60] Harry W. Wells and I. E. Gray. Summer upwelling off the northeast coast of North Carolina. Limnology and Oceanography, 5(1):108–109, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Wangersky:1965:PCO

[WG65] Peter J. Wangersky and Donald C. Gordon, Jr. Particulate carbonate, organic carbon, and Mn^{+}+ in the open ocean. Limnology and Oceanography, 10(4):544–550, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Watt:1963:TSP

[WH63a] W. D. Watt and F. R. Hayes. Tracer study of the phosphorus cycle in sea water. Limnology and Oceanography, 8(2):276–285, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Watts:1963:URC

[WH63b] J. R. Watts and R. S. Harvey. Uptake and retention of Cs^{137} by a blue-green alga in continuous flow and batch culture systems. Limnology and Oceanography, 8(1):45–49, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Wright:1965:UOS

[WH65] Richard T. Wright and John E. Hobbie. The uptake of organic solutes in lake water. Limnology and Oceanography, 10(1):22–28, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Wright:1967:HBS

[WH67] John C. Wright and Ross M. Horrall. Heat budget studies on the Madison River, Yellowstone National Park. Limnology and Oceanography, 12(4):578–583, April 1967. CODEN LIOCAH. ISSN 0024-

Wheeler:1967:CDD

[Whe67a] E. H. Wheeler, Jr. Copepod detritus in the deep sea. Limnology and Oceanography, 12(4):697–702, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Whetten:1967:LCW

[Whe67b] John T. Whetten. Lake Chelan, Washington: bottom and subbottom topography. Limnology and Oceanography, 12(2):253–259, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Whitney:1960:PCI

[Whi60] G. G. Whitney, Jr. Procedure for and comments on ice point tests of deep sea reversing thermometers. Limnology and Oceanography, 5 (3):232–235, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Whitley:1962:NBS

[Whi62] L. Stephen Whitley. New bottom sampler for use in shallow streams. Limnology and Oceanography, 7(2):265–266, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1969:EOP

[Whi69] M. Whitfield. Eh as an operational parameter in estuarine studies. Limnology and Oceanography, 14(4):547–558, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Wieser:1960:BSB

[Wie60a] Wolfgang Wieser. Benthic studies in Buzzards Bay II. The meiofauna. Limnology and Oceanography, 5(2):121–137, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Wieser:1960:E

[Wie60b] Wolfgang Wieser. Erratum. Limnology and Oceanography, 5(2):348, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Williams:1960:UCC

[Wil60] Louis G. Williams. Uptake of cesium^{137} by cells and detritus of Euglena and Chlorella. Limnology and Oceanography, 5(3):301–311, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Wilson:1961:MOC

[Wil61] Ronald F. Wilson. Measurement of organic carbon in sea water. Limnology and Oceanography, 6(3):259–261, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Willliams:1963:UNC

[Wil63] Richard B. Willliams. Use of netting to collect motile benthic algae. Limnology and Oceanography, 8(3):360–361, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Williams:1966:DPR

[Wil66] Louis G. Williams. Dominant planktonic rotifers of major waterways of the United States. Limnology and Oceanography, 11(1):83– 91, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Williams:1968:BDP

[Wil68] George C. Williams. Bathymetric distribution of planktonic fish eggs in Long Island Sound. Limnology and Oceanography, 13(2):382–385, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Williams:1969:WOO

[Wil69a] P. J. Leb. Williams. The wet oxidation of organic matter in seawater. Limnology and Oceanography, 14(2):292–297, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Williams:1969:ACD

[Wil69b] P. M. Williams. The association of copper with dissolved organic matter in seawater. Limnology and Oceanography, 14(1):156–158, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Williams:1969:DDO

[Wil69c] P. M. Williams. The determination of dissolved organic carbon in seawater: a comparison of two methods. Limnology and Oceanography, 14(2):297–298, March 1969. CODEN LIOCAH. ISSN 0024-3590.

Wisner:1962:MTS

[Wis62] Robert L. Wisner. Midwater trawl surveys. Limnology and Oceanography, 7(S1):xxxix–xli, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Webb:1967:SRD

[WJ67] Kenneth L. Webb and R. E. Johannes. Studies of the release of dissolved free amino acids by marine zooplankton. Limnology and Oceanography, 12(3):376–382, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Wieser:1961:EPS

[WK61] Wolfgang Wieser and John Kanwisher. Ecological and physiological studies on marine nematodes from a small salt marsh near Woods Hole, Massachusetts. Limnology and Oceanography, 6(3):262–270, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Whitehouse:1966:SBS

[WL66] John W. Whitehouse and Brian G. Lewis. The separation of benthos from stream samples by flotation with carbon tetrachloride. Limnology and Oceanography, 11(1):124–126, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Wilkniss:1968:DFS

[WL68] P. K. Wilkniss and V. J. Linnenbom. The determination of fluorine in seawater by photon activation analysis. Limnology and Oceanography, 13(3):530–533, July 1968. CODEN LIOCAH. ISSN 0024-3590.

Whitworth:1969:ETC

[WL69] Walter R. Whitworth and Thomas H. Lane. Effects of toxicants on community metabolism in pools. Limnology and Oceanography, 14 (1):53–58, January 1969. CODEN LIOCAH. ISSN 0024-3590.

Wooster:1969:RS

[WLD69] Warren S. Wooster, Arthur J. Lee, and Gunter Dietrich. Redefinition of salinity. Limnology and Oceanography, 14(3):437–438, May 1969. CODEN LIOCAH. ISSN 0024-3590.

Wigley:1964:SQC

[WM64] Roland L. Wigley and A. D. McIntyre. Some quantitative comparisons of offshore meiobenthos and macrobenthos south of Martha’s Vineyard. Limnology and Oceanography, 9(4):485–493, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Williams:1966:PPC

[WM66] Richard B. Williams and Marianne B. Murdoch. Phytoplankton production and chlorophyll concentration in the Beaufort Channel, North Carolina. Limnology and Oceanography, 11(1):73–82, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Weber:1967:PSD

[WM67a] Cornelius I. Weber and Donald R. Moore. Phytoplankton, seston, and dissolved organic carbon in the Little Miami River at Cincinnati, Ohio. Limnology and Oceanography, 12(2):311–318, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Wright:1967:PSM

[WM67b] John C. Wright and I. K. Mills. Productivity studies on the Madison River, Yellowstone National Park. Limnology and Oceanography, 12 (4):568–577, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Wolfson:1965:OCT

[Wol65] Fay H. Wolfson. The optical comparator as a tool in plankton research. Limnology and Oceanography, 10(1):156–157, April 1965. CO-DEN LIOCAH. ISSN 0024-3590.

Wooster:1961:YCP

[Woo61] Warren S. Wooster. Yearly changes in the Peru Current. Limnology and Oceanography, 6(2):222–226, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Wood:1962:MPS

[Woo62] E. J. Ferguson Wood. A method for phytoplankton study. Limnology and Oceanography, 7(1):32–35, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Wood:1963:PGM

[Woo63] E. J. Ferguson Wood. Professor L. G. M. Baas Becking. Limnology and Oceanography, 8(2):312, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Wood:1965:CCS

[Woo65a] Langley Wood. A controlled conditions system (CCS) for continuously flowing seawater. Limnology and Oceanography, 10(3):475–477, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Woodcock:1965:MPI

[Woo65b] Alfred H. Woodcock. Melt patterns in ice over shallow waters. Limnology and Oceanography, 10(S1):R290–R297, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Wright:1960:LCF

[Wri60] John C. Wright. The limnology of Canyon Ferry Reservoir: III. Some observations on the density dependence of photosynthesis and its cause. Limnology and Oceanography, 5(4):356–361, April 1960. CODEN LIOCAH. ISSN 0024-3590.

Wright:1961:LCF

[Wri61] John C. Wright. The limnology of Canyon Ferry Reservoir IV. The estimation of primary production from physical limnological data. Limnology and Oceanography, 6(3):330–337, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Wright:1964:DPC

[Wri64] Richard T. Wright. Dynamics of a phytoplankton community in an ice-covered lake. Limnology and Oceanography, 9(2):163–178, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Wright:1965:PDP

[Wri65] John C. Wright. The population dynamics and production of Daphnia in Canyon Ferry Reservoir, Montana. Limnology and Oceanography, 10(4):583–590, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Whitford:1961:ECM

[WS61] L. A. Whitford and G. J. Schumacher. Effect of current on mineral uptake and respiration by a fresh-water alga. Limnology and Oceanography, 6(4):423–425, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Williams:1962:PDM

[WS62] Louis G. Williams and Carol Scott. Principal diatoms of major waterways of the United States. Limnology and Oceanography, 7(3): 363–379, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Williams:1966:CBA

[WS66] P. M. Williams and P. M. Strack. Complexes of boric acid with organic cis-diols in seawater. Limnology and Oceanography, 11(3): 401–404, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Watabe:1966:ETG

[WW66] Norimitsu Watabe and Karl M. Wilbur. Effects of temperature on growth, calcification, and coccolith form in Coccolithus huxleyi (Coccolithineae). Limnology and Oceanography, 11(4):567–575, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Werner:1967:SGB

[WW67] A. E. Werner and M. Waldichuk. A sampler for gases in bottom sediments. Limnology and Oceanography, 12(3):158–161, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Wells:1968:RVR

[WW68a] J. M. Wells and J. E. Warinner. A recording volumetric respirometer for aquatic animals. Limnology and Oceanography, 13(2):376–378, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Wiebe:1968:PPE

[WW68b] Peter Howard Wiebe and Peter Howard Wiebe. Plankton patchiness: effects on repeated net tows. Limnology and Oceanography, 13(2): 315–321, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Wyrtki:1965:ASV

[Wyr65] Klaus Wyrtki. The annual and semiannual variation of sea surface temperature in the north Pacific Ocean. Limnology and Oceanography, 10(3):307–313, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Wyrtki:1966:BRR

[Wyr66] Klaus Wyrtki. Book review: Reid, J. L., Jr. 1965. Intermediate Waters of the Pacific Ocean. Johns Hopkins Oceanographic Studies, No. 2. Johns Hopkins Press, Baltimore, Maryland 21218. 85 p. $8.50. Limnology and Oceanography, 11(2):313–316, April 1966. CODEN LIOCAH. ISSN 0024-3590.

Wyrtki:1967:BRK

[Wyr67] Klaus Wyrtki. Book review: Krauss, W. 1966. Methoden und Ergebnisse der Theoretischen Ozeanographie, Band II: Interne Wellen. Gebrüder Borntraeger, Berlin. 248 p. $45.00. Limnology and Oceanography, 12(4):724, April 1967. CODEN LIOCAH. ISSN 0024-3590.

Yaalon:1964:CCR

[Yaa64] Dan H. Yaalon. Chemical changes in rain-fed marsh waters during the dry season. Limnology and Oceanography, 9(2):218–223, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Young:1969:NLW

[YBF69] A. W. Young, R. W. Buddemeier, and A. W. Fairhall. A new 60-liter water sampler built from a beer keg. Limnology and Oceanography, 14(4):634–637, July 1969. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1964:FAV

[YBW64] C. S. Yentsch, R. H. Backus, and Asa Wing. Factors affecting the vertical distribution of bioluminescence in the euphotic zone. Limnology and Oceanography, 9(4):519–524, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1962:MVL

[Yen62] Charles S. Yentsch. Measurement of visible light absorption by particulate matter in the ocean. Limnology and Oceanography, 7(2): 207–217, April 1962. CODEN LIOCAH. ISSN 0024-3590.

Yount:1961:NSC

[You61] James L. Yount. A note on stability in central Florida lakes, with discussion of the effect of hurricanes. Limnology and Oceanography, 6(3):322–325, April 1961. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1963:EPD

[YR63] Charles S. Yentsch and Carol A. Reichert. The effects of prolonged darkness on photosynthesis, respiration, and chlorophyll in the marine flagellate, dunaliella euchlora. Limnology and Oceanography, 8 (3):338–342, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Zaret:1969:PBP

[Zar69] Thomas M. Zaret. Predation-balanced polymorphism of Ceriodaphnia cornuta Sars. Limnology and Oceanography, 14(2):301–303, March 1969. CODEN LIOCAH. ISSN 0024-3590.

ZoBell:1965:NRM

[ZB65] Claude E. ZoBell and Keith M. Budge. Nitrate reduction by marine bacteria at increased hydrostatic pressures. Limnology and Oceanography, 10(2):207–214, April 1965. CODEN LIOCAH. ISSN 0024-3590.

Zeiss:1963:EPD

[Zei63] F. Ralph Zeiss, Jr. Efects of population densities on zooplankton respiration rates. Limnology and Oceanography, 8(1):110–115, April 1963. CODEN LIOCAH. ISSN 0024-3590.

Zeigler:1964:HSG

[Zei64] John M. Zeigler. The hydrography and sediments of the Gulf of Venezuela. Limnology and Oceanography, 9(3):397–411, April 1964. CODEN LIOCAH. ISSN 0024-3590.

Ziebell:1968:SRR

[ZMB68] Charles D. Ziebell, William J. Mcconnell, and Howard A. Baldwin. A sonic remote recovery device for submerged equipment. Limnology and Oceanography, 13(1):198–200, April 1968. CODEN LIOCAH. ISSN 0024-3590.

Zeigler:1965:ADP

[ZTTG65] John M. Zeigler, Sherwood D. Tuttle, Herman J. Tasha, and Graham S. Giese. The age and development of the Provincelands Hook, Outer Cape Cod, Massachusetts. Limnology and Oceanography, 10 (S1):R298–R311, April 1965. CODEN LIOCAH. ISSN 0024-3590.